### v1

# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

[ Scale = MatchLowercase, Path = fonts/, Extension = .ttf , UprightFont = *-Regular, ] 11institutetext: Institute of Theoretical Astrophysics, University of Oslo, P.O.Box 1029 Blindern, N-0315 Oslo, Norway
11email: herman.sletmoen@astro.uio.no
SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver
Herman Sletmoen
(Received $X$ / Accepted $X$)
SymBoltz is a new Julia package that solves the linear Einstein-Boltzmann equations. It features a symbolic-numeric interface for specifying equations, is free of approximation switching schemes and is compatible with automatic differentiation. Cosmological models are built from replaceable physical components in a way that scales well in model space. The modeler should simply write down their equations, and SymBoltz solves them and eliminates much of the friction in the process. SymBoltz enables up to $100\times$ shorter model definitions compared to browsing equivalent files in CLASS. Symbolic knowledge enables powerful automation of tasks, such as separating computational stages like the background and perturbations, generating the Jacobian matrix and its sparsity pattern, and interpolating arbitrary expressions from the solution. Modern implicit solvers integrate the full stiff equations at all times, reducing slowdowns by taking long time steps, reusing the Jacobian and LU-factorizing it over several time steps, and using fast linear system solvers. Automatic differentiation gives exact derivatives of any output with respect to any input, which is important for gradient-based Markov chain Monte Carlo methods in large parameter spaces, training of emulators, Fisher forecasting and sensitivity analysis. These features are useful in their own rights, but also reinforce each other in a synergy. Results agree with established codes like CLASS and CAMB. With more work, SymBoltz can grow into an integrated symbolic-numeric cosmological modeling environment with a large library of models that delivers differentiable output as fast as other codes. SymBoltz is available at https://github.com/hersle/SymBoltz.jl with single-command installation and extensive documentation, and welcomes questions, suggestions and contributions.
Key Words.:
Cosmology: theory - Methods: numerical
[…]
Cosmology is at a crossroads (Freedman, 2017). Despite the enormous success of the $\mathrm{\Lambda CDM}$ model in explaining many observations, the increasing precision of modern observations are revealing tensions between theory and observations. These suggest that the current standard model is not the ultimate truth.
On the theoretical side, this drives a search for the true cosmological model through modifications to $\mathrm{\Lambda CDM}$ (Bull et al., 2016). Efficiently exploring the space of models benefits from numerical tools that are easy to modify. Minimizing friction in the modeling process encourages relaxing model-dependent approximations, creating user-friendly interfaces and structuring codes in modular components that are easy to replace. Some of the most important such tools in the cosmological modeling toolbox are linear Einstein-Boltzmann solvers ( ?Boltzmann codes?) like CAMB (Lewis et al., 2000) and CLASS (Lesgourgues, 2011a).
On the observational frontier, next-generation surveys like the Square Kilometer Array (Dewdney et al., 2009), Vera C. Rubin Observatory (LSST Science Collaboration, 2009), Dark Energy Spectroscopic Instrument (DESI Collaboration, 2016), Simon’s Observatory (The Simons Observatory Collaboration, 2019) and Euclid (Euclid Collaboration, 2025) promise even more precise data. Setting models apart with upcoming data involves both theoretical model parameters and experimental nuisance parameters that live side-by-side in large $O(100)$-dimensional spaces (Piras et al., 2024). In high dimensions, modern Markov chain Monte Carlo methods like Hamiltonian Monte Carlo and the No-U-Turn Sampler (Hoffman & Gelman, 2011) beat the traditional Metropolis-Hastings algorithm (Hastings, 1970). These methods explore parameter space more efficiently, but need both the likelihood and its derivatives with respect to parameters. Differentiability is also increasingly important in other applications, such as forward-modeling field-level inference using simulations with initial conditions from Boltzmann solvers (Seljak et al., 2017). Automatic differentiation is a way to compute derivatives that is more accurate and can be faster than finite differences, which is an approximate and brute-force method. A differentiable Boltzmann solver is therefore a cornerstone in the next-generation cosmological modeling toolbox.
SymBoltz is a new Boltzmann solver that aims to fill these gaps, featuring a convenient symbolic-numeric interface, approximation-freeness and differentiability. At its core, a Boltzmann code solves the gravitational (Einstein) equations for a gravitational theory coupled to some particle species described by thermodynamic (Boltzmann) equations up to first perturbative order around a homogeneous and isotropic universe. It can predict cosmic microwave background (CMB), baryon acoustic oscillation (BAO) or supernova (SN) observations; or be used in longer computations, such as to generate initial conditions to non-linear $N$-body simulations of large-scale structure.
This article is structured as follows. Section 1 revisits the historical development of Boltzmann codes, sketches their structure, reviews methods for computing derivatives and motivates SymBoltz. Section 2 maps out the architecture and main features of SymBoltz and how it differs from other codes. Section 3 shows some example usage. Section 4 discusses synergies and tradeoffs in the design. Section 5 concludes with the current state of the code and future potential. The appendices list equations, implementation details, comparison settings and testing methods.
1 History and motivation
Peebles & Yu (1970) were first to numerically integrate a comprehensive set of linearized Einstein-Boltzmann equations. Their work was refined over several years, and Ma & Bertschinger (1995) established the groundwork for modern Boltzmann solvers with their code COSMICS. Seljak & Zaldarriaga (1996) soon realized that one can integrate photon multipoles by parts to reduce many differential equations to one integral solution and released CMBFAST. Their line-of-sight integration method has become a standard technique that greatly speeds up the calculation, but requires truncating the multipoles in the differential equations. This Fortran codebase has since evolved into CAMB [^1] https://camb.info/ written by Lewis et al. (2000), which is one of the two most used and maintained Boltzmann solvers today. Doran (2005a) ported CMBFAST to C++ with the fork CMBEASY, which structured the code in an object-oriented fashion and improved user-friendliness, but this project is abandoned today.
Lesgourgues (2011a); Blas et al. (2011) started the second major family of Boltzmann solvers with the birth of CLASS [^2] http://class-code.net in C. It improved performance, user-friendliness, code flexibility, ease of modification and control over precision parameters. It was the first cross-check of CAMB from an independent branch and boosted the scientific accuracy of the Planck mission.
Since then a healthy arms race have fueled refinements to CAMB and CLASS. Both codes have spawned many forks for studying alternative models and performing custom calculations. Today they are very well-made, efficient and reliable tools.
More recently, the market has seen an influx of alternative solvers that target new numerical techniques, GPU parallelization, differentiability, interactivity and symbolic computation. Refregier et al. (2017) published the first symbolic-numeric Boltzmann code PyCosmo [^3] https://pypi.org/project/PyCosmo/, which automatically generates efficient C++ ODE code from user-provided symbolic equations, performs sparsity optimization on its Jacobian matrix and avoids the use of approximation schemes to speed up and stabilize the integration. Hahn et al. (2024) published the first differentiable Boltzmann solver DISCO-EB [^4] https://github.com/ohahn/DISCO-EB in the JAX framework in Python, and relaxes approximation schemes to avoid complications and overhead when switching ODEs on GPUs. Bolt.jl [^5] https://github.com/xzackli/Bolt.jl by Li et al. (2023) accomplishes much of the same in Julia. SymBoltz.jl [^6] https://github.com/hersle/SymBoltz.jl is inspired by some of these developments.
1.1 Structure of traditional Boltzmann solvers
The full Einstein-Boltzmann equations are partial differential equations that linearize to ordinary differential equations.
In principle, the core task of a Boltzmann solver is thus very simple: to solve these ordinary differential equations (ODEs) ${\mathrm{d}\mathbf{u}}/{\mathrm{d}\tau}=\mathbf{f}(\tau,\mathbf{u};\mathbf{p})$ for some initial conditions $\mathbf{u}(\tau_{0})$ and parameters $\mathbf{p}$, including several perturbation wavenumbers, and compute desired quantities from their solution, like power spectra.
In practice, however, several properties complicate this task. First, the equations separate into computational stages that benefit substantially from being solved sequentially for performance and stability, such as the background and perturbations stages. It is common to solve each stage with interpolated input from the previous stage, and joining these can be cumbersome and fragment code. Second, the set of equations is very long and convoluted. Ideally, parts of the equations should be easily replaceable to accommodate different cosmological (sub)models. Organizing a code with a suitable modular structure that scales well in model space is hard. Third, tension between wildly different timescales that coexist in the equations make them extremely stiff and intractable for standard explicit ODE solvers. This stiffness must be massaged away in the equations or dealt with numerically in other ways. Fourth, the size of differential equations is very large and increases for more accurate description of relativistic species. Typical models with accurate treatment of photons and neutrinos need $O(100)$ equations. Fifth, the perturbations must be solved for many different wavenumbers $k$, and tradeoffs between performance and accuracy must be made.
To overcome these challenges, Boltzmann solvers are typically written in low-level high-performance languages like C, C++ and Fortran. Codes are usually tightly adapted to the pipeline-like computational structure of the problem (e.g. input $\rightarrow$ background $\rightarrow$ thermodynamics $\rightarrow$ perturbations $\rightarrow\ldots\rightarrow$ output in section 3.3 in Lesgourgues (2011a)). This makes sense for programmers and computers, but not always for modelers.
Traditional codes have nonexistent or only very thin abstraction layers around them. To modify them, users must work directly in the low-level numerical interface and should be prepared to get their hands dirty. For example, to implement a new species, one must generally modify the code in many places: input handling for any new parameters, new background equations, new thermodynamics equations, new perturbation equations, joining between each of these stages, output handling, and possibly more. This leads to fragmented code where the changes related to one species are scattered throughout the code, making it challenging and unintuitive to navigate and decipher.
More importantly, this structure scales poorly in model space. Each module becomes more polluted as more components are added. Even if they are deactivated by if statements at runtime, their mere presence in the source code reduces its readability. In turn, the whole code grows into a complex and mysterious beast.
One can alleviate this problem to some extent by instead forking the code for modified models, so the main code base is not polluted. But this just moves the problem. Forking duplicates the entire code base even though only small parts are modified. Forks are often abandoned and do not receive upstream improvements. They are also incompatible with each other unless one merges them into one code, but then one is back to the first problem.
The two-language problem amplifies the problem, as data analysis usually takes place in slower high-level languages like Python. This shapes Boltzmann solvers to rigid pipelines that must compute everything in one shot and avoid interception at all costs, in order to maximize performance in the low-level language before passing the data back to the high-level language. Some features that are really just post-processing of the ODE solutions are appended to the pipeline even if they are peripheral to the core task of an Einstein-Boltzmann solver – which is to solve the Einstein-Boltzmann equations. This includes non-linear boosting and CMB lensing, for example.
These code smells lead to big monolithic ?black boxes? that scale poorly in model space, and whose complexity grows to incorporate features beyond their original core scope. This complexity has even fueled development of specialized AI assistants for CLASS (Casas et al., 2025). While such tools are helpful, we think they are a symptom of unnecessary complexity and poor scaling with the number of models. Of course, there are also some advantages to this structure, as we discuss in section 4.
1.2 Boltzmann solver approximation schemes
The Einstein-Boltzmann equations are infamously stiff. This is a property of differential equations (3) that means their numerical solution is unstable and requires tiny step sizes with standard explicit Runge-Kutta methods. Stiffness can arise when multiple and very different time scales appear in the same problem. This is very common in cosmology, where particles interact very rapidly in a universe that expands very slowly, particularly in the tightly coupled baryon-photon fluid, for example. Stiff equations are practically impossible to integrate with explicit solvers and require special treatment.
For a long time, Boltzmann solvers have massaged away stiffness with several approximation schemes [^7] Here ?approximations? refers to schemes that switch between different equations at different times. It excludes techniques like multipole truncation and line-of-sight integration, which SymBoltz also uses. in the equations (e.g. Doran (2005b); Blas et al. (2011)):
  • tight-coupling approximation (TCA),
  • ultra-relativistic fluid approximation (UFA),
  • radiation streaming approximation (RSA),
  • Saha approximation.
These usually involve switching from one set of equations to another when some control variable that measures the applicability of the approximation crosses a threshold. This can change the unknowns in the ODE and require reinitializing it. The approximations help stability and sometimes performance of the equations and allow the use of explicit solvers.
However, approximations put much more strain on the modeler to derive and validate versions of their equations in different regimes. Furthermore, this process should generally be repeated when further changes are made to the model. The numerics also become more complicated: timeseries from each ODE solution must be stitched together, tolerances are duplicated for each separate ODE system that may respond differently to them, ODE integrators must be reinitialized, and so on.
It is a misunderstanding in the literature that it is ?difficult? or ?impossible? to solve the Einstein-Boltzmann equations without using approximations, or that they are ?unavoidable? (Ma & Bertschinger, 1995; Lewis, 2025; Lesgourgues, 2011a, respectively). This statement must only be understood in the context of explicit solvers! Implicit integrators can solve stiff equations, but seem largely underutilized by Boltzmann solvers, perhaps because they are scarce and harder to implement. CLASS has an implicit solver, but does not permit disabling all approximations Blas et al. (2011). On the contrary, new life has been breathed into this field recently, as development of modern implicit solvers and techniques like automatic and symbolic differentiation make implicit solvers more feasible and powerful.
1.3 Differentiation methods
Derivatives are important in computing. Cosmological applications are no exception. For example, algorithms that optimize likelihoods and Markov chain Monte Carlo (MCMC) samplers for Bayesian parameter inference can take advantage of derivatives of the likelihood with respect to each parameter to intelligently step in a direction where the likelihood or probability increases. In machine learning, the same applies when training neural networks emulators for cosmological observables as functions of cosmological parameters by minimizing scalar loss functions. Some cosmologies are parametrized as boundary-value problems with the shooting method, and use nonlinear root solvers like Newton’s method that need Jacobians. Jacobians are also used by implicit ODE solvers to solve for values at the next time step. Fisher forecasting uses the Hessian (double derivative) of the likelihood with respect to parameters to predict how strong parameter constraints that can be placed by data with given errors. Boltzmann solvers typically save computational resources by interpolating spectra computed on coarse grids of $k$ and $l$ to finer grids, and this can be made more precise with knowledge of derivatives with respect to $k$ and $l$. These are all examples where Boltzmann solvers or applications that use them need derivatives.
There are at least four ways to compute derivatives. Manual differentiation is the exact pen-and-paper method with differentiation rules, but is limited to simple expressions and by human error. Symbolic differentiation automates this process with computer algebra systems, but is inherently symbolic and cannot differentiate arbitrary programs nested with control flow through conditional statements and loops that depend on numerical values. Finite differentiation approximates the derivative $f^{\prime}(x)\approx(f(x+\epsilon/2)-f(x-\epsilon/2))\,/\,\epsilon$ with a small $\epsilon>0$ (here using central differences) by simply evaluating the program several times. This can differentiate arbitrary programs, but is approximate, introduces the step size $\epsilon$ as a hyperparameter that must be tuned for both accuracy and stability, and is a brute-force approach that requires $O(n)$ ($2n$ using central differences) evaluations to compute the gradient of an $n$-variate scalar $f$. Automatic differentiation (e.g. Griewank & Walther, 2008) can be understood by viewing any computer program as a (big) composite function
$\boldsymbol{f}=\boldsymbol{f}_{N}\circ\boldsymbol{f}_{N-1}\circ\cdots\circ\boldsymbol{f}_{2}\circ\boldsymbol{f}_{1}=\boldsymbol{f}_{N}(\boldsymbol{f}_{N-1}(\cdots\boldsymbol{f}_{2}(\boldsymbol{f}_{1})))$ (1)
of (many) elementary operations $\boldsymbol{f}_{i}:\mathbb{R}^{a_{i}}\rightarrow\mathbb{R}^{b_{i}}$ (think of $\boldsymbol{f}_{i}$ as the $i$-th line of code). It then numerically evaluates the chain rule
$\boldsymbol{J}=\boldsymbol{J}_{N}\cdot\boldsymbol{J}_{N-1}\cdots\boldsymbol{J}_{2}\cdot\boldsymbol{J}_{1}.$ (2)
through the Jacobian $(\boldsymbol{J}_{n})_{ij}=\partial f_{n,i}/\partial f_{n-1,j}$ of every elementary operation to accumulate the derivative of the entire program. This is numerically exact, free of precision parameters, usually requires fewer operations than finite differences, but is perhaps less intuitive, harder to implement and needs the source code to interpret the program (1) in the non-standard way (2).
Notably, while the function (1) must be evaluated inside-to-outside (right-to-left), the chain rule (2) is an associative matrix product that can be evaluated in any order. This generally changes the number of operations and is a more open-ended computational problem. Forward-mode automatic differentiation seeds $\boldsymbol{J}_{1}=\boldsymbol{1}$ (the derivative of the input with respect to itself) and multiplies $\boldsymbol{J}_{N}(\boldsymbol{J}_{N-1}(\cdots(\boldsymbol{J}_{2}(\boldsymbol{J}_{1}))))$ by ?pushing? every column of $\boldsymbol{J}_{1}$ forward through the product in the same evaluation order as $\boldsymbol{f}$. Reverse-mode first computes $\boldsymbol{f}$ in a forward pass, then seeds $\boldsymbol{J}_{N}=\boldsymbol{1}$ (the derivative of the output with respect to itself) and multiplies $((((\boldsymbol{J}_{N})\boldsymbol{J}_{N-1})\cdots)\boldsymbol{J}_{2})\boldsymbol{J}_{1}$ by ?pulling? every row of $\boldsymbol{J}_{N}$ backwards through the product. This usually makes forward-mode faster when $\boldsymbol{f}:\mathbb{R}^{a}\rightarrow\mathbb{R}^{b}$ has more outputs $b\gg a$, and reverse-mode better when there are more inputs $a\gg b$.
In practice, both modes are implemented using techniques like operator overloading or source code transformation. The former overloads every operation to accept a special number type that propagates both values and derivatives, such as dual numbers for forward-mode (e.g. Revels et al., 2016). The latter transforms the code for $\boldsymbol{f}$ into another code that computes $\boldsymbol{J}$. In any case, automatic differentiation requires access to the source code!
GravityBaryonsPhotonsMasslessneutrinosMassiveneutrinosCold darkmatterCosmo-logicalconstant Symbolic model BackgroundPerturbations Numerical problem BackgroundPerturbations Solution object CompileSolve
Figure 1: SymBoltz represents cosmological models with equations in symbolic form and grouped by physical components like the spacetime metric, gravity and particle species. This is compiled to a numerical problem that splits the model into background and perturbation stages and generates efficient numerical code for ODEs and their Jacobians. The problem is then solved and the result stored in a solution object that gives convenient access to any quantity defined in the symbolic model.
2 Code architecture and main features
SymBoltz is built somewhat differently than other Boltzmann codes. In short, it uses a symbolic-numeric abstraction interface where users enter high-level symbolic equations, and automatically compiles them to fast low-level numerical functions. It cures stiffness with implicit ODE solvers, and uses no approximation schemes to keep models simple, elegant and extensible. It is differentiable, so one can easily get accurate derivatives of any output with respect to any input. This provides rapid model prototyping, helpful abstractions and automation of typical chores when modifying other codes. SymBoltz encourages interactive usage and pursues a modular design that lets users integrate whatever they need from the package into their own applications. The end goal is to prioritize the modeler, who should be able to just write down the model equations while SymBoltz takes care of the rest.
SymBoltz is written in the Julia programming language (Bezanson et al., 2017), which attracts development of modern numerical methods and aims to resolve the two-language problem. The symbolic-numeric interface is built on the ModelingToolkit.jl (Ma et al., 2022) and Symbolics.jl (Gowda et al., 2022) packages. The compiled functions are integrated by implicit ODE solvers in DifferentialEquations.jl (Rackauckas & Nie, 2017). Everything works with automatic differentiation through ForwardDiff.jl (Revels et al., 2016), for example.
The next subsections present the design of SymBoltz in more detail, revolving around the three key features in the title. Other minor implementation details are given in appendix A. This paper is as of SymBoltz version 0.8.1. Please refer to the package documentation for definitive up-to-date information.
2.1 Symbolic-numeric interface
The core of SymBoltz is built around a symbolic-numeric interface shown in fig. 1. Variables and equations are specified in a high-level user-friendly symbolic modeling language, then ultimately compiled down to efficient low-level numerical functions that are integrated by ODE solvers. This fits the interactive and just-in-time compiled nature of Julia. A key realization is that knowledge of symbolic equations makes it possible to analyze their structure programmatically, enabling powerful transformations of the equations and automation of chores.
SymBoltz turns the layout of traditional Boltzmann solvers inside-out. While other codes are built like rigid pipelines following computational stages like the background and perturbations, SymBoltz is primarily structured around physical components of the Einstein-Boltzmann system. Full cosmological models are built by joining submodels for the metric, gravitational theory, photons, baryons, dark matter, dark energy, neutrinos and other species. Each component is a chunk of related variables and internal equations, and is unaware of other components. Interactions (e.g. Compton scattering or sourcing of gravity) are equations that connect components. This structure is made possible by the symbolic interface and would slow a purely numerical code.
Separation of computational stages is secondary and done automatically (see section 2.1.3). This makes it possible to write everything related to one component in one place, instead of scattering it across separate modules for input, background, perturbations and so on. One component can be replaced by another to construct a different cosmological model without polluting a big global model. A large part of SymBoltz is thus simply devoted to building a well-organized library of such physical components.
This modular structure scales better in model space and makes it easier to assemble both extended and reduced models. While many Boltzmann solvers hardcode baryons and photons because they are fundamental to the code, SymBoltz works naturally even with pure models with non-interacting radiation, matter and dark energy. For example, this can help understand how a modified gravity theory responds to different species without complex interactions or thermodynamics cluttering the picture.
Maybe this structure is unfamiliar to some. As a compromise, SymBoltz also provides an ?unstructured? version of the $\mathrm{\Lambda CDM}$ model with all equations and variables in one big system. This makes it easy-peasy to change anything in the model, and really shows the power of the symbolic interface: SymBoltz defines the full $\mathrm{\Lambda CDM}$ model (with GR, baryons, RECFAST recombination, photons, cosmological constant, massive and massless neutrinos) in just 277 lines of code, while the equivalent code that one must realistically browse in CLASS is spread over 10 files with 27721 lines! [^8] Counting input.{h,c}, background.{h,c}, thermodynamics.{h,c}, perturbations.{h,c} and wrap_recfast.{h,c} with wc -l. SymBoltz has less features than CLASS, but this is not even remotely close to making up for the $100\times$ simplification.
Let us see which helpful features this interface provides.
2.1.1 Automatic numerical code generation
SymBoltz automatically compiles symbolic equations to numerical code for ordinary differential equations (ODEs)
$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}\tau}=\mathbf{f}(\tau,\mathbf{u}).$ (3)
The generated code is fast and prevents users unfamiliar with the language or package from writing slow numerical code. If necessary, one can escape the standard code generation and include arbitrary numerical code, but the user is then fully responsible for its performance. This accommodates functions that cannot be written as straightforward equations, for example if one must solve a nonlinear equation for the minimum of a potential or interpolate tabulated data. The code generation deals with chores like dynamically allocating indices for each ODE state $u_{i}$, and performs optimizations such as common subexpression elimination. This is helpful as Boltzmann solvers tend to have big $\boldsymbol{f}$, and particularly with implicit ODE solvers that call $\boldsymbol{f}$ often.
2.1.2 Automatic handling of observed variables
In general, an ODE (3) admits two types of variables: unknowns $\mathbf{u}(\tau)$ are the variables that are integrated with respect to time, and observed variables are functions of the unknowns. The Einstein-Boltzmann equations are commonly formulated with many observed variables.
For example, consider the metric and gravity equations in sections A.1 and A.2 sourced by some arbitrary $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$. Here $a(\tau)$ and $\Phi(\tau,k)$ are the only unknown (differential) variables that the ODE solver must solve for, while one can ?observe? $z(\tau)$, $\mathscr{H}(\tau)$, $H(\tau)$, $\chi(\tau)$ and $\Psi(\tau,k)$ from the unknowns. Of course, it is always possible to eliminate all observeds by explicitly inserting them into the equations for the unknowns. But this kills readability, as observed variables are helpful and natural intermediate definitions that break up the system, and one may want to extract them from the solution as well. Furthermore, modified models can change the sets of unknown and observed variables (e.g. modified gravity can change the constraint equation for $\Psi$ into a differential equation). It is easier to compose models when variables do not have to be hardcoded as either unknown or observed.
SymBoltz reads in an entire system of equations like that defined by appendix A and automatically figures out which variables are unknown and observed. It then generates numerical code for solving the ODE for the unknowns (see section 2.1.1), and can lazily compute any observed variable or expression thereof from the solution of the unknowns.
The bottom line is that the user can trivially use and obtain any variable anywhere just by referring to it, whether it is unknown or observed. In other solvers one must look up the desired expression for observed variables and recompute them manually. This is tedious, error-prone and can tempts users to neglect parts of expressions (e.g. approximating $\Psi\approx\Phi$ by neglecting anisotropic stress).
2.1.3 Automatic stage separation and splining of unknowns
In principle, one can solve the Einstein-Boltzmann equations by integrating the entire system (i.e. background and perturbations) at the same time. However, the system can be broken down into sequential computational stages that each depend only on those before it. To alleviate stiffness in each stage, separate concerns by solving every equation only once (e.g. avoid recomputing the background for every perturbation mode) and to improve performance, it is common to solve the system stage-by-stage and spline variables from one stage as input to the next.
To illustrate, again consider the general relativistic equations in section A.2 sourced by some given $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$. Clearly $\Phi(\tau,k)$ and $\Psi(\tau,k)$ depend on $a(\tau)$, but $a(\tau)$ does not depend on $\Phi(\tau,k)$ or $\Psi(\tau,k)$ (this just reflects the perturbative nature of the problem). One can first solve for only $a(\tau)$ in the background, then spline $a(\tau)$ and use it as input for solving for $\Phi(\tau,k)$ and $\Psi(\tau,k)$ in the perturbations, instead of solving for all three together and repeatedly solve for $a(\tau)$ for every $k$.
SymBoltz uses the same stage separation strategy, but automates and strengthens it with its knowledge of the symbolic equations. First, the full symbolic model is split into background and perturbation systems. Cubic Hermite splines are then constructed for all unknown variables (like $a(\tau)$) to pass their solution from one stage to the next. This type of splines is perfect for interpolating ODEs, as cubic Hermite splines increase accuracy by taking both $\boldsymbol{u}(\tau)$ and $\boldsymbol{u}^{\prime}(\tau)$ into account, and $\boldsymbol{u}^{\prime}(\tau)$ is known analytically by definition (3). In contrast, observed variables (like $\Psi(\tau,k)$) are computed from the (splined) unknowns because their derivatives are not known analytically. Splining them directly would be less accurate. Modelers can simply write down a new background variable and access it in the perturbations for free. They can access any variable anywhere as if solving the entire set of equations at once, while still benefiting from splining under the hood.
The separation into background and perturbation stages just reflects the perturbative structure of the linearized Einstein-Boltzmann equations, where each order depends only on those before it, and is always guaranteed. However, they can often be broken further down: most thermodynamics (recombination) models can be separated from the background, and some variables have integral solutions (e.g. the optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\bar{\tau})\mathrm{d}\bar{\tau}$ or line-of-sight integration) that can be computed in isolation after solving the differential equations. Over time, SymBoltz aims to extend the automatic background-perturbation separation by automatically incorporating more or all stages of the equations at hand. [^9] All stages can in principle be identified from a directed dependency graph between variables (e.g. from the Jacobian matrix). Cycles in the graph correspond to interdependent equations that must be integrated together in one ODE system (e.g. background and perturbations). Leaves with differentiated variables correspond to integral solutions. This feature is a major advantage with access to the symbolic equations.
2.1.4 Automatic solution interpolation
SymBoltz integrates equations in conformal time $\tau$ for several given wavenumbers $k$, and stores the result in an object that wraps independent ODE solutions for the background and perturbations for every $k$. Conveniently, this solution object can be queried for any variable or symbolic expression
$x(\tau,k)=x(\boldsymbol{u}(\tau,k_{i})).$ (4)
at any $\tau$ and $k$. The solution object automatically expands $x$ in terms of unknowns $\boldsymbol{u}$ (if $x$ is observed), and interpolates between solved wavenumbers $k_{i}$ (if $x$ is perturbative) and in $\tau$ using the ODE solver’s built-in interpolation method (or as in section 2.1.3 if $x$ is splined). This provides the modeler easy access to any variable and expression defined in the model. The solution interpolation is also incorporated into convenient plot recipes for trivial visualization of any expressions for any $\tau$ and $k$.
2.1.5 Automatic Jacobian generation and sparsity detection
Just like SymBoltz generates code for $\boldsymbol{f}$ in the ODE (3), it can use the same symbolic equations to generate its Jacobian
$J_{ij}=\frac{\partial f_{i}}{\partial u_{j}}.$ (5)
Jacobians are crucial for solving stiff ODEs accurately and efficiently with implicit solvers (see section 2.2)! Manually coding them is an extremely tiresome and error-prone undertaking that must be repeated for model modifications. Evaluating them numerically with finite differences is approximate and expensive. Automation of this procedure is a powerful feature that again enables the modeler to focus solely on the main equations.
Another benefit is that analytical knowledge of the Jacobian lets one find its exact sparsity pattern. Numerical approaches to this are less robust and can mistake local zeros for global zeros. This permits constant factorization of sparse Jacobians, which improves performance of implicit solvers (see section 2.2). Full support for sparse Jacobians is still ongoing work, however.
SymBoltz can fall back from symbolic to automatic differentiation for the Jacobian. Finite differences can also be used.
2.1.6 Automatic change of variable (future work)
SymBoltz consistently formulates all differential equations with respect to conformal time $\tau$ as the independent variable. This is perhaps the most common parametrization in the literature. It is very natural because the equations are autonomous with respect to $\tau$ (i.e. $f(\tau,\boldsymbol{u})\rightarrow f(\boldsymbol{u})$, except some multipole truncation schemes that use $1/(k\tau)$, but these are unphysical).
Some other codes inconsistently use different independent variables in different stages. This requires manual translation into cosmic time $t$, redshift $z$, scale factor $a$ or its logarithm $b=\log a$, for example, involving differential transformations like $\mathrm{d}t=a(\tau)\,\mathrm{d}\tau$, $\mathrm{d}z=-(a^{\prime}(\tau)/a(\tau)^{2})\,\mathrm{d}\tau$, $\mathrm{d}a=a^{\prime}(\tau)\,\mathrm{d}\tau$ and $\mathrm{d}b=(a^{\prime}(\tau)/a(\tau))\,\mathrm{d}\tau$. This is another mechanical and error-prone process.
In the future, SymBoltz could automate change to another independent variable (which should be one-to-one with $\tau$) by transforming the symbolic expressions. This can be particularly useful to compute observables as a direct function of redshift, although one can invert $z(\tau)$. This is common in fits to observations.
2.1.7 Automatic unit handling (future work)
Most variables in SymBoltz are defined in internal dimensionless units (see appendix A). In the future, SymBoltz could use its symbolic pre-processing for more powerful unit features. For example, equations could be checked for dimensional validity to catch modeling mistakes. Input quantities could be automatically transformed from the user’s preferred (dimensionful) units (e.g. Mpc) to internal (dimensionless) units, and then back to the user’s units for output.
2.1.8 Automatic gauge transformation (future work)
SymBoltz is currently formulated only in the conformal Newtonian gauge. In the future, SymBoltz could conveniently transform the symbolic equations to other gauges automatically, such as the synchronous gauge used in CAMB and CLASS.
2.1.9 Automatic initial conditions (future work)
SymBoltz currently uses adiabatic initial conditions for the perturbations. In the future, SymBoltz could generate initial conditions for arbitrary models from the equations. This typically involves mechanical procedures like perturbation theory or power series expansion that could be automated.
2.1.10 Automatic approximation schemes (future work)
SymBoltz is initially designed to be free of approximation schemes (see section 2.2). In the future, SymBoltz could provide convenience utilities to help derive approximations to the equations. For example, it could automatically expand equations in power series of some smallness parameter. This could accelerate derivation of approximations for modified models to improve their performance with less effort.
2.2 Approximation-freeness
SymBoltz treats stiffness in the Einstein-Boltzmann equations with modern implicit ODE solvers that integrate the full equations at all times. It is therefore free of approximation schemes, such as the TCA, UFA, RSA and Saha approximation. This is much friendlier to the modeler, who now simply has to provide a single set of equations instead of deriving approximations, validating them and dealing with related chores.
However, implicit ODE solvers are more expensive than explicit methods. In particular, at every time step they must solve a nonlinear system of equations for the next unknowns. This is usually done with Newton’s method that requires the Jacobian of the nonlinear system, which in turn involves the ODE Jacobian (5). Automatic, accurate and efficient computation of the Jacobian (see section 2.1.5) makes this less of a problem.
As the size of the ODE and its Jacobian increases (e.g. larger $l_{\text{max}}$), solving the large system of equations at every time step becomes a bottleneck. One should then transition from dense to sparse matrix methods (this is still ongoing work) and use performant matrix methods when solving the nonlinear system. By default, SymBoltz uses the Rodas4P (Steinebach, 1995) solver for the background (stiff when thermodynamics is included) and KenCarp4 (Kennedy & Carpenter, 2003) for the perturbations. These are both implicit solvers from DifferentialEquations.jl. We discuss this aspect further in section 4.
In the long run, SymBoltz could also implement approximation schemes to maximize speed (see section 2.1.10). However, the goal is to keep the code primarily approximation-free and for all approximations to be secondary and optional.
2.3 Differentiability
SymBoltz is compatible with automatic differentiation. One can obtain the derivative of any output quantity, such as all ODE variables and derived spectra, with respect to any combination of input parameters, like the reduced Hubble parameter $h$ or cold dark matter density parameter $\Omega_{c0}$. Automatic differentiation is also used several places internally, such as for computing the Jacobian for implicit ODE solvers (as an alternative to the fully symbolic approach) and in the shooting method (which is a nonlinear equation solver that wraps around the ODE solver).
Currently SymBoltz is only well-tested with forward-mode dual numbers through ForwardDiff.jl. In particular, support for reverse-mode is very attractive for scalar loss applications (e.g. likelihoods), but left for future work (see section 4.3).
Figure 2: Included plot recipes in SymBoltz make it trivial to visualize any symbolic variable or expression thereof from a solution of the Einstein-Boltzmann equations. This plot was made with one short line of code per subplot. Wavenumbers $k$ are in units of $H_{0}/c$.
3 Examples
The best way to illustrate the features in section 2 is perhaps through some examples. The full code is as of SymBoltz version 0.8.1 and available in a notebook in the project repository.
3.1 Basic usage workflow
Most usage follows a model–problem–solution workflow: {codebox}
using SymBoltzM = ΛCDM(lmax = 10)pars = Dict( M.γ.T₀ => 2.7, M.b.Ω₀ => 0.05, M.b.rec.Yp => 0.25, M.ν.Neff => 3.0, M.c.Ω₀ => 0.27, M.h.m_eV => 0.06, M.I.ln_As1e10 => 3.0, M.I.ns => 0.96, M.g.h => 0.7)prob = CosmologyProblem(M, pars)ks = [4, 40, 400, 4000] # k / (H₀/c)sol = solve(prob, ks)The model–problem–solution split achieves three distinct goals.
First, a symbolic representation of the $\mathrm{\Lambda CDM}$ model M is created. This is a standalone object designed to be interactively inspected and modified independently of numerics (see section 3.2). It contains every variable, parameter and equation of the cosmological model structured as a graph of submodels for each logically distinct (physical) component: the metric $g$, the gravitational theory $G$, photons $\gamma$, massless neutrinos $\nu$, massive neutrinos $h$, cold dark matter $c$, baryons $b$, the cosmological constant $\Lambda$ and inflation $I$. For example, equations(M) shows all model equations; equations(M.G) shows only the gravitational ones; M.g.a gives a handle to the scale factor variable $a(\tau)$ that ?belongs? to the metric $g$; M.b.Ω₀ is the density parameter $\Omega_{b0}$ in the baryon component $b$; and parameters(M) lists parameters of the model that the user may set. Everything displays with -compatibility in notebooks to encourage interactive use.
Second, the symbolic model is compiled to a numerical problem prob with parameters pars. This is an expensive operation where the ?magic? in section 2.1 happens: equations are checked for consistency and split into background and perturbation stages; observed and unknown variables are distinguished; fast ODE code is generated; background unknowns in the perturbations are replaced by splines; the Jacobian matrix is generated in numerical/analytical and dense/sparse form; and initial values are computed. One can customize whether stages should be separated, how the Jacobian should be generated, and so on. This is a separate step because it performs final transformations on the model M once the user has finished modifying and committed to it.
Third, solve(prob, ks) solves the background and perturbations for the requested wavenumbers ks. Omitting ks solves only the background. The resulting solution object sol provides convenient access to all model variables. Internally, it stores the values of all ODE unknowns at the time steps taken by the (adaptive) solvers and for every requested wavenumber. However, it can be queried with any time, wavenumber and symbolic expression, and will automatically compute it from the unknowns and interpolate between stored times and wavenumbers (see sections 2.1.2, 2.1.3 and 2.1.4). For example, calling sol(1.0, 2.0, g.Φ+g.Ψ) will compute $\Phi+\Psi$ at $k=1\,H_{0}/c$ and $\tau=2\,H_{0}^{-1}$ by expanding it in terms of unknowns according to the equations in section A.2. It is also possible to compute arbitrary expressions over grids of $k$ and $\tau$. The convenient solution interface is also integrated into plot recipes for effortless visualization, as shown in fig. 2.
Other Boltzmann solvers save only a hardcoded set of variables, such as only the unknowns. Observed variables must generally be recomputed manually, even though they are already expressed somewhere in the code. This is cumbersome, introduces more potential for user error and adds unnecessary friction in the modeling process. SymBoltz has higher ambitions than simply spitting out a table with some selected variables, and is designed to interactively make modifications to the model with minimal changes and easily inspect its impact on any output.
Figure 3: Matter and CMB ($T$, TE and $E$) power spectra computed by SymBoltz (SB; colored lines) compared to CLASS (CL; grey dashes) with relative errors $\text{rel.err.}=P_{k}^{\text{SB}}/P_{k}^{\text{CL}}-1$ and $\text{rel.err.}=C_{l}^{\text{SB}}/C_{l}^{\text{CL}}-1$ for the $\Lambda\text{CDM}$ model. CLASS uses the precision parameters in appendix B.
3.2 Modifying models
Suppose we want to replace the cosmological constant species $\Lambda$ in the model M with another dark energy species. A well-known example is dynamical $w_{0}w_{a}$ dark energy (Chevallier & Polarski, 2001; Linder, 2003) with equation of state [^10] SymBoltz already includes $w_{0}w_{a}$ as an available species.
$w(\tau)=w_{0}+w_{a}(1-a(\tau)).$ (6)
In this case, the continuity equation
$\rho^{\prime}(\tau)=-3\mathscr{H}(\tau)\rho(\tau)(1+w(\tau))$ (7)
admits the analytical solution (by an ansatz of the same form)
$\rho(\tau)=\rho(\tau_{0})a(\tau)^{-3(1+w_{0}+w_{a})}\exp(-3w_{a}(1-a(\tau))).$ (8)
To implement this species including perturbations following de Putter et al. (2010), simply write down the symbolic variables, parameters, equations and initial conditions: {codebox}
g, τ, k = M.g, M.τ, M.ka, ℋ, Φ, Ψ = g.a, g.ℋ, g.Φ, g.ΨD = Differential(τ)@parameters w₀ wₐ cₛ² Ω₀ ρ₀@variables ρ(τ) P(τ) w(τ) cₐ²(τ) δ(τ,k) θ(τ,k) σ(τ,k)eqs = [ w ~ w₀ + wₐ*(1-a) ρ₀ ~ 3*Ω₀ / (8*Num(π)) ρ ~ ρ₀ * a^(-3(1+w₀+wₐ)) * exp(-3wₐ*(1-a)) P ~ w * ρ cₐ² ~ w - 1/(3ℋ) * D(w)/(1+w) D(δ) ~ 3ℋ*(w-cₛ²)*δ - (1+w) * ( (1+9(ℋ/k)^2*(cₛ²-cₐ²))*θ + 3*D(Φ)) D(θ) ~ (3cₛ²-1)*ℋ*θ + k^2*cₛ²*δ/(1+w) + k^2*Ψ σ ~ 0]initialization_eqs = [ δ ~ -3//2 * (1+w) * Ψ θ ~ 1//2 * k^2*τ * Ψ]X = System(eqs, τ; initialization_eqs, name = :X)Everything is packed down into the $w_{0}w_{a}$ component X. Note that SymBoltz encourages Unicode symbols to maximize similarity between equations and code (e.g. Ω₀ over Omega_0) and to easily display in compatible environments (e.g. notebooks).
This is all the user has to do! Variables are automatically split into and carried across the background and perturbation stages, hooks for setting input parameters and getting arbitrary output variables are automatically available, and so on. The modification simply consists of writing down the equations verbatim. It could not have been more compact and to the point.
A similar modification to CLASS is more involved, for example. It would require at least: reading new parameters in input.c; declaring new background and perturbation variables in background.h and perturbations.h; solving background equations, storing desired output and coupling them to gravity in background.c; and recomputing or looking up background variables, solving perturbation equations and storing desired output and coupling them to gravity in perturbations.c. The changes should be if-guarded correctly to ensure the code works both with and without $w_{0}w_{a}$.
A full $w_{0}w_{a}\text{CDM}$ model and problem can now be built by replacing the cosmological constant species Λ in $\Lambda\text{CDM}$ by the $w_{0}w_{a}$ species X: {codebox}
M = ΛCDM(Λ = X, name = :w₀wₐCDM)pars[M.X.w₀] = -0.9pars[M.X.wₐ] = 0.2pars[M.X.cₛ²] = 1.0prob = CosmologyProblem(M, pars)We proceed with this new model and problem.
3.3 Computing power spectra
SymBoltz can compute the matter power spectrum $P(k)$ and the angular CMB power spectra $C_{l}^{\text{TT}}$, $C_{l}^{\text{TE}}$ and $C_{l}^{\text{EE}}$: {codebox}
using Unitful, UnitfulAstro # for Mpc unitks = 10 .^ range(-4, 1, length=200) / u"Mpc"Ps = spectrum_matter(prob, ks)ls = 10:10:2000Cls = spectrum_cmb([:TT, :TE, :EE], prob, ls)Calling spectrum_matter and spectrum_cmb with prob will automatically select a grid of $k$ to solve the perturbations for, and interpolate between them when computing the spectra. This is a common interpolation trick in Boltzmann solvers to integrate fewer perturbation modes. The functions can be called with the solution sol instead, but will then interpolate only between the $k$ that sol was solved for, leaving it to the user to ensure sufficient sampling density. Figure 3 shows that the spectra agree with CLASS to 11‰-11% or better for the chosen parameters.
3.4 Differentiable Fisher forecasting
Fisher forecasting is a powerful technique for predicting how strong parameter constraints that can be placed by data with given errors. It requires accurate derivatives and is a nice application for automatic differentiation.
Near a peak $\boldsymbol{\theta}_{0}$, where derivatives vanish, a log-likelihood function of parameters $\boldsymbol{\theta}$ is approximated by the Taylor series
$\log L(\boldsymbol{\theta})\approx\log L(\boldsymbol{\theta}_{0})-\sum_{i,j}F_{ij}(\theta_{i}-\bar{\theta}_{i})(\theta_{j}-\bar{\theta}_{j}),$ (9)
where $F$ is the Fisher (information) matrix with elements
$F_{ij}=-\frac{1}{2}\frac{\partial^{2}\log L}{\partial\theta_{i}\,\partial\theta_{j}}\bigg\rvert_{\,\boldsymbol{\theta}=\boldsymbol{\theta}_{0}}.$ (10)
Intuitively, $F$ measures how sharp the peak is, or how sensitive $\log L$ is to changes in different directions in parameter space. Fisher forecasting is powered by the Cramér-Rao bound, which guarantees that $F^{-1}_{ij}$ is a lower bound $\big\lvert C_{ij}\big\rvert\geq\big\lvert F^{-1}_{ij}\big\rvert$ for the covariance $C_{ij}$ between model parameters $\theta_{i}$ and $\theta_{j}$. The inequality is more saturated the better the likelihood approximation (9) is (i.e. the ?more Gaussian? the probability distribution is, for which the expansion is exact). Thus, computing $F_{ij}$ at a peak and inverting it gives the tightest parameter constraints one can hope for.
Since $F_{ij}$ depends on derivatives of $\log L$, Fisher forecasting traditionally involves error-prone finite differences and step size tuning. This problem is avoided with automatic differentiation.
To demonstrate differentiable Fisher forecasting with SymBoltz, we make the best possible CMB (TT) measurement: one of $\bar{C}_{l}$ over the full sky with errors only due to cosmic variance
$\sigma_{l}=\sqrt{\frac{2}{2l+1}}\bar{C}_{l}.$ (11)
Assuming the measurements of each $\bar{C}_{l}$ are normally distributed and uncorrelated, the log-likelihood for this experiment is $\chi^{2}$:
$\log L(\boldsymbol{\theta})=-\frac{1}{2}\sum_{l}\left(\frac{C_{l}(\boldsymbol{\theta})-\bar{C}_{l}}{\sigma_{l}}\right)^{2}.$ (12)
In this case, the Fisher matrix (10) becomes
$F_{ij}=\sum_{l}\frac{\partial C_{l}}{\partial\theta_{i}}\frac{1}{\sigma_{l}^{2}}\frac{\partial C_{l}}{\partial\theta_{j}}\bigg\rvert_{\,\boldsymbol{\theta}=\boldsymbol{\theta}_{0}},$ (13)
evaluated in some fiducial cosmology $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$.
The derivatives $\partial C_{l}/\partial\theta_{i}$ are hard to compute and perfect candidates for automatic differentiation with ForwardDiff.jl: {codebox}
using ForwardDiff: jacobianvary = [ M.g.h, M.c.Ω₀, M.b.Ω₀, M.b.YHe, M.ν.Neff, M.h.m_eV, M.X.w₀, M.X.wₐ, M.I.ln_As1e10, M.I.ns,]genprob = parameter_updater(prob, vary)ls, ls′ = 100:1:1000, 100:25:1000Cl(θ) = spectrum_cmb(:TT, genprob(θ), ls, ls′)θ₀ = map(par -> pars[par], vary)dCl_dθ = jacobian(Cl, θ₀)Here vary orders the subset of parameters to differentiate with respect to, and genprob is a function that generates a new problem with updated parameters θ. Then Cl specifies a function that computes $C_{l}(\boldsymbol{\theta})$ for ls, interpolating from a coarser grid ls′. This machinery just converts from SymBoltz’ parameter-to-value mapping to a function of an array of parameter values, as jacobian requires a pure mathematical function $\boldsymbol{f}:\mathbb{R}^{a}\rightarrow\mathbb{R}^{b}$ to differentiate. Finally $\partial C_{l}/\partial\theta_{i}$ is computed with forward-mode automatic differentiation and stored in dCl_dθ, as shown in fig. 4.
Figure 4: Normalized derivatives $(\partial C_{l}/\partial\theta_{i})/C_{l}$ of a CMB $T$ power spectrum with respect to cosmological parameters $\theta_{i}$ from SymBoltz and automatic differentiation (AD; colored lines) versus CLASS and central finite differences (FD; gray dashes). CLASS uses the precision parameters in appendix B and finite differences with $5\%$ relative step sizes.
Figure 5: Marginalized 68% and 95% 2D confidence ellipses for parameter constraints from a Fisher forecast on a cosmic variance-dominated CMB $T$-only survey using the derivatives in fig. 4.
It is now trivial to compute the Fisher matrix (13). Inverting it forecasts the constraints in fig. 5. They are in excellent agreement with finite difference results from CLASS, but these required significant tuning of precision parameters and step sizes.
3.5 Differentiable MCMC sampling with supernova data
Finally, we show that SymBoltz can output differentiable results for gradient-based Markov chain Monte Carlo (MCMC) samplers, like the No-U-Turn Sampler (NUTS). This state-of-the-art method uses likelihood gradients to move more intelligently in parameter space than the ?blind? Metropolis-Hastings algorithm.
We predict apparent magnitudes $m(z)$ of Type Ia supernovae as a function of redshift $z$ and fit them to 1048 observations $(z_{i},m_{i})$ from the Pantheon dataset (Scolnic et al., 2018). Such supernovae are standard candles with constant absolute magnitude $M\approx-19.3$. Their apparent magnitudes are $m(z)=M+\mu(z)$, where $\mu(z)=5\lg(d_{L}(z)/10\text{ pc})$ is the distance modulus from the background-derived luminosity distance [^11] This expression is valid for any $\Omega_{k0}$ with complex $\operatorname{sinc}(x)=\sin(x)/x$, but can be split into branches for positive, negative and zero $\Omega_{k0}$.
$d_{L}(z)=\frac{c}{H_{0}}\frac{\chi(z)}{a(z)}\operatorname{sinc}\Big(H_{0}\sqrt{-\Omega_{k0}}\,\chi(z)\Big).$ (14)
We define the likelihood by a multivariate normal distribution of the 1048 supernovae with (constant) covariance matrix $C$ given in detail by Scolnic et al. (2018). To compute it, we interface SymBoltz with the probabilistic programming framework Turing.jl (Fjelde et al., 2025), which allows a high-level description of the probabilistic model equivalent to the log-likelihood
$\log L=-\frac{1}{2}\sum_{i,j}\big(m(z_{i})-m_{i}\big)C^{-1}_{ij}\big(m(z_{j})-m_{j}\big).$ (15)
We use the NUTS sampler in Turing. The code for this example can be found in the paper’s notebook. For the flat $w_{0}\text{CDM}$ model ($\Omega_{k0}=0$, $w_{a}=0$, fixed $\Omega_{r0}$) it gives the constraints in fig. 6. The results are not interesting, but show that differentiable parameter inference works with SymBoltz. Performance improvements are needed to make this viable also for perturbation-derived spectra.
Figure 6: Parameter inference on 1048 Type Ia supernovae from the full Pantheon supernova dataset, using 5000 MCMC samples with the gradient-based NUTS sampler in Turing.jl and differentiable theory predictions from SymBoltz.jl.
4 Discussion of design synergies and tradeoffs
A symbolic-numeric interface, lack of approximations and differentiability are interesting features in their own rights. They also constitute a self-reinforcing synergy explained in fig. 7. Like any Boltzmann solver, SymBoltz makes some tradeoffs in its design.
Symbolic-numericinterfaceFree ofapprox-imationsAutomaticdifferentiationDetect sparsity of JJCalculate $J$ robustlyEasy to specify model
Figure 7: A symbolic-numeric interface, approximation-freeness and differentiability are three important features of SymBoltz that form a synergy. For example: automatic differentiation helps to calculate the ODE Jacobian robustly; which is needed by implicit solvers to integrate stiff ODEs without approximations; which makes it easier to write down models in a simple symbolic form; which can be used to detect the sparsity pattern of the Jacobian; which in turn increases the efficiency of the implicit ODE solver. This synergy creates a self-reinforcing design.
4.1 Symbolic vs. numeric interface
As explained in sections 1.1 and 2.1, modifications to most Boltzmann solvers are made by editing the low-level numerical source code directly, while changes to SymBoltz are made in a high-level symbolic interface that abstracts away internal details.
We think three properties make the linearized Einstein-Boltzmann equations particularly attractive for symbolic abstraction. First, on the ?outside?, they are fundamentally one (large) system of equations with a predictable general structure. Second, on the ?inside?, they are complicated to solve in multiple stages due to their structure, stiffness and need for speed. Third, they are heavily subject to modifications as the true cosmological model is unknown. To simplify the modeling process, these aspects make it valuable to abstract the ?inside? from the ?outside?.
Purely numerical codes generally put the programmer first. They are adapted to the computational structure of the problem. Users have full freedom to tweak problem-specific details, such as implementing complex approximation schemes. Without abstraction layers, the code says exactly what it does. This can be worthwhile for custom-tailoring codes to ?workhorse? cosmological models like $\mathrm{\Lambda CDM}$ for maximum performance and demanding MCMC analyses. However, it can result in an overwhelming monolithic design that is hard to read and modify.
The symbolic interface in SymBoltz prioritizes the modeler. Its design matches the physical structure of the model. Manual chores are automated with compile-time logic on the equations. Users can more easily change equations without understanding internals provided that they can be written in a straightforward form that is compatible with the symbolic interface. This is not a major restriction for the Einstein-Boltzmann equations, as they have a plain and predictable structure when linearized and free of approximations. Nevertheless, to bridge this gap, SymBoltz lets users call arbitrary numerical functions that escape the symbolic-to-numeric compilation, if necessary.
This abstraction layer works only if fits the structure of the underlying equations. A key to envision the design was to look past the computational pipeline structure of Boltzmann solvers (e.g. background $\rightarrow$ perturbations $\rightarrow\ldots$), and rather organize the code primarily by physical components like the metric, gravitational theory and particle species. Components are self-contained and joined into full cosmological models. As more models are added, the complexity of this modular design stays constant, while monolithic ?all-in-one? codes become complicated.
This structure can be unfamiliar to those used to the layout of other codes. The modular structure scales best with a growing number of submodels for individual physical components. But SymBoltz also provides a full unstructured $\mathrm{\Lambda CDM}$ model where everything is merged into one large system, which is easier to work with if one wants to modify arbitrary small parts of the equations. This leaves the choice to the user, who can select between the best of both worlds.
Mixing symbolic and numerical computing can be susceptible to performance problems. Numerics should be fast and symbolics is usually slow, so ?perceived performance? suffers if the two are interleaved. But SymBoltz disentangles the expensive symbolic operations in section 2.1 from numerics, so they are performed only before and after performance-critical tasks.
We hope the design of SymBoltz provides a platform for easy exploration of alternative models, regardless of which sector of the equations one is interested in modifying.
4.2 Approximation-freeness vs. performance
As seen in section 1.2, traditional codes reduce stiffness by approximating the equations for explicit integrators, while SymBoltz solves the full stiff system with implicit methods.
Without approximations, SymBoltz needs only one set of equations to solve. These equations are easy to write down in pure symbolic form, which pairs nicely with SymBoltz’ high-level symbolic interface. With approximations, traditional solvers need more complicated infrastructure to switch between versions of the same equations. This fits well to a low-level and fully numerical code that maximizes implementation freedom, like CLASS.
Approximation schemes do not only alleviate stiffness, but also improve performance by reducing the ODE size. However, Lesgourgues & Tram (2011); Hahn et al. (2024) found that the TCA and UFA provide only marginal speedups compared to use of implicit solvers. On the other hand, the RSA can provide significant speedups with high $l_{\text{max}}$ (Lesgourgues & Tram, 2011; Moser et al., 2022). However, more load is put on the modeler to derive, implement and validate multiple approximations. This process must generally be repeated for modified models that can easily reintroduce stiffness or invalidate the approximations.
Implicit solvers take more expensive steps than explicit ones. They solve a nonlinear system for the next unknowns at every time step, which gets costly as the ODE size grows (e.g. higher $l_{\text{max}}$). This is often done with Newton’s method, which iterates over linear matrix solutions $A\boldsymbol{x}=b$, where $A$ involves the ODE Jacobian $J$. But several factors mitigate this slowdown.
First and foremost, implicit solvers take longer and fewer steps due to better stability properties. This is highly solver-dependent, however: two different implicit solvers can perform very differently on one stiff problem. DifferentialEquations.jl lets SymBoltz easily switch between a large suite of implicit solvers. The background (with stiff thermodynamics) is solved robustly with Rodas4P/Rodas5P, and the perturbations efficiently with TRBDF2/KenCarp4/Rodas5P for low/medium/high precision. Notably, CLASS includes and defaults to a custom implicit solver ndf15 and performs great (Lesgourgues & Tram, 2011). In contrast, SymBoltz can choose between a range of compatible solvers to fulfill different performance and precision requirements.
Second, a key optimization for implicit solvers is to compute $J$ efficiently, reuse it over time steps and LU-factorize $A$ to speed up successive linear system solutions until Newton’s method no longer converges due to outdated Jacobians. The TRBDF2 and KenCarp4 solvers do this. When needed, $J$ is updated with the method in section 2.2, which is cheaper than finite differences. The linear matrix solver used in Newton’s method also matters, as they are among the most thoroughly optimized numerical algorithms. DifferentialEquations.jl interoperates with LinearSolve.jl, which lets SymBoltz easily swap the linear solver in the ODE solver. On our hardware we find that performance improves by $5\times$ with a recursive LU factorization algorithm from RecursiveFactorization.jl or Intel’s (proprietary) Math Kernel Library MTK.jl over Julia’s default OpenBLAS backend.
Third, sparse matrix methods can speed up the linear solver as the system grows in size and the fraction of zeros increases. SymBoltz can find the exact sparsity pattern from the symbolic Jacobian in section 2.1.5, and LinearSolve.jl provides several sparse matrix methods. Support for this is still ongoing work.
Although SymBoltz sacrifices some performance to get rid of approximations, these countermeasures make the impact less severe than one might fear. In return, the approximation-free structure is a major simplification and pairs well with the high-level interface for simple symbolic equations. There is still room to improve performance without resorting to approximations by tuning precision tolerances, supporting sparse matrix methods, exploring implicit-explicit (IMEX) solvers (e.g. Kennedy & Carpenter (2003)) and smarter sampling of $k$ and $l$, for example.
Further performance optimizations and comparisons are left for future work. SymBoltz is fast enough for single runs and interactive usage, but not yet for MCMC analyses with perturbation-derived quantities. Of course, SymBoltz can also implement approximation schemes in the future, but its symbolic nature can make this harder than in other codes. As a counterweight to other codes, SymBoltz’ primary mission will be to solve full stiff equations and keep approximations optional and secondary.
4.3 Forward-mode vs. reverse-mode automatic differentiation
At this time SymBoltz is only well-tested with forward-mode automatic differentiation. However, we saw in section 1.3 that reverse-mode is more attractive for applications with more inputs than outputs. This is the case for popular applications with scalar likelihood or loss functions, such as MCMC parameter inference and training neural network emulators. Reverse-mode could be particularly powerful for parameter inference with next-generation surveys, where experimental nuisance parameters must also be sampled in large $\mathcal{O}(100)$-dimensional parameter spaces, even if eventually marginalized over. However, forward-mode has better characteristics when performing sensitivity analyses. For example, computing $\partial P(k;\boldsymbol{\theta})/\partial\theta_{i}$ or $\partial C_{l}(\boldsymbol{\theta})/\partial\theta_{i}$ (e.g. Fisher forecasting in section 3.4) is faster with forward-mode, since one typically wants $\mathcal{O}(100\text{-}1000)$ $k$ or $l$, but only has $\mathcal{O}(10)$ parameters.
Of course, automatic differentiation is just an additional feature. One can run the code with or without automatic differentiation, or use finite differences instead. It does not pose any tradeoff, but robust support for both forward-mode and reverse-mode remains a future goal that would make SymBoltz more powerful for different applications. This situation will improve as the differentiable Julia ecosystem continues to evolve.
5 Conclusion and future potential
SymBoltz is a fresh Julia package for solving the linearized Einstein-Boltzmann equations. It relaxes all approximation switching schemes found in other codes and solves a single set of stiff equations at all times with implicit integrators, and combats the performance loss with modern and efficient implicit ODE solvers, optimized linear system solvers and (soon) sparse matrix methods. It is differentiable, so one can get accurate derivatives of any output quantity with respect to any input parameter. This enables modelers to rapidly prototype new models by straightforwardly writing down new variables and equations.
Version 0.8.1 features the metric in the conformal Newtonian gauge, General Relativity, cold dark matter, photons, baryons and RECFAST recombination, massless and massive neutrinos, the cosmological constant and $w_{0}w_{a}$ dark energy, and computes luminosity distances and matter and CMB spectra. It also has some rudimentary models for Brans-Dicke gravity, quintessence dark energy and curved geometry, but these are not complete yet. We think the modular design makes it very easy to add new models and compute other quantities.
Forward-mode automatic differentiation is well-tested, while support for reverse-mode is a future goal. This would make SymBoltz very powerful for scalar loss applications like MCMCs.
With some more work, SymBoltz could grow into a fully integrated symbolic-numeric and differentiable cosmological modeling environment. For example, numerical code generation could be extended from linear equations to non-linear $N$-body simulations in a consistent and unified framework. This could alleviate the $N$-language problem ($N\gtrsim 2$) prevalent in cosmological modeling, which glues together convenient high-level languages like Python with performant low-level languages like C and Fortran, and sometimes symbolic work in a CAS like Mathematica.
SymBoltz shows that it is possible to create a symbolic-numeric, approximation-free and differentiable Boltzmann solver with modern numerical techniques. It excels at model simplicity and ease of modification. More work is needed to make it as feature-complete and fast as CAMB and CLASS that have been refined over many years. A major goal is to compute values and derivatives of perturbation-derived spectra fast enough for use with gradient-based MCMC samplers. Implicit ODE solvers, nonlinear and linear equation solvers, sparse matrix methods and differentiability are all demanding numerical techniques. Making them work robustly and efficiently together is challenging, but the current state shows promise. SymBoltz is built on (and has contributed to) several evolving Julia packages, and will continue to grow both on its own and with improvements to its dependencies.
SymBoltz is easy to install from https://github.com/hersle/SymBoltz.jl. Documentation is available there, and the code is tested with continuous integration to ensure it remains correct and up-to-date (see appendix C). Anyone is encouraged to ask questions, give feedback, open issues and contribute pull requests in the repository! We hope SymBoltz offers valuable competition and new ideas on the Boltzmann solver market.
Acknowledgments
I thank Aayush Sabharwal and Christopher Rackauckas for developing and maintaining ModelingToolkit.jl and DifferentialEquations.jl and answering questions. I thank Hans A. Winther for helpful suggestions, testing the code and giving feedback on it and this manuscript. I thank Julien Lesgourgues, Thomas Tram and others for developing CLASS, which has inspired SymBoltz.
Appendix A List of equations and practical implementation details
This appendix summarizes the equations that define the standard $\mathrm{\Lambda CDM}$ model in SymBoltz and comments on their practical implementation. We hope this can be a useful reference for others. The list is structured like SymBoltz with one component per subsection. Variables are in units where $G=c=H_{0}=1$? and temperatures are in $\mathrm{K}$, unless otherwise is stated. These units are chosen because $G$, $c$ and $H_{0}$ can be divided out from the Einstein equations as natural units, but this requires some conversion in the recombination equations that depend explicitly on $H_{0}$. In other words, times are in units of $1/H_{0}$, distances in $c/H_{0}$ and masses in $c^{3}/H_{0}G$. When $G$, $c$ or and $H_{0}$ appear explicitly in the equations, they are in SI units and used only to convert from SI units into the dimensionless units. The equations closely follow the conventions in the seminal paper by Ma & Bertschinger (1995) and very closely matches the source code of SymBoltz with Unicode characters. It is far outside the scope of this paper to derive the equations and explain the meaning of every variable.
The independent variable is conformal time $\tau$, and all derivatives $^{\prime}=\mathrm{d}/\mathrm{d}\tau$ are with respect to it (in units of $1/H_{0}$). By default, integration starts from the early time $\tau=\tau_{i}=10^{-6}$, and is terminated when the scale factor crosses $a=1$, and the corresponding time $\tau=\tau_{0}$ is labeled today.
A.1 Metric and spacetime $(g)$
SymBoltz is currently only formulated in the conformal Newtonian gauge, with this metric and related quantities:
$g_{0i}=g_{i0}=-a^{2}(1+2\Psi)\delta_{0i},\qquad g_{ij}=a^{2}(1-2\Phi)\delta_{ij},\qquad z=\frac{1}{a}-1,\qquad\mathscr{H}=\frac{a^{\prime}}{a},\qquad H=\frac{\mathscr{H}}{a},\qquad\chi=\tau_{0}-\tau.$
Here $\mathscr{H}$ and $H$ are the conformal and cosmic Hubble factors (in units where $\mathscr{H}_{0}=H_{0}=1$). The scale factor $a$ is related to redshift $z$, and $\chi$ is the lookback time from today that appears in some integral solutions. SymBoltz is currently restricted to a flat spacetime.
A.2 General relativity $(G)$
The gravitational theory of the $\mathrm{\Lambda CDM}$ model is General Relativity governed by the Einstein field equations $G_{\mu\nu}=8\pi T_{\mu\nu}$. By default, SymBoltz solves for the metric variables $a$, $\Phi$ and $\Psi$ with $(\mu,\nu)=(0,0)$ in the background (1st Friedmann equation) and $(\mu,\nu)=\{(0,0),(i,j)\}$ in the perturbations:
$a^{\prime}=\sqrt{\frac{8\pi}{3}\rho}\,a^{2},\qquad\Phi^{\prime}=-\mathscr{H}\Psi-\frac{k^{2}}{3\mathscr{H}}\Phi-\frac{4\pi}{3}\frac{a^{2}}{\mathscr{H}}{\delta\rho},\qquad\Psi=-\Phi-12\pi\bigg(\frac{a}{k}\bigg)^{2}\Pi.$
Note that it is also possible to evolve other redundant combinations of the Einstein equations (such as the acceleration equation). The equations are coupled to total densities $\rho$, $\delta\rho$, pressures $P$ and anisotropic stresses $\Pi$ for an arbitrary set of species $s$ that are present in the cosmological model:
$\rho=\sum_{s}\rho_{s},\qquad P=\sum_{s}P_{s},\qquad\delta\rho=\sum_{s}\delta\rho_{s}=\sum_{s}\delta_{s}\rho_{s},\qquad\delta P=\sum_{s}\delta P_{s}=\sum_{s}\delta_{s}\rho_{s}c_{s,s}^{2},\qquad\Pi=\sum_{s}\Pi_{s}=\sum_{s}(\rho_{s}+P_{s})\sigma_{s}.$
We emphasize that the gravity component is completely unaware of all particle species and makes no assumptions about them. It only reacts to total stress-energy components. All species must therefore define $\rho$, $P$, $\delta\rho$, $\delta P$ and $\Pi$ explicitly, even if zero. This requirement is somewhat pedantic, but helps isolate components from each other for greater reuse when composing models.
The scale factor $a$ is initialized as the nonlinear solution of the Friedmann equation constrained to $\mathscr{H}=1/\tau$ (motivated by its radiation-dominated solution $a=\sqrt{\Omega_{r0}}\,\tau$). The constraint potential is initialized to $\Psi=20C/(15+4f_{\nu})$ with the (arbitrary) integration constant $C=1/2$ and initial energy density fraction $f_{\nu}=(\rho_{\nu}+\rho_{h})/(\rho_{\nu}+\rho_{h}+\rho_{\gamma})$ of all (massless and massive) neutrino species relative to all species that are radiative at early times. The evolved potential $\Phi$ is initialized accordingly from the constraint equation (the found solution is close to $\Phi=(1+2f_{\nu}/5)\Psi$, but providing both $\Phi$ and $\Psi$ explicitly leads to overdetermined initialization equations due to the constraint equation).
The next sections present the $\mathrm{\Lambda CDM}$ section of the ?library of species? that are available in SymBoltz.
A.3 Cold dark matter $(c)$
Cold dark matter is a non-relativistic and non-interacting species that follows very simple equations:
$w=0,\qquad{c_{s}^{2}}=w,\qquad P=0,\qquad\rho=\frac{\rho_{0}}{a^{3}},\qquad\delta^{\prime}=-\theta+3\Phi^{\prime},\qquad\theta^{\prime}=-\mathscr{H}\theta+k^{2}\Psi,\qquad u=\frac{\theta}{k},\qquad\sigma=0.$
Initial conditions are adiabatic with $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$. The species is parametrized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today.
A.4 Baryons $(b)$
Baryons are also non-relativistic, but interact with photons through Compton scattering and are subject to recombination physics. This significantly complicates their behavior. SymBoltz currently implements equations from RECFAST [^12] https://www.astro.ubc.ca/people/scott/recfast.html version 1.5.2 (Seager et al., 1999, 2000; Wong et al., 2008; Scott & Moss, 2009):
$w=0,\qquad P=0,\qquad\rho=\frac{\rho_{0}}{a^{3}},\qquad{f_{\mathrm{He}}}=\frac{{Y_{\mathrm{He}}}}{\frac{m_{\mathrm{He}}}{m_{\mathrm{H}}}\big(1-{Y_{\mathrm{He}}}\big)},\qquad n_{\mathrm{H}}=\frac{(1-Y_{\mathrm{He}})\rho}{m_{\mathrm{H}}},\qquad n_{\mathrm{He}}=f_{\mathrm{He}}n_{\mathrm{H}},\qquad{c_{s}^{2}}=\frac{k_{B}}{\mu c^{2}}\bigg({T_{b}}-\frac{T_{b}^{\prime}}{3\mathscr{H}}\bigg),$
$\beta=\frac{1}{k_{B}T_{b}},\qquad T_{b}^{\prime}=-2\mathscr{H}{T_{b}}-\frac{a}{H_{0}}\frac{8}{3}\frac{T_{\gamma}^{4}{X_{\mathrm{e}}}}{1+{f_{\mathrm{He}}}+{X_{\mathrm{e}}}}\big(T_{b}-T_{\gamma}\big),\qquad{\mu}=\frac{m_{\mathrm{H}}}{1+\big(\frac{m_{\mathrm{H}}}{m_{\mathrm{He}}}-1\big){Y_{\mathrm{He}}}+\big(1-{Y_{\mathrm{He}}}\big){X_{\mathrm{e}}}},\qquad\kappa^{\prime}=-\frac{a}{H_{0}}n_{\mathrm{e}}\sigma_{T}c,$
${v}=-\kappa^{\prime}e^{-{\kappa}},\qquad n_{\mathrm{e}}=X_{\mathrm{e}}n_{\mathrm{H}},\qquad{X_{\mathrm{e}}}={X_{\mathrm{H}}^{+}}+{X_{\mathrm{He}}^{++}}+{f_{\mathrm{He}}}{X_{\mathrm{He}}^{+}}+{X_{\mathrm{e}}^{\mathrm{re}_{1}}}+{X_{\mathrm{e}}^{\mathrm{re}_{2}}},\qquad X_{\mathrm{H}}^{+\prime}=-\frac{a}{H_{0}}{C_{\mathrm{H}}}\Big({\alpha}_{\mathrm{H}}n_{\mathrm{e}}{X_{\mathrm{H}}^{+}}-{\beta_{\mathrm{H}}}e^{-{\beta_{b}}E_{\mathrm{H}}^{2s,1s}}\big(1-{X_{\mathrm{H}}^{+}}\big)\Big),$
${X_{\mathrm{He}_{1}}^{+\prime}}=-\frac{a}{H_{0}}C_{\mathrm{He}_{1}}\Big(\alpha_{\mathrm{He}_{1}}n_{\mathrm{e}}X_{\mathrm{He}}^{+}-\beta_{\mathrm{He}_{1}}e^{-\beta_{b}E_{\mathrm{He}_{1}}^{2s,1s}}\big(1-X_{\mathrm{He}}^{+}\big)\Big),\qquad X_{\mathrm{He}_{3}}^{+\prime}=-\frac{a}{H_{0}}C_{\mathrm{He}_{3}}\Big(n_{\mathrm{e}}{\alpha_{\mathrm{He}_{3}}}{X_{\mathrm{He}}^{+}}-3{{\beta}_{\mathrm{He}_{3}}}e^{-{\beta_{b}}E_{\mathrm{He}_{3}}^{2s,1s}}\big(1-{X_{\mathrm{He}}^{+}}\big)\Big),$
${X_{\mathrm{He}}^{+\prime}}={X_{\mathrm{He}_{1}}^{+\prime}}+{X_{\mathrm{He}_{3}}^{+\prime}},\qquad{X_{\mathrm{He}}^{++}}=\frac{2{f_{\mathrm{He}}}{R_{\mathrm{He}}^{+}}}{\bigg(1+f_{\mathrm{He}}+R_{\mathrm{He}}^{+}\bigg)\bigg(1+\sqrt{1+\frac{4{f_{\mathrm{He}}}{R_{\mathrm{He}}^{+}}}{(1+f_{\mathrm{He}}+R_{\mathrm{He}}^{+})^{2}}}\bigg)},\qquad R_{\mathrm{He}^{+}}=\frac{\exp\big({-\beta E_{\mathrm{He}^{+}}^{\infty,1s}}\big)}{n_{\mathrm{H}}\lambda_{\mathrm{e}}^{3}},\qquad\lambda_{\mathrm{e}}=\frac{h}{\sqrt{2\pi m_{\mathrm{e}}/\beta}},$
${X_{\mathrm{e}}^{\mathrm{re}_{1}}}=\frac{1+f_{\mathrm{He}}}{2}+\frac{1+f_{\mathrm{He}}}{2}\tanh\bigg(\frac{4}{3}\frac{(1+z^{\mathrm{re}_{1}})^{3/2}-(1+z)^{3/2}}{(1+z^{\mathrm{re}_{1}})^{1/2}}\bigg),\qquad{X_{\mathrm{e}}^{\mathrm{re}_{2}}}=\frac{f_{\mathrm{He}}}{2}+\frac{f_{\mathrm{He}}}{2}\tanh\bigg(\frac{4}{3}\frac{(1+{z^{\mathrm{re}_{2}}})^{3/2}-(1+z)^{3/2}}{(1+{z^{\mathrm{re}_{2}}})^{1/2}}\bigg),$
$\delta^{\prime}=-\theta-3\mathscr{H}c_{s}^{2}\delta+3\Phi^{\prime},\qquad\theta^{\prime}=-\mathscr{H}\theta+k^{2}{c_{s}^{2}}\delta+k^{2}\Psi-\frac{4}{3}\kappa^{\prime}\frac{\rho_{\gamma}}{\rho_{b}}\big(\theta_{\gamma}-\theta_{b}\big),\qquad u=\frac{\theta}{k},\qquad\sigma=0.$
Transition rates and coefficients related to recombination of Hydrogen include fitting functions that emulate the results of more accurate and expensive computations (here $\ln(a)$ is the logarithm of the scale factor, while $(Fa)$ is an unrelated fudge factor):
$\alpha_{\mathrm{H}}=10^{-19}(Fa)\frac{\Big(\frac{T_{b}}{T_{0}}\Big)^{b}}{1+c\Big(\frac{T_{b}}{T_{0}}\Big)^{d}},\qquad\beta_{\mathrm{H}}=\frac{\alpha_{\mathrm{H}}}{\lambda_{\mathrm{e}}^{3}}\exp\big({-\beta E_{\mathrm{H}}^{\infty,2s}}\big),$
$K_{\mathrm{H}}=\Bigg(1+A_{1}\exp\Bigg({-\bigg(\frac{\ln(a)-\ln(a_{1})}{w_{1}}\bigg)^{2}}\Bigg)+A_{2}\exp\Bigg({-\bigg(\frac{\ln(a)-\ln(a_{2})}{w_{2}}\bigg)^{2}}\Bigg)\Bigg)\frac{\big(\lambda_{\mathrm{H}}^{2s,1s}\big)^{3}}{8\pi H},\qquad C_{\mathrm{H}}=\frac{1+K_{\mathrm{H}}\Lambda_{\mathrm{H}}n_{\mathrm{H}}(1-X_{\mathrm{H}}^{+})}{1+K_{\mathrm{H}}(\Lambda_{\mathrm{H}}+\beta_{\mathrm{H}})n_{\mathrm{H}}(1-X_{\mathrm{H}}^{+})}.$
Helium rates and coefficients are even more complicated. First, Helium includes contributions from singlet states ($\mathrm{He}_{1}$}):
$\alpha_{\mathrm{He}_{1}}=\frac{q_{1}}{\sqrt{\frac{T_{b}}{T_{2}}}\Big(1+\sqrt{\frac{T_{b}}{T_{2}}}\Big)^{1-p_{1}}\Big(1+\sqrt{\frac{T_{b}}{T_{1}}}\Big)^{1+p_{1}}},\qquad\beta_{\mathrm{He}_{1}}=4\frac{\alpha_{\mathrm{He}_{1}}}{\lambda_{\mathrm{e}}^{3}}\exp\big({-βE_{\mathrm{He}_{1}}^{\infty,2s}}\big),\qquad K_{\mathrm{He}_{1}}=\frac{1}{K_{\mathrm{He}_{1}^{0}}^{-1}+K_{\mathrm{He}_{1}^{1}}^{-1}+K_{\mathrm{He}_{1}^{2}}^{-1}},$
$K_{\mathrm{He}_{1}^{0}}^{-1}=\frac{8\pi H}{\big(\lambda_{\mathrm{He}_{1}}^{2p,1s}\big)^{3}},\qquad K_{\mathrm{He}_{1}^{1}}^{-1}=-\exp(-\tau_{\mathrm{He}_{1}})K_{\mathrm{He}_{1}^{0}}^{-1},\qquad K_{\mathrm{He}_{1}^{2}}^{-1}=\frac{A_{2p_{1}}}{3\big(1+0.36\,\gamma_{2p_{1}}^{0.86}\big)n_{\mathrm{He}}(1-X_{\mathrm{He}}^{+})},\qquad\tau_{\mathrm{He}_{1}}=\frac{3A_{2p_{1}}n_{\mathrm{He}}(1-X_{\mathrm{He}}^{+})}{K_{\mathrm{He}_{1}^{0}}^{-1}},$
$\gamma_{2p_{1}}=\frac{3A_{2p_{1}}f_{\mathrm{He}}c^{2}(1-X_{\mathrm{He}}^{+})}{8\pi\sigma_{{\mathrm{He}_{1}}}\sqrt{\frac{2\pi}{\beta m_{\mathrm{He}}c^{2}}}\big(f_{\mathrm{He}_{1}}^{2p,1s}\big)^{3}(1-X_{\mathrm{H}}^{+})},\qquad C_{\mathrm{He}_{1}}=\frac{\exp\big({-\beta E_{\mathrm{He}_{1}}^{2p,2s}}\big)+K_{\mathrm{He}_{1}}\Lambda_{\mathrm{He}_{1}}n_{\mathrm{He}}(1-X_{\mathrm{He}}^{+})}{\exp\big({-\beta E_{\mathrm{He}_{1}}^{2p,2s}}\big)+K_{\mathrm{He}_{1}}(\Lambda_{\mathrm{He}_{1}}+\beta_{\mathrm{He}_{1}})n_{\mathrm{He}}(1-X_{\mathrm{He}}^{+})}.$
Second, Helium also includes contributions from triplet states ($\mathrm{He}_{3}$}):
$\alpha_{\mathrm{He}_{3}}=\frac{q_{3}}{\sqrt{\frac{T_{b}}{T_{2}}}\Big(1+\sqrt{\frac{T_{b}}{T_{2}}}\Big)^{1-p_{3}}\Big(1+\sqrt{\frac{T_{b}}{T_{1}}}\Big)^{1+p_{3}}},\qquad\beta_{\mathrm{He}_{3}}=\frac{4}{3}\frac{\alpha_{\mathrm{He}_{3}}}{\lambda_{\mathrm{e}}^{3}}\exp\big({-\beta E_{\mathrm{He}_{3}}^{\infty,2s}}\big),\qquad\tau_{\mathrm{He}_{3}}=\frac{3A_{2p_{3}}n_{\mathrm{He}}(1-X_{\mathrm{He}}^{+})\big(\lambda_{\mathrm{He}_{3}}^{2p,1s}\big)^{3}}{8\pi H},$
$\gamma_{2p_{3}}=\frac{3A_{2p_{3}}f_{\mathrm{He}}c^{2}(1-X_{\mathrm{He}}^{+})}{8\pi\sigma_{{\mathrm{He}_{3}}}\sqrt{\frac{2\pi}{\beta m_{\mathrm{He}}c^{2}}}\big(f_{\mathrm{He}_{3}}^{2p,1s}\big)^{3}(1-X_{\mathrm{H}}^{+})},\qquad C_{\mathrm{He}_{3}}=\frac{A_{2p_{3}}\bigg(\frac{1-\exp({-\tau_{\mathrm{He}_{3}}})}{\tau_{\mathrm{He}_{3}}}+\frac{1}{3\big(1+0.66\,\gamma_{2p_{3}}^{0.9}\big)}\bigg)\exp\big({-\beta E_{\mathrm{He}_{3}}^{2p,2s}}\big)}{A_{2p_{3}}\bigg(\frac{1-\exp({-\tau_{\mathrm{He}_{3}}})}{\tau_{\mathrm{He}_{3}}}+\frac{1}{3\big(1+0.66\,\gamma_{2p_{3}}^{0.9}\big)}\bigg)\exp\big({-\beta E_{\mathrm{He}_{3}}^{2p,2s}}\big)+\beta_{\mathrm{He}_{3}}}.$
Every variable that does not occur on the left side of an equation is either a constant or a parameter. This includes $Y_{\mathrm{He}}$, fudge factors and wavenumbers, frequencies and energies for atomic transitions. Some important variables are the baryon temperature $T_{b}$, photon temperature $T_{\gamma}$, mean molecular weight $\mu$, baryon sound speed $c_{s}^{2}$, optical depth $\kappa$, visibility function $v$ and the free electron fraction $X_{\mathrm{e}}$ (conventionally relative to Hydrogen, so $X_{\mathrm{e}}>1$ in presence of Helium). Please consult the code and RECFAST references cited above for more details.
Unlike other RECFAST implementations, SymBoltz does not approximate the stiff Peebles equations at early times by Saha approximations (although $X_{\mathrm{He}}^{++}$ is given by a Saha equation at all times). This is not necessary with a good implicit ODE solver. SymBoltz sets $C_{\mathrm{H}}=C_{\mathrm{He}_{1}}=1$ when $X_{\mathrm{e}}\gtrsim 0.99$ to avoid numerical instability at early times. Atomic calculations are done in SI units and converted to SymBoltz’ dimensionless units by factors of $H_{0}$ in SI units. The differential equation for $T_{b}^{\prime}$ is very stiff and sensitive to $T_{b}-T_{\gamma}$, but $T_{b}\approx T_{\gamma}$ in the early universe, so we rewrite it to a more stable differential equation for $\Delta T^{\prime}=T_{b}^{\prime}-T_{\gamma}^{\prime}$ instead, initialize $\Delta T=0$ and observe $T_{b}=\Delta T+T_{\gamma}$. The optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\tau^{\prime})\mathrm{d}\tau^{\prime}$ is really a line-of-sine integral into the past, but is integrated together with the background ODEs by initializing $\kappa(\tau_{i})=0$ to an arbitrary value, integrating the differential equation for $\kappa^{\prime}$ and subtracting the final value of $\kappa(\tau_{0})$ (i.e. $\int_{\tau_{0}}^{\tau}=\int_{\tau_{0}}^{\tau_{i}}+\int_{\tau_{i}}^{\tau}=\int_{\tau_{i}}^{\tau}-\int_{\tau_{i}}^{\tau_{0}}$). There is no tight-coupling approximation.
Initial conditions are full ionization $X_{\mathrm{H}}^{+}=X_{\mathrm{He}}^{+}=1$, thermal equilibrium $T_{b}=T_{\gamma}$ ($\Delta T=0$), the arbitrary $\kappa=0$, and adiabatic perturbations $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$. The baryon species is parametrized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today and the primordial Helium mass fraction $Y_{\mathrm{He}}$.
SymBoltz solves thermodynamics equations together with the background equations, while some other codes treat these as separate stages. There is no meaningful performance improvement from doing this, as the size of the background (and thermodynamics) ODEs is so small. This makes a clear distinction between the background with all 0th order equations of motion, and the perturbations with all 1st order equations. It also makes it easy to create exotic models where the thermodynamics couple to the background, for example.
Note that RECFAST uses fitting functions to emulate the results of more physically accurate and expensive simulations. These are tuned to work for the $\mathrm{\Lambda CDM}$ model. SymBoltz would therefore benefit from including more physically accurate recombination models for safer use with modified models.
A.5 Photons $(\gamma)$
Photons are massless and therefore ultra-relativistic. Unlike non-relativistic particles, one must account for the direction $\cos\theta=\boldsymbol{p}\cdot\boldsymbol{k}/\lvert\boldsymbol{p}\rvert\lvert\boldsymbol{k}\rvert$ of their momenta $\boldsymbol{p}$ relative to the Fourier wavenumber $\boldsymbol{k}$. This results in a theoretically infinite hierarchy of equations for Legendre multipoles $l$, which in practice must be truncated at some maximum multipole $l_{\text{max}}$:
$T=\frac{{T_{0}}}{a},\qquad w=\frac{1}{3},\qquad{c_{s}^{2}}=w,\qquad P=\frac{\rho}{3},\qquad\rho=\frac{\rho_{0}}{a^{4}},\qquad\delta=F_{0},\qquad\theta=\frac{3}{4}kF_{1},\qquad u=\frac{\theta}{k},\qquad\sigma=\frac{F_{2}}{2},$
$F_{0}^{\prime}=-kF_{1}+4\Phi^{\prime},\qquad F_{1}^{\prime}=\frac{k}{3}\big(F_{0}-2F_{2}+4\Psi\big)+\frac{4}{3}\frac{\kappa^{\prime}}{k}\big(\theta_{\gamma}-{\theta_{b}}\big),$
$F_{l}^{\prime}=\frac{k}{2l+1}\big(lF_{l-1}-(l+1)F_{l+1}\big)+F_{l}{\kappa^{\prime}}-\delta_{l,2}\frac{\kappa^{\prime}}{10}\Pi,\qquad F_{l_{\text{max}}}^{\prime}=kF_{l_{\text{max}}-1}-\frac{l_{\text{max}}+1}{\tau}F_{l_{\text{max}}}+{\kappa^{\prime}}F_{l_{\text{max}}},$
$G_{0}^{\prime}=-kG_{1}+{\kappa^{\prime}}{G_{0}}-\frac{\kappa^{\prime}}{2}\Pi,\qquad G_{l}^{\prime}=\frac{k}{2l+1}\big(lG_{l-1}-(l+1)G_{l+1}\big)+\kappa^{\prime}G_{l}-\delta_{l,2}\frac{{\kappa^{\prime}}}{10}\Pi,$
$G_{l_{\text{max}}}^{\prime}=kG_{l_{\text{max}}-1}-\frac{l_{\text{max}}+1}{\tau}G_{l_{\text{max}}}+{\kappa^{\prime}}G_{l_{\text{max}}},\qquad\Pi=F_{2}+G_{0}+G_{2},\qquad{{\Theta}_{l}}=\frac{F_{l}}{4}.$
The equations for $F_{l}^{\prime}$ and $G_{l}^{\prime}$ apply for $2\leq l<l_{\text{max}}$. There are no tight-coupling, radiation-streaming or ultra-relativistic fluid approximations. Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e. $\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e. $\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=-\frac{8}{15}\frac{k}{\kappa^{\prime}}F_{1}$, $G_{0}=\frac{5}{16}F_{2}$, $G_{1}=-\frac{1}{16}\frac{k}{\kappa^{\prime}}F_{2}$, $G_{2}=\frac{1}{16}F_{2}$, and $F_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}F_{l-1}$ and $G_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}G_{l-1}$ for $3\leq l\leq l_{\text{max}}$. The species is parametrized by its temperature $T_{0}$ today, which in turn sets the density parameters $\Omega_{0}=\frac{\pi^{2}}{15}\frac{(k_{B}T_{0})^{4}}{(\hbar c)^{3}}\frac{8\pi G}{3H_{0}^{2}}$ and $\rho_{0}=\frac{8\pi}{3}\Omega_{0}$ today.
A.6 Massless neutrinos $(\nu)$
Massless neutrinos behave similarly to photons, but decouple from interactions with other species in the very early universe. One must only account for this interaction in initial conditions, while their evolution equations are a simpler case of the photons’:
$T=\frac{T_{0}}{a},\qquad w=\frac{1}{3},\qquad{c_{s}^{2}}=\frac{1}{3},\qquad P=\frac{\rho}{3},\qquad\rho=\frac{\rho_{0}}{a^{4}},\qquad\delta={F_{0}},\qquad\theta=\frac{3}{4}kF_{1},\qquad\sigma=\frac{F_{2}}{2},$
$F_{0}^{\prime}=-kF_{1}+4\Phi^{\prime},\qquad F_{1}^{\prime}=\frac{k}{3}\big({F_{0}}-2F_{2}+4\Psi\big),\qquad F_{l}^{\prime}=\frac{k}{2l+1}\big(lF_{l-1}-(l+1)F_{l+1}\big),\qquad F_{l_{\text{max}}}^{\prime}=kF_{l_{\text{max}}-1}-\frac{l_{\text{max}}+1}{\tau}F_{l_{\text{max}}}.$
The equations for $F_{l}^{\prime}$ apply for $2\leq l<l_{\text{max}}$. There is no ultra-relativistic fluid approximation. Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e. $\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e. $\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=\frac{2}{15}(k\tau)^{2}\Psi$ and $F_{l}=\frac{l}{2l+1}k\tau F_{l-1}$. The species is parametrized by the effective number $N_{\text{eff}}$, the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ and temperature $T_{0}$ today. If photons are present, they default to $T_{\nu 0}=\big(\frac{4}{11}\big)^{1/3}T_{\gamma 0}$ and $\Omega_{\nu 0}=N_{\text{eff}}\frac{7}{8}\big(\frac{4}{11}\big)^{4/3}\Omega_{\gamma 0}$.
A.7 Massive neutrinos $(h)$
Massive neutrinos are the most complicated species in the $\mathrm{\Lambda CDM}$ model (alongside baryon recombination). In essence, the species we have looked at so far have Boltzmann equations where the momenta of their distribution function can be integrated out analytically in non-relativistic and ultra-relativistic limits. This means that their stress-energy components are linked by trivial equations of state and sound speeds, for example, and their effect can be parametrized by a simple density parameter $\Omega_{0}$.
On the other hand, massive neutrinos have intermediate masses that fall between the non-relativistic and ultra-relativistic limits. Integrals over their distribution function must be computed numerically. This is very expensive if done naively, and it is extremely important to choose a quadrature scheme that minimizes the number of sampled points. Fortunately, the momentum integrals have a structure that can be exploited: they are all in the form weighted form $I[g(x)]=\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)$, where $f(x)=1/(e^{x}+1)$ is the equilibrium distribution function and $g(x)$ is an arbitrary function of the dimensionless momentum $x=pc/k_{B}T$ (see Ma & Bertschinger (1995) for more details). One can generally approximate $I[g(x)]\approx\sum_{i}W_{i}\,g(x_{i})$ with a weighted quadrature scheme with points $x_{i}$ and weights $W_{i}$ (more on this after the equations). In other words, the integral operator $\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)$ is effectively replaced by the discrete summation operator $\sum_{i}\!W_{i}$ for some weights $W_{i}$.
On top of this, perturbations are also expanded in Legendre multipoles $l$ up to a cutoff $l_{\text{max}}$:
$T=\frac{{T_{0}}}{a},\qquad x=\frac{pc}{k_{B}T},\qquad y=\frac{mc^{2}}{k_{B}T},\qquad E_{i}=\sqrt{x_{i}^{2}+y^{2}},\qquad f=\frac{1}{1+e^{x}},\qquad\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}=-\frac{x}{1+e^{-x}},$
$I_{\rho}=\sum_{i}\!W_{i}E_{i},\qquad I_{P}=\sum_{i}\!W_{i}\frac{x_{i}^{2}}{E_{i}},\qquad\rho=\frac{N}{\pi^{2}}\frac{(k_{B}T)^{4}}{(\hbar c)^{3}}\frac{G}{(H_{0}c)^{2}}I_{\rho},\qquad P=\frac{N}{3\pi^{2}}\frac{(k_{B}T)^{4}}{(\hbar c)^{3}}\frac{G}{(H_{0}c)^{2}}I_{P},\qquad w=\frac{P}{\rho},$
$\psi_{i,0}^{\prime}=-k\frac{x_{i}}{E_{i}}\psi_{i,1}-\Phi^{\prime}\left(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\right)_{i},\qquad\psi_{i,1}^{\prime}=\frac{k}{3}\frac{x_{i}}{E_{i}}\big(\psi_{i,0}-2\psi_{i,2}\big)-\frac{k}{3}\frac{E_{i}}{x_{i}}\Psi\left(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\right)_{i},$
$\psi_{i,l}^{\prime}=\frac{k}{2l+1}\frac{x_{i}}{E_{i}}\big(l\psi_{i,l-1}-(l+1)\psi_{i,l+1}\big),\qquad\psi_{i,l_{\text{max}}+1}=\frac{2l_{\text{max}}+1}{k\tau}\frac{E_{i}}{x_{i}}\psi_{i,l_{\text{max}}}-\psi_{i,l_{\text{max}}-1},$
$I_{0}=\sum_{i}\!W_{i}E_{i}\psi_{i,0},\qquad I_{1}=\sum_{i}\!W_{i}x_{i}\psi_{i,1},\qquad I_{2}=\sum_{i}\!W_{i}\frac{x_{i}^{2}}{E_{i}}\psi_{i,2},\qquad\delta=\frac{I_{0}}{I_{\rho}},\qquad\sigma=\frac{2I_{2}}{3I_{\rho}+I_{P}},\qquad u=\frac{3I_{1}}{3I_{\rho}+I_{P}},\qquad\theta=ku.$
Initial conditions are $\psi_{i,0}=-\frac{1}{4}(-2\Psi)\big(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\big)_{i}$, $\psi_{i,1}=-\frac{1}{3}\frac{E_{i}}{x_{i}}\frac{1}{2}k\tau\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\big)_{i}$, $\psi_{i,2}=-\frac{1}{2}\frac{1}{15}(k\tau)^{2}\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\big)_{i}$ and $\psi_{i,l}=0$. This integrates to adiabatic $\delta/(1+w)$, $\theta$ and $\sigma$ similarly to massless neutrinos. Free parameters are the temperature today $T_{0}$, mass $m$ of a single neutrino and degeneracy factor $N=\sum_{i=1}^{N}m_{i}/m$ for describing multiple neutrinos with equal mass. The degeneracy factor defaults to $N=3$, and the temperature to $T_{h0}=\big(\frac{4}{11}\big)^{1/3}T_{\gamma 0}$ if photons are present, as for massless neutrinos.
Here the equation for $\psi_{i,l}^{\prime}$ applies for $2\leq l\leq l_{\text{max}}$, and expressions $g_{i}=g(x_{i})$ indexed by $i$ are evaluated with the momentum quadrature point $x=x_{i}$. The reduction to dimensionless momenta $x=pc/k_{B}T$ (the argument of $\exp$ in $f$) is deliberate because it makes numerics more well-defined and the quadrature scheme independent of $m$ and all other cosmological parameters.
SymBoltz automatically computes momentum bins $x_{i}$ and quadrature weights $W_{i}$ with $N$-point Gaussian quadrature. First, by default, the following substitution is applied to the momentum integral:
$\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)=\int_{u(0)}^{u(\infty)}\mathrm{d}u\,x^{\prime}(u)x(u)^{2}f(x(u))g(x(u))\quad\text{with}\quad u(x)=\frac{1}{1+\frac{x}{L}}.$
This substitution achieves two things: the scaling $x/L$ brings the dominant integral contributions well within $x/L\ll 1$ if $L$ is chosen to be a characteristic decay ?length? of the distribution function, and the rational part $1/(1+x/L)$ maps the infinite domain $x\in(0,\infty)$ to the finite domain $u\in(0,1)$, which can be integrated numerically. The substituted integrand is then passed to an adaptive algorithm in QuadGK.jl [^13] https://github.com/JuliaMath/QuadGK.jl that computes quadrature points $u_{i}$ and weights $W_{i}$ by performing weighted integrals against several test functions $g(x)$. Finally, the corresponding momenta $x_{i}=x(u_{i})$ are returned along with the weights $W_{i}$, from which one can approximate the integral $I[g(x)]\approx\sum_{i}W_{i}g(x_{i})$ against any $g(x)$.
SymBoltz tests this numerical quadrature scheme against the analytical result $I[x^{n-2}]=\int_{0}^{\infty}\mathrm{d}x\,x^{n}/(e^{x}+1)=(1-2^{-n})\zeta(n+1)\Gamma(n+1)$ for $2\leq n\leq 8$. We assume this to be a reasonable test for the integrals encountered in the equations above. Agreement is excellent with $L=100$, which yields relative errors below $10^{-6+n-N}$ for all $2\leq n\leq 8$ and $1\leq N\leq 5$. SymBoltz defaults to $N=4$ momenta, for which this relative error is less than $10^{-6}$ for $n\leq 4$, for example. It also agrees well with CLASS using default settings.
Note that this momentum quadrature strategy is generic with respect to the distribution function $f(x)$ and substitution $u(x)$, so it can easily be modified for other particle species whose distribution function cannot be integrated out.
CAMB (Lewis, 2025) and CLASS (Lesgourgues & Tram, 2011) apply similar weighted quadrature strategies. They also get away with only a handful of sampled momenta, but the precise details of the computation differ slightly. For reference, here are points and weights computed by SymBoltz for $1\leq N\leq 8$ momenta:
NN $x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ $x_{5}$ $x_{6}$ $x_{7}$ $x_{8}$
1 3.12273
2 2.07807 5.94834
3 1.56110 4.22902 8.86258
4 1.24461 3.30909 6.57536 11.80351
5 1.02955 2.71805 5.27853 9.03363 14.74043
6 0.87373 2.30142 4.41479 7.39595 11.55088 17.65818
7 0.75572 1.99028 3.79064 6.27323 9.60568 14.09716 20.54878
8 0.66337 1.74848 3.31580 5.44442 8.24194 11.87257 16.65452 23.40806
NN $W_{1}$ $W_{2}$ $W_{3}$ $W_{4}$ $W_{5}$ $W_{6}$ $W_{7}$ $W_{8}$
1 1.80309
2 1.30306 0.50002
3 0.84813 0.88596 0.06899
4 0.55272 0.99943 0.24384 0.00709
5 0.36868 0.95311 0.43658 0.04409 0.00063
6 0.25275 0.84165 0.58496 0.11736 0.00632 0.00005
7 0.17792 0.71541 0.67227 0.21284 0.02386 0.00079 0.00000
8 0.12832 0.59663 0.70569 0.31144 0.05685 0.00406 0.00009 0.00000
A.8 Cosmological constant $(\Lambda)$
The cosmological constant is equivalent to a very simple species without perturbations:
$w=-1,\qquad\rho=\rho_{0},\qquad P=-\rho,\qquad\delta=0,\qquad\theta=0,\qquad\sigma=0,\qquad u=0.$
It is parametrized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today. This is set to $\Omega_{\Lambda 0}=1-\sum_{s\neq\Lambda}\Omega_{s0}$ if all species $s$ have a $\Omega_{s0}$ parameter and General Relativity is the theory of gravity. This constraint comes from the 1st Friedmann equation today.
A.9 Primordial power spectrum ($I$)
SymBoltz computes the inflationary primordial power spectrum parametrized by the amplitude $A_{s}$ and tilt $n_{s}$:
$P_{0}(k)=\frac{2\pi^{2}}{k^{3}}A_{s}\bigg(\frac{k}{k_{\text{p}}}\bigg)^{n_{s}-1}.$
A.10 Matter power spectrum
SymBoltz computes the matter power spectrum for some desired set of species $s$, which are presumably matter-like at late times (e.g. $s\in\{c,b,h\}$):
$P(k,\tau)=P_{0}(k)\big|\Delta(\tau,k)\big|^{2}\quad\text{with}\quad\Delta=\delta+\frac{3\mathscr{H}}{k^{2}}\theta=\frac{\sum_{s}\delta\rho_{s}}{\sum_{s}\rho_{s}}+\frac{3\mathscr{H}}{k^{2}}\frac{\sum_{s}(\rho_{s}+P_{s})\theta_{s}}{\sum_{s}(\rho_{s}+P_{s})}.$
Here $\Delta$ is the total gauge-independent overdensity with total $\delta$ and $\theta$ computed by summing the components of the energy-momentum tensor that are additive.
A.11 CMB power spectrum and line-of-sight integration
SymBoltz finds photon temperature and polarization multipoles today for any $l$ by computing the line-of-sight integrals
$\Theta_{l}^{\mathrm{T}}\big(\tau_{0},k\big)=\int_{\tau_{i}}^{\tau_{0}}S_{T}(\tau,k)j_{l}\big(k(\tau_{0}-\tau)\big)\mathrm{d}\tau\quad\text{with}\quad S_{T}=v\bigg(\frac{\delta_{\gamma}}{4}+\Psi+\frac{\Pi_{\gamma}}{16}\bigg)+e^{-\kappa}\big(\Psi+\Phi\big)^{\prime}+\frac{\big(vu_{b}\big)^{\prime}}{k}+\frac{3}{16k^{2}}\big(v\Pi_{\gamma}\big)^{\prime\prime},$
$\Theta_{l}^{\mathrm{E}}\big(\tau_{0},k\big)=\sqrt{\frac{(l+2)!}{(l-2)!}}\int_{\tau_{i}}^{\tau_{0}}S_{E}\big(\tau,k\big)\frac{j_{l}\big(k(\tau_{0}-\tau)\big)}{\big(k(\tau_{0}-\tau)\big)^{2}}\mathrm{d}\tau\quad\text{with}\quad S_{E}=\frac{3}{16}v\Pi_{\gamma}.$
As first suggested by Seljak & Zaldarriaga (1996), this approach enables cheap computation for any $l$ after integrating the perturbation ODEs with only a few $l\leq l_{\text{max}}$. This drastically speeds up the computation over including all $l$ in an enormous set of coupled perturbation ODEs. SymBoltz performs the integrals with the trapezoid method using the substitution $u(\tau)=\tanh(\tau)$, which adds more points in the early universe when sampled uniformly, using $768$ points by default. Here $j_{l}$ are the spherical Bessel functions of the first kind. SymBoltz is not yet generalized to non-flat geometries, where they are replaced by hyperspherical functions. The cross-correlated angular spectrum between $\mathrm{A},\mathrm{B}\in\{\mathrm{T},\mathrm{E}\}$ is then computed from
$C_{l}^{\mathrm{AB}}=\frac{2\pi}{l(l+1)}D_{l}^{\mathrm{AB}}=\frac{2}{\pi}\int_{0}^{\infty}\mathrm{d}kk^{2}P_{0}(k)\,\Theta_{l}^{A}(\tau_{0},k)\,\Theta_{l}^{B}(\tau_{0},k).$
This integral is also performed with the trapezoid method. The point $(k,\Theta)=(0,0)$ is included manually, for which the numerical solution to the perturbation ODEs is ill-defined. By default, the $\Theta_{l}$ are sampled on a fine grid of wavenumbers with spacing $\Delta k=2\pi/2\tau_{0}$, which interpolates from solved perturbation modes on a coarse grid $\Delta k=8/\tau_{0}$. Both grids range between $0.1l_{\text{min}}/\tau_{0}\leq k\leq 3l_{\text{max}}/\tau_{0}$, where $l_{\text{min}}$ and $l_{\text{max}}$ are the angular spectrum’s minimum and maximum requested multipoles.
Appendix B Precision parameters for CLASS
When comparing results to CLASS in section 3, CLASS is configured with the following non-default precision parameters: {codebox}
background_Nloga = 6000tight_coupling_trigger_tau_c_over_tau_h = 1e-2tight_coupling_trigger_tau_c_over_tau_k = 1e-3radiation_streaming_approximation = 3ur_fluid_approximation = 3ncdm_fluid_approximation = 3We also set l_max_g, l_max_pol_g, l_max_ur, l_max_ncdm to the same $l_{\text{max}}$ used by SymBoltz’ model. These settings disable as many approximations as possible and reduces the impact of the tight-coupling approximation, which cannot be disabled. Oddly, we find that the parameter background_Nloga must be decreased from the default value 40000 to make the derivatives in fig. 4 stable. This parameter controls the number of points used for splining background functions in the perturbations. The default value of this parameter was changed from 3000 to 40000 in 2023, but we suspect that the increased density in points makes the splines susceptible to oscillations from numerical noise. These settings are important for good agreement between the Fisher forecasts in section 3.4. We used CLASS version 3.3.1.
Appendix C Testing and comparison to CLASS
SymBoltz’ code repository is set up with continuous integration that runs several tests and builds updated documentation pages every time changes to the code are committed. In particular, this compares the solution for $\mathrm{\Lambda CDM}$ with CLASS for many variables solved by the background, thermodynamics and perturbations (using the options write_background, write_thermodynamics and k_output_values). These are the basis for all derived quantities like luminosity distances, matter and CMB power spectra, which are also compared. The checks pass when the quantities agree within a small tolerance. We do not compare directly against more codes like CAMB, but CLASS has already been compared extensively with CAMB with excellent agreement (Lesgourgues, 2011b). The comparison takes a lot of space and is not included here, but is found in the documentation linked from SymBoltz’ repository.
Another test checks that integration of the background and perturbations equations are stable throughout parameter space. As the equations are very stiff and SymBoltz does not rely on approximations for relieving it, one could imagine that the integration would be stable for some parameter values and unstable for others. The test creates a box in parameter space $\pm 50\%$ around a fiducial set of realistic parameter values, draws several sets of parameter values from that space with Latin hypercube sampling (to efficiently cover parameter space) and integrates the background and perturbations for each such set. All parameter samples are found to integrate successfully without warnings and errors.

### v2

@@
 It features a symbolic-numeric interface for specifying equations, is free of approximation switching schemes and is compatible with automatic differentiation.
-Cosmological models are built from replaceable physical components in a way that scales well in model space.
-The modeler should simply write down their equations, and SymBoltz solves them and eliminates much of the friction in the process.
-SymBoltz enables up to $100\times$ shorter model definitions compared to browsing equivalent files in CLASS.
-Symbolic knowledge enables powerful automation of tasks, such as separating computational stages like the background and perturbations, generating the Jacobian matrix and its sparsity pattern, and interpolating arbitrary expressions from the solution.
-Modern implicit solvers integrate the full stiff equations at all times, reducing slowdowns by taking long time steps, reusing the Jacobian and LU-factorizing it over several time steps, and using fast linear system solvers.
+Cosmological models are built from replaceable physical components in a way that scales well in model space, or alternatively written as one compact system of equations.
+The modeler should simply write down their equations, and SymBoltz solves them and eliminates friction in the modeling process.
+Symbolic knowledge enables powerful automation of tasks, such as separating computational stages like the background and perturbations, generating the analytical Jacobian matrix and its sparsity pattern, and interpolating arbitrary variables from the solution.
+Implicit solvers integrate the full stiff equations at all times without approximations, which greatly simplifies the code.
+Performance remains as good as in existing approximation-based codes due to high-order implicit methods that take long time steps, fast generated code, optimal handling of the Jacobian and efficient sparse matrix methods.
 Automatic differentiation gives exact derivatives of any output with respect to any input, which is important for gradient-based Markov chain Monte Carlo methods in large parameter spaces, training of emulators, Fisher forecasting and sensitivity analysis.
-These features are useful in their own rights, but also reinforce each other in a synergy.
-Results agree with established codes like CLASS and CAMB.
-With more work, SymBoltz can grow into an integrated symbolic-numeric cosmological modeling environment with a large library of models that delivers differentiable output as fast as other codes.
-SymBoltz is available at https://github.com/hersle/SymBoltz.jl with single-command installation and extensive documentation, and welcomes questions, suggestions and contributions.
+The main features form a synergy that reinforces the design of the code.
+Results agree with established codes to $0.1\%$ with standard precision.
+More work is needed to implement additional features and for fast reverse-mode automatic differentiation of scalar loss functions.
+SymBoltz is available at https://github.com/hersle/SymBoltz.jl with single-command installation and extensive documentation, and welcomes all contributions.
 Key Words.:
@@
 Efficiently exploring the space of models benefits from numerical tools that are easy to modify.
-Minimizing friction in the modeling process encourages relaxing model-dependent approximations, creating user-friendly interfaces and structuring codes in modular components that are easy to replace.
+Minimizing friction in the modeling process encourages relaxing model-dependent approximations, creating user-friendly interfaces and structuring codes with modular components that are easy to replace.
 Some of the most important such tools in the cosmological modeling toolbox are linear Einstein-Boltzmann solvers ( ?Boltzmann codes?) like CAMB (Lewis et al., 2000) and CLASS (Lesgourgues, 2011a).
 On the observational frontier, next-generation surveys like the Square Kilometer Array (Dewdney et al., 2009), Vera C.
-Rubin Observatory (LSST Science Collaboration, 2009), Dark Energy Spectroscopic Instrument (DESI Collaboration, 2016), Simon’s Observatory (The Simons Observatory Collaboration, 2019) and Euclid (Euclid Collaboration, 2025) promise even more precise data.
-Setting models apart with upcoming data involves both theoretical model parameters and experimental nuisance parameters that live side-by-side in large $O(100)$-dimensional spaces (Piras et al., 2024).
-In high dimensions, modern Markov chain Monte Carlo methods like Hamiltonian Monte Carlo and the No-U-Turn Sampler (Hoffman & Gelman, 2011) beat the traditional Metropolis-Hastings algorithm (Hastings, 1970).
-These methods explore parameter space more efficiently, but need both the likelihood and its derivatives with respect to parameters.
-Differentiability is also increasingly important in other applications, such as forward-modeling field-level inference using simulations with initial conditions from Boltzmann solvers (Seljak et al., 2017).
-Automatic differentiation is a way to compute derivatives that is more accurate and can be faster than finite differences, which is an approximate and brute-force method.
-A differentiable Boltzmann solver is therefore a cornerstone in the next-generation cosmological modeling toolbox.
-SymBoltz is a new Boltzmann solver that aims to fill these gaps, featuring a convenient symbolic-numeric interface, approximation-freeness and differentiability.
-At its core, a Boltzmann code solves the gravitational (Einstein) equations for a gravitational theory coupled to some particle species described by thermodynamic (Boltzmann) equations up to first perturbative order around a homogeneous and isotropic universe.
-It can predict cosmic microwave background (CMB), baryon acoustic oscillation (BAO) or supernova (SN) observations; or be used in longer computations, such as to generate initial conditions to non-linear $N$-body simulations of large-scale structure.
+Rubin Observatory (LSST Science Collaboration, 2009), Dark Energy Spectroscopic Instrument (DESI Collaboration, 2016), Simon’s Observatory (The Simons Observatory Collaboration, 2019) and Euclid (Euclid Collaboration, 2025) promise more precise data.
+Setting models apart with upcoming data involves both theoretical model parameters and experimental nuisance parameters that coexist in large $O(100)$-dimensional spaces (Piras et al., 2024).
+In high dimensions, modern Markov chain Monte Carlo methods like Hamiltonian Monte Carlo and the No-U-Turn Sampler (Hoffman and Gelman, 2011) outperform the traditional Metropolis-Hastings algorithm (Hastings, 1970).
+They explore parameter space more efficiently, but need both the likelihood and its gradient with respect to parameters.
+Emulators have become popular to get this from existing non-differentiable codes by training differentiable neural networks to reproduce their output (e.g Bolliet et al., 2024; Bonici et al., 2024).
+Differentiability is also used in forward-modeling field-level inference with simulations and initial conditions from Boltzmann solvers, for example (Seljak et al., 2017).
+Automatic differentiation can compute derivatives more accurately and faster than approximate and brute-force finite differences.
+A differentiable Boltzmann solver is a cornerstone in the modern cosmological modeling toolbox.
+SymBoltz is a new Boltzmann solver that aims to meet these needs, featuring a convenient symbolic-numeric interface, approximation-freeness and differentiability.
+At its core, a Boltzmann code solves the Einstein equations for a gravitational theory coupled to some particle species described by Boltzmann equations up to first perturbative order around a homogeneous and isotropic universe.
+For example, it predicts cosmic microwave background (CMB), baryon acoustic oscillation (BAO) or supernova (SN) observations, or generates initial conditions for non-linear $N$-body simulations of large-scale structure.
 This article is structured as follows.
-Section 1 revisits the historical development of Boltzmann codes, sketches their structure, reviews methods for computing derivatives and motivates SymBoltz.
-Section 2 maps out the architecture and main features of SymBoltz and how it differs from other codes.
-Section 3 shows some example usage.
-Section 4 discusses synergies and tradeoffs in the design.
-Section 5 concludes with the current state of the code and future potential.
-The appendices list equations, implementation details, comparison settings and testing methods.
+Section 1 motivates SymBoltz by reviewing the historical development of Boltzmann codes and methods for computing derivatives.
+Section 2 describes the structure and main features of SymBoltz and how it differs from other codes.
+Section 3 shows example usage.
+Section 4 compares performance to an existing code.
+Section 5 describes possibilities for future work.
+Section 6 concludes with the current state of the code.
+Appendices A, B and C list equations, implementation details, comparison settings and testing methods.
 1 History and motivation
-Peebles & Yu (1970) were first to numerically integrate a comprehensive set of linearized Einstein-Boltzmann equations.
-Their work was refined over several years, and Ma & Bertschinger (1995) established the groundwork for modern Boltzmann solvers with their code COSMICS.
-Seljak & Zaldarriaga (1996) soon realized that one can integrate photon multipoles by parts to reduce many differential equations to one integral solution and released CMBFAST.
+Peebles and Yu (1970) were first to numerically integrate a comprehensive set of linear Einstein-Boltzmann equations.
+Their work was refined over several years, and Ma and Bertschinger (1995) established the groundwork for modern Boltzmann solvers with their code COSMICS.
+Seljak and Zaldarriaga (1996) soon realized that one can integrate photon multipoles by parts to reduce many differential equations to one integral solution and released CMBFAST.
 Their line-of-sight integration method has become a standard technique that greatly speeds up the calculation, but requires truncating the multipoles in the differential equations.
@@
 Today they are very well-made, efficient and reliable tools.
-More recently, the market has seen an influx of alternative solvers that target new numerical techniques, GPU parallelization, differentiability, interactivity and symbolic computation.
-Refregier et al.
-(2017) published the first symbolic-numeric Boltzmann code PyCosmo [^3] https://pypi.org/project/PyCosmo/, which automatically generates efficient C++ ODE code from user-provided symbolic equations, performs sparsity optimization on its Jacobian matrix and avoids the use of approximation schemes to speed up and stabilize the integration.
-Hahn et al.
-(2024) published the first differentiable Boltzmann solver DISCO-EB [^4] https://github.com/ohahn/DISCO-EB in the JAX framework in Python, and relaxes approximation schemes to avoid complications and overhead when switching ODEs on GPUs.
+Recently, several alternative solvers have appeared on the market.
+They target new numerical techniques, lack of approximations, differentiability, GPU parallelization, interactivity and symbolic computation.
+PyCosmo [^3] https://pypi.org/project/PyCosmo/ by Refregier et al.
+(2018) in Python was the first symbolic-numeric Boltzmann code that generates fast C++ code from user-provided symbolic equations, optimizes the sparse and analytical Jacobian matrix and avoids approximation schemes.
 Bolt.jl [^4] https://github.com/xzackli/Bolt.jl by Li et al.
-(2023) accomplishes much of the same in Julia.
-SymBoltz.jl [^6] https://github.com/hersle/SymBoltz.jl is inspired by some of these developments.
+(2023) in Julia is also approximation-free, supports forward-mode automatic differentiation and uses similar differential equations solvers as SymBoltz.
+DISCO-EB [^5] https://github.com/ohahn/DISCO-EB by Hahn et al.
+(2024) in the JAX framework in Python is also differentiable and relaxes approximation schemes to avoid overhead from switching equations on GPUs.
+SymBoltz.jl [^6] https://github.com/hersle/SymBoltz.jl is inspired by all these developments.
 1.1 Structure of traditional Boltzmann solvers
 The full Einstein-Boltzmann equations are partial differential equations that linearize to ordinary differential equations.
-In principle, the core task of a Boltzmann solver is thus very simple: to solve these ordinary differential equations (ODEs) ${\mathrm{d}\mathbf{u}}/{\mathrm{d}\tau}=\mathbf{f}(\tau,\mathbf{u};\mathbf{p})$ for some initial conditions $\mathbf{u}(\tau_{0})$ and parameters $\mathbf{p}$, including several perturbation wavenumbers, and compute desired quantities from their solution, like power spectra.
+In principle, the core task of a Boltzmann solver is thus very simple: to solve these ordinary differential equations (ODEs) ${\mathrm{d}\boldsymbol{u}}/{\mathrm{d}\tau}=\boldsymbol{f}(\tau,\boldsymbol{u},\boldsymbol{p})$ for some initial conditions $\boldsymbol{u}(\tau_{0})$ and parameters $\boldsymbol{p}$, including several perturbation wavenumbers, and compute desired quantities from their solution, like power spectra.
 In practice, however, several properties complicate this task.
@@
 Second, the set of equations is very long and convoluted.
-Ideally, parts of the equations should be easily replaceable to accommodate different cosmological (sub)models.
+Ideally, parts of the equations should be easily replaceable to accommodate different submodels for gravity and particle species.
 Organizing a code with a suitable modular structure that scales well in model space is hard.
-Third, tension between wildly different timescales that coexist in the equations make them extremely stiff and intractable for standard explicit ODE solvers.
+Third, wildly different timescales coexist in the equations and make them extremely stiff and intractable to solve with standard explicit ODE solvers.
 This stiffness must be massaged away in the equations or dealt with numerically in other ways.
@@
 Typical models with accurate treatment of photons and neutrinos need $O(100)$ equations.
-Fifth, the perturbations must be solved for many different wavenumbers $k$, and tradeoffs between performance and accuracy must be made.
-To overcome these challenges, Boltzmann solvers are typically written in low-level high-performance languages like C, C++ and Fortran.
-Codes are usually tightly adapted to the pipeline-like computational structure of the problem (e.g.
-input $\rightarrow$ background $\rightarrow$ thermodynamics $\rightarrow$ perturbations $\rightarrow\ldots\rightarrow$ output in section 3.3 in Lesgourgues (2011a)).
-This makes sense for programmers and computers, but not always for modelers.
-Traditional codes have nonexistent or only very thin abstraction layers around them.
-To modify them, users must work directly in the low-level numerical interface and should be prepared to get their hands dirty.
-For example, to implement a new species, one must generally modify the code in many places: input handling for any new parameters, new background equations, new thermodynamics equations, new perturbation equations, joining between each of these stages, output handling, and possibly more.
-This leads to fragmented code where the changes related to one species are scattered throughout the code, making it challenging and unintuitive to navigate and decipher.
-More importantly, this structure scales poorly in model space.
-Each module becomes more polluted as more components are added.
-Even if they are deactivated by if statements at runtime, their mere presence in the source code reduces its readability.
-In turn, the whole code grows into a complex and mysterious beast.
+Fifth, the perturbations must be solved for many different wavenumbers $k$, and tradeoffs between performance and precision must be made.
+To overcome these challenges, most Boltzmann solvers are written in low-level high-performance languages like C, C++ and Fortran.
+They are tightly adapted to the pipeline-like computational structure of the problem (e.g.
+input $\rightarrow$ background $\rightarrow$ thermodynamics $\rightarrow$ perturbations $\rightarrow\ldots\rightarrow$ output in Lesgourgues (2011a)).
+This makes sense for programmers, but does not necessarily provide the simplest interface for modelers.
+Traditional codes like CAMB and CLASS have nonexistent or thin abstraction layers.
+To modify them, users must work directly in the low-level numerical code and have a good understanding of its internal structure.
+For example, to implement a new species, one must often modify the code in many places: input handling for new parameters, new background equations, new thermodynamics equations, new perturbation equations, joining each of these stages, output handling, and so on.
+This leads to fragmented code where changes related to one species are scattered throughout the code.
+This structure scales poorly in model space.
+As more species or gravitational theories are added, each module becomes intertwined with code from other physical components.
+Even if deactivated by if statements at runtime, their mere presence in the source code increases its complexity and reduces its readability.
 One can alleviate this problem to some extent by instead forking the code for modified models, so the main code base is not polluted.
@@
 They are also incompatible with each other unless one merges them into one code, but then one is back to the first problem.
-The two-language problem amplifies the problem, as data analysis usually takes place in slower high-level languages like Python.
-This shapes Boltzmann solvers to rigid pipelines that must compute everything in one shot and avoid interception at all costs, in order to maximize performance in the low-level language before passing the data back to the high-level language.
-Some features that are really just post-processing of the ODE solutions are appended to the pipeline even if they are peripheral to the core task of an Einstein-Boltzmann solver – which is to solve the Einstein-Boltzmann equations.
-This includes non-linear boosting and CMB lensing, for example.
-These code smells lead to big monolithic ?black boxes?
-that scale poorly in model space, and whose complexity grows to incorporate features beyond their original core scope.
-This complexity has even fueled development of specialized AI assistants for CLASS (Casas et al., 2025).
-While such tools are helpful, we think they are a symptom of unnecessary complexity and poor scaling with the number of models.
-Of course, there are also some advantages to this structure, as we discuss in section 4.
+The two-language problem amplifies the problem, as data analysis often happens in slow high-level languages like Python.
+This shapes Boltzmann solvers to rigid pipelines that must compute everything at once and avoid interception at all costs, in order to maximize performance in the low-level language before passing output back to the high-level language.
+Some features are really just post-processing of the ODE solutions, but are appended to the pipeline even if peripheral to the core task of an Einstein-Boltzmann solver – solving the Einstein-Boltzmann equations.
+Features like non-linear boosting and CMB lensing could be made as smaller and interoperating modular packages, but instead become part of ?all-in-one?
+Boltzmann codes.
+These patterns lead to big monolithic Boltzmann solvers.
+They become increasingly complex as they incorporate more models and features beyond their original core scope.
+This complexity has even driven development of specialized AI assistants for CLASS (Casas et al., 2025).
+While existing Boltzmann solvers are impressively well-engineered and such tools are only helpful, we think these are symptoms of unnecessary complexity.
 1.2 Boltzmann solver approximation schemes
-The Einstein-Boltzmann equations are infamously stiff.
-This is a property of differential equations (3) that means their numerical solution is unstable and requires tiny step sizes with standard explicit Runge-Kutta methods.
+The Einstein-Boltzmann equations are infamously stiff (Nadkarni-Ghosh and Refregier, 2017).
+This property of differential equations means their numerical solution is unstable with standard explicit integrators and requires tiny step sizes, making them hard to solve.
 Stiffness can arise when multiple and very different time scales appear in the same problem.
@@
 Doran (2005b); Blas et al.
-(2011)):
+(2011); Lewis (2025)):
   • tight-coupling approximation (TCA),
@@
   • radiation streaming approximation (RSA),
+  • non-cold dark matter fluid approximation (NCDMFA),
   • Saha approximation.
-These usually involve switching from one set of equations to another when some control variable that measures the applicability of the approximation crosses a threshold.
-This can change the unknowns in the ODE and require reinitializing it.
-The approximations help stability and sometimes performance of the equations and allow the use of explicit solvers.
-However, approximations put much more strain on the modeler to derive and validate versions of their equations in different regimes.
-Furthermore, this process should generally be repeated when further changes are made to the model.
-The numerics also become more complicated: timeseries from each ODE solution must be stitched together, tolerances are duplicated for each separate ODE system that may respond differently to them, ODE integrators must be reinitialized, and so on.
-It is a misunderstanding in the literature that it is ?difficult?
-or ?impossible?
-to solve the Einstein-Boltzmann equations without using approximations, or that they are ?unavoidable?
-(Ma & Bertschinger, 1995; Lewis, 2025; Lesgourgues, 2011a, respectively).
-This statement must only be understood in the context of explicit solvers!
-Implicit integrators can solve stiff equations, but seem largely underutilized by Boltzmann solvers, perhaps because they are scarce and harder to implement.
-CLASS has an implicit solver, but does not permit disabling all approximations Blas et al.
-(2011).
-On the contrary, new life has been breathed into this field recently, as development of modern implicit solvers and techniques like automatic and symbolic differentiation make implicit solvers more feasible and powerful.
+They involve switching from one set of equations to another when some control variable measuring the applicability of the approximation crosses a threshold.
+This can change the variables in the ODE and require reinitializing it.
+Approximations can help both speed and stability and allow the use of explicit solvers.
+However, approximations put much more strain on modelers to derive and validate several versions of the equations in different regimes.
+This process may have to be repeated when the model is modified.
+The numerics also become more complicated: timeseries from each ODE solution must be stitched together, each separate ODE system can use different tolerances, ODE integrators must be reinitialized, and so on.
+Another way to integrate stiff equations is to use appropriate implicit solvers.
+PyCosmo first solved the full stiff Einstein-Boltzmann equations with an implicit integrator, followed by Bolt.jl and DISCO-EB.
+CLASS also has an implicit solver, but does not permit disabling the tight-coupling approximation, for example.
+Historically, implicit solvers may have been underutilized because they are harder to implement than explicit solvers and have a reputation for being slower, and due to iterative evolution from initial Boltzmann codes that were centered around approximations.
+However, new life has recently been breathed into the field of implicit solvers, with development of new methods combined with techniques like automatic and symbolic differentiation that make them more feasible and powerful (e.g.
+Steinebach, 2023; Ekanathan et al., 2025).
 1.3 Differentiation methods
@@
 Cosmological applications are no exception.
-For example, algorithms that optimize likelihoods and Markov chain Monte Carlo (MCMC) samplers for Bayesian parameter inference can take advantage of derivatives of the likelihood with respect to each parameter to intelligently step in a direction where the likelihood or probability increases.
-In machine learning, the same applies when training neural networks emulators for cosmological observables as functions of cosmological parameters by minimizing scalar loss functions.
+For example, algorithms that optimize likelihoods and Markov chain Monte Carlo (MCMC) samplers for Bayesian parameter inference can take advantage of derivatives of the likelihood with respect to each parameter to intelligently step in a direction where the likelihood increases.
+In machine learning, the same applies when training neural networks emulators for cosmological observables as functions of parameters by minimizing a scalar loss.
 Some cosmologies are parametrized as boundary-value problems with the shooting method, and use nonlinear root solvers like Newton’s method that need Jacobians.
-Jacobians are also used by implicit ODE solvers to solve for values at the next time step.
-Fisher forecasting uses the Hessian (double derivative) of the likelihood with respect to parameters to predict how strong parameter constraints that can be placed by data with given errors.
-Boltzmann solvers typically save computational resources by interpolating spectra computed on coarse grids of $k$ and $l$ to finer grids, and this can be made more precise with knowledge of derivatives with respect to $k$ and $l$.
-These are all examples where Boltzmann solvers or applications that use them need derivatives.
+Implicit ODE solvers also use Jacobians to solve for values at the next time step.
+Fisher forecasting uses the Hessian (double derivative) of the likelihood with respect to parameters to predict how strong parameter constraints that can be placed by data with some uncertainty.
+Boltzmann solvers save time by interpolating spectra computed on coarse grids of $k$ and $l$ to finer grids, and this can be made more precise with derivatives with respect to $k$ and $l$.
+These are all examples where (applications of) Boltzmann solvers need derivatives.
 There are at least four ways to compute derivatives.
-Manual differentiation is the exact pen-and-paper method with differentiation rules, but is limited to simple expressions and by human error.
-Symbolic differentiation automates this process with computer algebra systems, but is inherently symbolic and cannot differentiate arbitrary programs nested with control flow through conditional statements and loops that depend on numerical values.
+Manual differentiation is human application of differentiation rules, but is limited to simple expressions and by human error.
+Symbolic differentiation automates this process with computer algebra systems, but is inherently symbolic and cannot differentiate arbitrary programs nested with control flow like conditional statements and loops that depend on numerical values.
 Finite differentiation approximates the derivative $f^{\prime}(x)\approx(f(x+\epsilon/2)-f(x-\epsilon/2))\,/\,\epsilon$ with a small $\epsilon>0$ (here using central differences) by simply evaluating the program several times.
-This can differentiate arbitrary programs, but is approximate, introduces the step size $\epsilon$ as a hyperparameter that must be tuned for both accuracy and stability, and is a brute-force approach that requires $O(n)$ ($2n$ using central differences) evaluations to compute the gradient of an $n$-variate scalar $f$.
-Automatic differentiation (e.g.
-Griewank & Walther, 2008) can be understood by viewing any computer program as a (big) composite function
+This can differentiate arbitrary programs, but is approximate, introduces the step size $\epsilon$ as a hyperparameter that must be tuned for both accuracy and stability, and is a brute-force approach that requires $O(n)$ evaluations ($2n$ using central differences) to compute the gradient of an $n$-variate scalar $f$.
+Automatic differentiation can be understood by viewing any computer program as a (big) composite function
 $\boldsymbol{f}=\boldsymbol{f}_{N}\circ\boldsymbol{f}_{N-1}\circ\cdots\circ\boldsymbol{f}_{2}\circ\boldsymbol{f}_{1}=\boldsymbol{f}_{N}(\boldsymbol{f}_{N-1}(\cdots\boldsymbol{f}_{2}(\boldsymbol{f}_{1})))$ (1)
@@
 through the Jacobian $(\boldsymbol{J}_{n})_{ij}=\partial f_{n,i}/\partial f_{n-1,j}$ of every elementary operation to accumulate the derivative of the entire program.
-This is numerically exact, free of precision parameters, usually requires fewer operations than finite differences, but is perhaps less intuitive, harder to implement and needs the source code to interpret the program (1) in the non-standard way (2).
+This is numerically exact, free of precision parameters, usually requires fewer operations than finite differences, but is perhaps less intuitive, harder to implement and needs the source code of the program (1) to interpret it in the non-standard way (2).
 Notably, while the function (1) must be evaluated inside-to-outside (right-to-left), the chain rule (2) is an associative matrix product that can be evaluated in any order.
@@
 In practice, both modes are implemented using techniques like operator overloading or source code transformation.
-The former overloads every operation to accept a special number type that propagates both values and derivatives, such as dual numbers for forward-mode (e.g.
+The former specializes every function on a particular ?dual number?
+that computes both the value and derivative of the function (e.g.
 Revels et al., 2016).
-The latter transforms the code for $\boldsymbol{f}$ into another code that computes $\boldsymbol{J}$.
+The latter analyzes the source code for $\boldsymbol{f}$ and transforms it into another code that computes $\boldsymbol{J}$.
 In any case, automatic differentiation requires access to the source code!
-GravityBaryonsPhotonsMasslessneutrinosMassiveneutrinosCold darkmatterCosmo-logicalconstant Symbolic model BackgroundPerturbations Numerical problem BackgroundPerturbations Solution object CompileSolve
-Figure 1: SymBoltz represents cosmological models with equations in symbolic form and grouped by physical components like the spacetime metric, gravity and particle species.
-This is compiled to a numerical problem that splits the model into background and perturbation stages and generates efficient numerical code for ODEs and their Jacobians.
-The problem is then solved and the result stored in a solution object that gives convenient access to any quantity defined in the symbolic model.
+GravityBaryonsPhotonsMasslessneutrinosMassiveneutrinosCold darkmatterCosmo-logicalconstant Symbolic model Background: $\boldsymbol{f}$, $\boldsymbol{J}$Perturbations: $\boldsymbol{f}$, $\boldsymbol{J}$ Numerical problem Background: $S(\tau)$Perturbations: $S(\tau,k)$ Solution object CompileSolve
+Figure 1: SymBoltz represents cosmological models with symbolic equations grouped in physical components for the metric, gravity and particle species.
+This is compiled to a numerical problem that splits equations into background and perturbation stages and generates fast code for ODE functions $\boldsymbol{f}$ and Jacobians $\boldsymbol{J}$.
+The problem is then solved and the result stored in a solution object that gives access to any variable $S$ in the model.
 2 Code architecture and main features
-SymBoltz is built somewhat differently than other Boltzmann codes.
-In short, it uses a symbolic-numeric abstraction interface where users enter high-level symbolic equations, and automatically compiles them to fast low-level numerical functions.
+SymBoltz is designed around three main features.
+In short, it has a symbolic-numeric abstraction interface where users enter high-level symbolic equations, and automatically compiles them to fast low-level numerical functions.
 It cures stiffness with implicit ODE solvers, and uses no approximation schemes to keep models simple, elegant and extensible.
-It is differentiable, so one can easily get accurate derivatives of any output with respect to any input.
-This provides rapid model prototyping, helpful abstractions and automation of typical chores when modifying other codes.
-SymBoltz encourages interactive usage and pursues a modular design that lets users integrate whatever they need from the package into their own applications.
-The end goal is to prioritize the modeler, who should be able to just write down the model equations while SymBoltz takes care of the rest.
-SymBoltz is written in the Julia programming language (Bezanson et al., 2017), which attracts development of modern numerical methods and aims to resolve the two-language problem.
-The symbolic-numeric interface is built on the ModelingToolkit.jl (Ma et al., 2022) and Symbolics.jl (Gowda et al., 2022) packages.
-The compiled functions are integrated by implicit ODE solvers in DifferentialEquations.jl (Rackauckas & Nie, 2017).
-Everything works with automatic differentiation through ForwardDiff.jl (Revels et al., 2016), for example.
-The next subsections present the design of SymBoltz in more detail, revolving around the three key features in the title.
-Other minor implementation details are given in appendix A.
-This paper is as of SymBoltz version 0.8.1.
-Please refer to the package documentation for definitive up-to-date information.
+It is differentiable, so one can get accurate derivatives of any output with respect to any input.
+This provides rapid model prototyping, helpful abstractions and automates tasks one must do manually when modifying other codes.
+SymBoltz encourages interactive use and pursues a modular design that lets users integrate what they need from the package into their own applications.
+The end goal is to prioritize the modeler, who should be able to just write down their equations while SymBoltz handles modeling chores.
+The code is written in the Julia programming language (Bezanson et al., 2017), with a rich ecosystem of scientific packages and an aim to resolve the two-language problem.
+The symbolic-numeric interface is built on ModelingToolkit.jl (Ma et al., 2022) and Symbolics.jl (Gowda et al., 2022).
+The compiled functions are integrated by implicit ODE solvers in OrdinaryDiffEq.jl (Rackauckas and Nie, 2017).
+Automatic differentiation works through ForwardDiff.jl (Revels et al., 2016).
+The next subsections describe SymBoltz’ three main features in more detail.
+Other implementation details are given in appendix A.
+This paper is as of SymBoltz version 0.14.1.
+Please see the package documentation for definitive up-to-date information.
 2.1 Symbolic-numeric interface
-The core of SymBoltz is built around a symbolic-numeric interface shown in fig.
+The core of SymBoltz is built around a symbolic-numeric interface illustrated in fig.
 1.
-Variables and equations are specified in a high-level user-friendly symbolic modeling language, then ultimately compiled down to efficient low-level numerical functions that are integrated by ODE solvers.
-This fits the interactive and just-in-time compiled nature of Julia.
-A key realization is that knowledge of symbolic equations makes it possible to analyze their structure programmatically, enabling powerful transformations of the equations and automation of chores.
-SymBoltz turns the layout of traditional Boltzmann solvers inside-out.
-While other codes are built like rigid pipelines following computational stages like the background and perturbations, SymBoltz is primarily structured around physical components of the Einstein-Boltzmann system.
-Full cosmological models are built by joining submodels for the metric, gravitational theory, photons, baryons, dark matter, dark energy, neutrinos and other species.
-Each component is a chunk of related variables and internal equations, and is unaware of other components.
+Variables and equations are specified in a high-level user-friendly symbolic modeling language, then compiled down to low-level numerical code that is integrated by ODE solvers.
+With knowledge of the symbolic equations, SymBoltz can analyze their structure programmatically to ease model specification, automate related boilerplate tasks and generate fast and stable code that avoids approximations.
+We think three properties motivate some kind of symbolic abstraction layer around the Einstein-Boltzmann equations.
+First, the equations are most easily written as one large ODE, but optimally solved as several ODEs in stages like the background, perturbations and line-of-sight integration.
+Mathematically, the Jacobian matrix of this single ODE encodes which variables are independent and belong to which stages.
+It is easiest for modelers to specify equations in one common system, where the perturbations can refer to variables in the background, for example.
+The separation into computational stages can be automated.
+Second, solving each stage is hard due to stiffness, performance requirements and its dependence on previous stages.
+To overcome this, it is helpful to automatically generate ODE code that is optimal in speed, accuracy and stability, generate the ODE Jacobian needed by implicit solvers and interpolate variables from previous stages.
+Modelers should not worry about this.
+Third, the equations are often modified because the true cosmology is unknown.
+This can be made easier by organizing equations in a way that reflects the physical structure of the model, so it is easy to replace a related subset of equations with another without duplicating the entire model.
+This could be slow to do at the numerical level, but can be done with no runtime penalty in an intermediate symbolic representation of the equations.
+Inspired by this, SymBoltz inverts the traditional layout of Boltzmann solvers.
+While most codes are built as pipelines that follow the background, perturbations and other stages of the equations, SymBoltz is primarily structured around the physical components that they consists of.
+Related variables and equations are grouped in distinct components for the metric, gravitational theory, photons, baryons, dark matter, dark energy, neutrinos and other species (see fig.
+1).
+This isolates everything related to one submodel to one place.
+Any set of such submodels are joined into a full cosmological model, like $\mathrm{\Lambda CDM}$.
 Interactions (e.g.
-Compton scattering or sourcing of gravity) are equations that connect components.
-This structure is made possible by the symbolic interface and would slow a purely numerical code.
-Separation of computational stages is secondary and done automatically (see section 2.1.3).
-This makes it possible to write everything related to one component in one place, instead of scattering it across separate modules for input, background, perturbations and so on.
-One component can be replaced by another to construct a different cosmological model without polluting a big global model.
-A large part of SymBoltz is thus simply devoted to building a well-organized library of such physical components.
-This modular structure scales better in model space and makes it easier to assemble both extended and reduced models.
-While many Boltzmann solvers hardcode baryons and photons because they are fundamental to the code, SymBoltz works naturally even with pure models with non-interacting radiation, matter and dark energy.
-For example, this can help understand how a modified gravity theory responds to different species without complex interactions or thermodynamics cluttering the picture.
-Maybe this structure is unfamiliar to some.
-As a compromise, SymBoltz also provides an ?unstructured?
-version of the $\mathrm{\Lambda CDM}$ model with all equations and variables in one big system.
-This makes it easy-peasy to change anything in the model, and really shows the power of the symbolic interface: SymBoltz defines the full $\mathrm{\Lambda CDM}$ model (with GR, baryons, RECFAST recombination, photons, cosmological constant, massive and massless neutrinos) in just 277 lines of code, while the equivalent code that one must realistically browse in CLASS is spread over 10 files with 27721 lines!
-[^8] Counting input.{h,c}, background.{h,c}, thermodynamics.{h,c}, perturbations.{h,c} and wrap_recfast.{h,c} with wc -l.
-SymBoltz has less features than CLASS, but this is not even remotely close to making up for the $100\times$ simplification.
-Let us see which helpful features this interface provides.
+Compton scattering or sourcing of gravity) are equations between components.
+Adding new submodels is easy, and SymBoltz is largely devoted to simply building a well-organized library of them.
+SymBoltz then takes a full cosmological model, separates it into stages and generates numerical code to solve each stage.
+This preprocessing does not slow the code at runtime; to the contrary the symbolic equations give extra information like the Jacobian that enhances speed and stability.
+Modifications automatically enjoy these benefits without extra user input.
+The modular component-based structure scales well in model space.
+For example, modified gravity or dark energy models can be written down as self-contained submodels and combined with other components to create many different extended cosmological models.
+The generic symbolic framework can also be used to build reduced models with non-interacting radiation, matter and dark energy species.
+When exploring modified gravity theories, it can be helpful to test such toy models to understand how gravity responds to pure fluids without coupling to baryons and photons.
+In contrast, the monolithic layout of traditional codes scatter changes related to one submodel across different modules for input/output, the background, perturbations and so on.
+As more and more submodels are added, the code is increasingly intertwined and fragmented.
+They are also not flexible enough to test simpler toy models, as baryons and photons are hardcoded as fundamental to the structure of the code.
+This design fits the computational structure better than the physical one.
+It can result in an overwhelming code that becomes hard to read and modify, and whose complexity scales poorly in model space.
+As an alternative to component-based modeling, SymBoltz also includes a version of the default $\mathrm{\Lambda CDM}$ model with all equations packed into one big ?unstructured?
+system.
+This makes it trivial to change anything in the equations, but sacrifices modularity.
+The symbolic interface makes it very compact: SymBoltz defines a full $\mathrm{\Lambda CDM}$ model in just 277 straightforward lines of code [^8] https://hersle.github.io/SymBoltz.jl/stable/unstructured/, while the equivalent code to browse in CLASS is spread over 10 files with 27721 lines [^9] Counting input.{h,c}, background.{h,c}, thermodynamics.{h,c}, perturbations.{h,c} and wrap_recfast.{h,c} with wc -l..
+PyCosmo’s symbolic interface also generates code for ODEs and sparse Jacobians, but does not focus on component-based modeling or automating tasks like stage separation, for example.
+SymBoltz’ symbolic-numeric interface simplifies solving and modifying the Einstein-Boltzmann equations.
+It aims to maximize speed, stability and convenience with minimal user input.
+The next sections go into detail on how this is done.
+Section 3.2 shows a concrete example and some of its features in action.
 2.1.1 Automatic numerical code generation
 SymBoltz automatically compiles symbolic equations to numerical code for ordinary differential equations (ODEs)
-$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}\tau}=\mathbf{f}(\tau,\mathbf{u}).$ (3)
-The generated code is fast and prevents users unfamiliar with the language or package from writing slow numerical code.
-If necessary, one can escape the standard code generation and include arbitrary numerical code, but the user is then fully responsible for its performance.
-This accommodates functions that cannot be written as straightforward equations, for example if one must solve a nonlinear equation for the minimum of a potential or interpolate tabulated data.
-The code generation deals with chores like dynamically allocating indices for each ODE state $u_{i}$, and performs optimizations such as common subexpression elimination.
-This is helpful as Boltzmann solvers tend to have big $\boldsymbol{f}$, and particularly with implicit ODE solvers that call $\boldsymbol{f}$ often.
+$\frac{\mathrm{d}\boldsymbol{u}}{\mathrm{d}\tau}=\boldsymbol{f}(\tau,\boldsymbol{u},\boldsymbol{p}).$ (3)
+The generated code is fast and prevents users unfamiliar with Julia or SymBoltz from writing slow code.
+If necessary, one can escape the standard code generation and call arbitrary numerical functions, for example to solve a nonlinear equation for the minimum of a potential or interpolate tabulated data.
+The code generation handles tasks like allocating indices for each ODE state $u_{i}$, and performs optimizations such as common subexpression elimination (see section 3.2).
+This is helpful as Boltzmann solvers tend to have big $\boldsymbol{f}$ and ODE solvers call $\boldsymbol{f}$ many times.
 2.1.2 Automatic handling of observed variables
-In general, an ODE (3) admits two types of variables: unknowns $\mathbf{u}(\tau)$ are the variables that are integrated with respect to time, and observed variables are functions of the unknowns.
-The Einstein-Boltzmann equations are commonly formulated with many observed variables.
+A general ODE (3) has two types of variables: ?unknowns?
+$\boldsymbol{u}(\tau)$ are integrated with respect to time, while ?observed?
+variables are any functions of the unknowns.
+The Einstein-Boltzmann equations are usually written with many observed variables.
 For example, consider the metric and gravity equations in sections A.1 and A.2 sourced by some arbitrary $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$.
-Here $a(\tau)$ and $\Phi(\tau,k)$ are the only unknown (differential) variables that the ODE solver must solve for, while one can ?observe?
-$z(\tau)$, $\mathscr{H}(\tau)$, $H(\tau)$, $\chi(\tau)$ and $\Psi(\tau,k)$ from the unknowns.
-Of course, it is always possible to eliminate all observeds by explicitly inserting them into the equations for the unknowns.
-But this kills readability, as observed variables are helpful and natural intermediate definitions that break up the system, and one may want to extract them from the solution as well.
+Here $a(\tau)$ and $\Phi(\tau,k)$ are the only unknown (differential) variables that the ODE is solved for, while one can ?observe?
+$z(\tau)$, $\mathscr{H}(\tau)$, $H(\tau)$ and $\Psi(\tau,k)$ from the unknowns.
+In addition, observed derivatives like $\mathscr{H}^{\prime}$ or $\Psi^{\prime}$ are not always trivial to express, but SymBoltz automatically expands them using the definitions of $\mathscr{H}$ or $\Psi$.
+Of course, one can eliminate all observeds by explicitly inserting them into the equations for the unknowns.
+But this reduces readability, as observed variables are helpful intermediate definitions that break up the equations, and one may want to extract them from the solution as well.
 Furthermore, modified models can change the sets of unknown and observed variables (e.g.
 modified gravity can change the constraint equation for $\Psi$ into a differential equation).
-It is easier to compose models when variables do not have to be hardcoded as either unknown or observed.
-SymBoltz reads in an entire system of equations like that defined by appendix A and automatically figures out which variables are unknown and observed.
-It then generates numerical code for solving the ODE for the unknowns (see section 2.1.1), and can lazily compute any observed variable or expression thereof from the solution of the unknowns.
-The bottom line is that the user can trivially use and obtain any variable anywhere just by referring to it, whether it is unknown or observed.
-In other solvers one must look up the desired expression for observed variables and recompute them manually.
-This is tedious, error-prone and can tempts users to neglect parts of expressions (e.g.
-approximating $\Psi\approx\Phi$ by neglecting anisotropic stress).
+It is easier to compose models when variables are not hardcoded as either unknown or observed.
+SymBoltz reads in a full system of equations like that defined by appendix A and separates variables into unknowns and observeds.
+It then generates numerical code for solving the ODE for the unknowns (see section 2.1.1), but can automatically recompute any observed variables from the solution of the unknowns.
+The bottom line is that users can easily use any variable anywhere by just referring to it, whether it is unknown or observed.
+In other solvers one may have to look up the expression for observed variables and recompute them manually.
+This is tedious, error-prone and can tempt users to make simplifying assumptions (e.g.
+approximating $\Psi\approx\Phi$ without anisotropic stress).
 2.1.3 Automatic stage separation and splining of unknowns
-In principle, one can solve the Einstein-Boltzmann equations by integrating the entire system (i.e.
-background and perturbations) at the same time.
+In principle, one can solve the Einstein-Boltzmann equations by integrating the entire system (e.g.
+background and perturbations) at once.
 However, the system can be broken down into sequential computational stages that each depend only on those before it.
 To alleviate stiffness in each stage, separate concerns by solving every equation only once (e.g.
-avoid recomputing the background for every perturbation mode) and to improve performance, it is common to solve the system stage-by-stage and spline variables from one stage as input to the next.
-To illustrate, again consider the general relativistic equations in section A.2 sourced by some given $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$.
-Clearly $\Phi(\tau,k)$ and $\Psi(\tau,k)$ depend on $a(\tau)$, but $a(\tau)$ does not depend on $\Phi(\tau,k)$ or $\Psi(\tau,k)$ (this just reflects the perturbative nature of the problem).
-One can first solve for only $a(\tau)$ in the background, then spline $a(\tau)$ and use it as input for solving for $\Phi(\tau,k)$ and $\Psi(\tau,k)$ in the perturbations, instead of solving for all three together and repeatedly solve for $a(\tau)$ for every $k$.
-SymBoltz uses the same stage separation strategy, but automates and strengthens it with its knowledge of the symbolic equations.
-First, the full symbolic model is split into background and perturbation systems.
-Cubic Hermite splines are then constructed for all unknown variables (like $a(\tau)$) to pass their solution from one stage to the next.
-This type of splines is perfect for interpolating ODEs, as cubic Hermite splines increase accuracy by taking both $\boldsymbol{u}(\tau)$ and $\boldsymbol{u}^{\prime}(\tau)$ into account, and $\boldsymbol{u}^{\prime}(\tau)$ is known analytically by definition (3).
-In contrast, observed variables (like $\Psi(\tau,k)$) are computed from the (splined) unknowns because their derivatives are not known analytically.
-Splining them directly would be less accurate.
-Modelers can simply write down a new background variable and access it in the perturbations for free.
-They can access any variable anywhere as if solving the entire set of equations at once, while still benefiting from splining under the hood.
-The separation into background and perturbation stages just reflects the perturbative structure of the linearized Einstein-Boltzmann equations, where each order depends only on those before it, and is always guaranteed.
+avoid recomputing the background for every perturbation mode) and to improve performance, all Boltzmann codes solve the system stage-by-stage and spline variables from one stage as input to the next.
+To illustrate, again consider the general relativistic equations in section A.2 sourced by some known $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$.
+Clearly $\Phi(\tau,k)$ and $\Psi(\tau,k)$ depend on $a(\tau)$, but $a(\tau)$ does not depend on $\Phi(\tau,k)$ or $\Psi(\tau,k)$, reflecting the perturbative nature of the problem.
+One can first solve for only $a(\tau)$ in the ?background?, then spline and look up $a(\tau)$ to solve for $\Phi(\tau,k)$ and $\Psi(\tau,k)$ in the ?perturbations?, instead of solving for all three together and repeatedly solve for $a(\tau)$ for every $k$.
+SymBoltz uses the same stage separation strategy as any other Boltzmann code, but automates it with its knowledge of the symbolic equations.
+First, all equations are split into background and perturbation stages.
+Cubic Hermite splines are then constructed for all background unknowns (like $a(\tau)$) and passed on to the perturbations.
+This spline type is optimal for interpolating ODEs, as Hermite splines take both $\boldsymbol{u}(\tau)$ and $\boldsymbol{u}^{\prime}(\tau)$ into account for better accuracy, and $\boldsymbol{u}^{\prime}(\tau)$ is known analytically from $\boldsymbol{u}(\tau)$ and definition (3).
+In contrast, observed variables (like $\mathscr{H}(\tau)$) are computed from the (splined) unknowns because their derivatives are not known directly, and splining them is less accurate.
+This makes any background variable available in the perturbations ?for free?.
+All stages of the Einstein-Boltzmann equations are solved as if it is written as one combined system of equations, while splining and stage separation are done automatically under the hood.
+The separation into background and perturbation stages is guaranteed and just reflects the perturbative structure of the linearized Einstein-Boltzmann equations, where each order depends only on those before it.
 However, they can often be broken further down: most thermodynamics (recombination) models can be separated from the background, and some variables have integral solutions (e.g.
 the optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\bar{\tau})\mathrm{d}\bar{\tau}$ or line-of-sight integration) that can be computed in isolation after solving the differential equations.
-Over time, SymBoltz aims to extend the automatic background-perturbation separation by automatically incorporating more or all stages of the equations at hand.
-[^9] All stages can in principle be identified from a directed dependency graph between variables (e.g.
-from the Jacobian matrix).
-Cycles in the graph correspond to interdependent equations that must be integrated together in one ODE system (e.g.
-background and perturbations).
-Leaves with differentiated variables correspond to integral solutions.
-This feature is a major advantage with access to the symbolic equations.
+In the future, SymBoltz’ background-perturbations separation could be extended to automatically split all smaller stages by symbolically inspecting the dependencies between variables in the equations.
 2.1.4 Automatic solution interpolation
-SymBoltz integrates equations in conformal time $\tau$ for several given wavenumbers $k$, and stores the result in an object that wraps independent ODE solutions for the background and perturbations for every $k$.
-Conveniently, this solution object can be queried for any variable or symbolic expression
-$x(\tau,k)=x(\boldsymbol{u}(\tau,k_{i})).$ (4)
-at any $\tau$ and $k$.
-The solution object automatically expands $x$ in terms of unknowns $\boldsymbol{u}$ (if $x$ is observed), and interpolates between solved wavenumbers $k_{i}$ (if $x$ is perturbative) and in $\tau$ using the ODE solver’s built-in interpolation method (or as in section 2.1.3 if $x$ is splined).
-This provides the modeler easy access to any variable and expression defined in the model.
-The solution interpolation is also incorporated into convenient plot recipes for trivial visualization of any expressions for any $\tau$ and $k$.
+SymBoltz integrates the background and several perturbation ODEs in conformal time $\tau$ for several wavenumbers $k$.
+All the results are stored in a special solution object.
+This solution object can be queried for any variable or symbolic expression
+$S(\tau)\quad\text{or}\quad S(\tau,k).$ (4)
+This looks up the background solution if $S$ is a background variable, or the perturbation solutions if it is perturbative.
+It interpolates from the times stored by the ODE solver to the requested $\tau$ using the solver’s dense interpolation.
+If $S$ is perturbative, it also interpolates from the solved perturbation $k$-modes to the requested $k$.
+If $S$ is an unknown, it is returned directly from the ODE solution.
+If $S$ is observed, it is instead recomputed automatically from the equation that defines it as a function of the unknowns.
+The result is that the user can access any variable in the model without having to do this manually.
+The solution interpolation is also incorporated into plot recipes for visualization of any variable as a function of $\tau$ and/or $k$.
 2.1.5 Automatic Jacobian generation and sparsity detection
-Just like SymBoltz generates code for $\boldsymbol{f}$ in the ODE (3), it can use the same symbolic equations to generate its Jacobian
+Just like SymBoltz generates code for $\boldsymbol{f}$ in the ODE (3), it uses the same symbolic equations to generate its Jacobian $\boldsymbol{J}$ with entries
 $J_{ij}=\frac{\partial f_{i}}{\partial u_{j}}.$ (5)
-Jacobians are crucial for solving stiff ODEs accurately and efficiently with implicit solvers (see section 2.2)!
-Manually coding them is an extremely tiresome and error-prone undertaking that must be repeated for model modifications.
-Evaluating them numerically with finite differences is approximate and expensive.
-Automation of this procedure is a powerful feature that again enables the modeler to focus solely on the main equations.
-Another benefit is that analytical knowledge of the Jacobian lets one find its exact sparsity pattern.
-Numerical approaches to this are less robust and can mistake local zeros for global zeros.
-This permits constant factorization of sparse Jacobians, which improves performance of implicit solvers (see section 2.2).
-Full support for sparse Jacobians is still ongoing work, however.
-SymBoltz can fall back from symbolic to automatic differentiation for the Jacobian. Finite differences can also be used.
-2.1.6 Automatic change of variable (future work)
-SymBoltz consistently formulates all differential equations with respect to conformal time $\tau$ as the independent variable.
-This is perhaps the most common parametrization in the literature.
-It is very natural because the equations are autonomous with respect to $\tau$ (i.e.
-$f(\tau,\boldsymbol{u})\rightarrow f(\boldsymbol{u})$, except some multipole truncation schemes that use $1/(k\tau)$, but these are unphysical).
-Some other codes inconsistently use different independent variables in different stages.
-This requires manual translation into cosmic time $t$, redshift $z$, scale factor $a$ or its logarithm $b=\log a$, for example, involving differential transformations like $\mathrm{d}t=a(\tau)\,\mathrm{d}\tau$, $\mathrm{d}z=-(a^{\prime}(\tau)/a(\tau)^{2})\,\mathrm{d}\tau$, $\mathrm{d}a=a^{\prime}(\tau)\,\mathrm{d}\tau$ and $\mathrm{d}b=(a^{\prime}(\tau)/a(\tau))\,\mathrm{d}\tau$.
-This is another mechanical and error-prone process.
-In the future, SymBoltz could automate change to another independent variable (which should be one-to-one with $\tau$) by transforming the symbolic expressions.
-This can be particularly useful to compute observables as a direct function of redshift, although one can invert $z(\tau)$.
-This is common in fits to observations.
-2.1.7 Automatic unit handling (future work)
-Most variables in SymBoltz are defined in internal dimensionless units (see appendix A).
-In the future, SymBoltz could use its symbolic pre-processing for more powerful unit features.
-For example, equations could be checked for dimensional validity to catch modeling mistakes.
-Input quantities could be automatically transformed from the user’s preferred (dimensionful) units (e.g.
-Mpc) to internal (dimensionless) units, and then back to the user’s units for output.
-2.1.8 Automatic gauge transformation (future work)
-SymBoltz is currently formulated only in the conformal Newtonian gauge.
-In the future, SymBoltz could conveniently transform the symbolic equations to other gauges automatically, such as the synchronous gauge used in CAMB and CLASS.
-2.1.9 Automatic initial conditions (future work)
-SymBoltz currently uses adiabatic initial conditions for the perturbations.
-In the future, SymBoltz could generate initial conditions for arbitrary models from the equations.
-This typically involves mechanical procedures like perturbation theory or power series expansion that could be automated.
-2.1.10 Automatic approximation schemes (future work)
-SymBoltz is initially designed to be free of approximation schemes (see section 2.2).
-In the future, SymBoltz could provide convenience utilities to help derive approximations to the equations.
-For example, it could automatically expand equations in power series of some smallness parameter.
-This could accelerate derivation of approximations for modified models to improve their performance with less effort.
+Jacobians are crucial for solving stiff ODEs with implicit solvers.
+Manually coding them is tiresome, error-prone, and must be repeated for new models.
+Numerical evaluation with finite differences is both approximate and slow.
+It can also be computed with automatic differentiation, but we find this slower and overkill compared to the analytical $\boldsymbol{J}$.
+Analytical Jacobians are as fast and stable as possible, and automatic generation lets the solver benefit from it while the modeler focuses on the high-level equations.
+SymBoltz stores the Jacobian in sparse form to boost performance when it has many zeros.
+From the analytical $\boldsymbol{J}$, SymBoltz precomputes its exact sparsity pattern (where $J_{ij}=0$).
+This is hard to specify manually, and numerical solvers cannot distinguish false local zeros (for some input) from true global zeros (for all inputs).
+Without approximations, SymBoltz then reuses the same sparsity structure throughout the integration, computes the nonzero $J_{ij}$ analytically and stores them directly in the sparse $\boldsymbol{J}$ with no extra cost from looking up sparse matrix indices.
+Easy, accurate, fast and sparse Jacobians are crucial for SymBoltz to remain performant without approximations! It is a major advantage with the symbolic approach, as in PyCosmo.
 2.2 Approximation-freeness
 SymBoltz treats stiffness in the Einstein-Boltzmann equations with modern implicit ODE solvers that integrate the full equations at all times.
-It is therefore free of approximation schemes, such as the TCA, UFA, RSA and Saha approximation.
-This is much friendlier to the modeler, who now simply has to provide a single set of equations instead of deriving approximations, validating them and dealing with related chores.
-However, implicit ODE solvers are more expensive than explicit methods.
-In particular, at every time step they must solve a nonlinear system of equations for the next unknowns.
-This is usually done with Newton’s method that requires the Jacobian of the nonlinear system, which in turn involves the ODE Jacobian (5).
-Automatic, accurate and efficient computation of the Jacobian (see section 2.1.5) makes this less of a problem.
-As the size of the ODE and its Jacobian increases (e.g.
-larger $l_{\text{max}}$), solving the large system of equations at every time step becomes a bottleneck.
-One should then transition from dense to sparse matrix methods (this is still ongoing work) and use performant matrix methods when solving the nonlinear system.
-By default, SymBoltz uses the Rodas4P (Steinebach, 1995) solver for the background (stiff when thermodynamics is included) and KenCarp4 (Kennedy & Carpenter, 2003) for the perturbations.
-These are both implicit solvers from DifferentialEquations.jl.
-We discuss this aspect further in section 4.
-In the long run, SymBoltz could also implement approximation schemes to maximize speed (see section 2.1.10).
-However, the goal is to keep the code primarily approximation-free and for all approximations to be secondary and optional.
+It is therefore free of approximation schemes, such as the TCA, UFA, RSA, NCDMFA and Saha approximation.
+This is friendlier to the modeler, who now simply has to provide a single set of equations instead of deriving approximations, validating them and dealing with related chores.
+This pairs well with a high-level equation-oriented symbolic interface.
+Implicit solvers take more expensive steps than explicit methods.
+At every step, they solve a generally nonlinear system of equations for the next unknowns $\boldsymbol{u}$.
+Using Newton’s method, it iteratively solves linear systems $\boldsymbol{W}\boldsymbol{u}=\boldsymbol{b}$ by LU-factorizing $\boldsymbol{W}=\boldsymbol{I}-\gamma h\boldsymbol{J}$, where $\gamma$ is a constant, $h$ is the time step and $\boldsymbol{J}$ is the ODE Jacobian (5).
+[^10] For example, the implicit Euler method $\boldsymbol{u}_{n+1}=\boldsymbol{u}_{n}+h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})$ solves $\boldsymbol{F}(\boldsymbol{u}_{n+1})=\boldsymbol{u}_{n+1}-h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})-\boldsymbol{u}_{n}=\boldsymbol{0}$ at every step with Newton’s method using its Jacobian $\boldsymbol{W}=\boldsymbol{\nabla}\boldsymbol{F}=\boldsymbol{I}-h\boldsymbol{J}(\tau_{n+1},\boldsymbol{u}_{n+1})$.
+LU-factorizing a dense $n\times n$ matrix from an $n$-dimensional ODE costs $O(n^{3})$ operations and bottlenecks larger ODEs (e.g.
+larger $l_{\text{max}}$).
+In return, implicit methods use more information about the ODE to take stable and long steps.
+To be fast, implicit solvers need several optimizations.
+They must compute not only $\boldsymbol{f}$, but also $\boldsymbol{J}$ efficiently.
+A common trick is to reuse an LU-factorized $\boldsymbol{W}$ over several steps, even if it really changes.
+Newton’s method only needs an approximate $\boldsymbol{W}$ to find the root, and $\boldsymbol{W}$ is recomputed only if convergence is slow due to an outdated $\boldsymbol{J}$ or a new step size $h$.
+This trades fewer LU-factorizations and $\boldsymbol{J}$ evaluations for more linear solve back-substitutions, which only cost $O(n^{2})$ operations.
+The next trick is to speed up the linear algebra with sparse matrices if $\boldsymbol{W}$ has many zeros, as for the perturbations.
+SymBoltz handles this by generating the analytical and sparse Jacobian (see section 2.1.5).
+SymBoltz generates code for all implicit solvers in OrdinaryDiffEq.jl [^11] https://github.com/SciML/OrdinaryDiffEq.jl/ and dense and sparse matrix methods in LinearSolve.jl [^12] https://github.com/SciML/LinearSolve.jl.
+We can test different solvers, which is useful as performance and stability of implicit methods and linear algebra operations are problem-dependent.
+Using $O(100)$ perturbation equations with more than $95\%$ sparsity, we find RFLUFactorization fastest for dense matrices, but KLUFactorization several times faster when exploiting sparsity.
+We tested these ODE solvers:
+  • Backward Differentiation Formula (BDF) methods like FBDF and QNDF (Shampine and Reichelt, 1997) are multi-step methods that use several past points to solve for the next step.
+They do not have Runge-Kutta ?stages?
+and solve just one implicit equation per step, so they scale well to large ODEs and can reuse Jacobians aggressively.
+They have variable order of 1-5, but are ?L-stable?
+only up to 2nd order.
+This makes them take small steps in stiff regimes without utilizing their high order.
+We find them to be slowest.
+The solvers ndf15 in CLASS and BDF2 in PyCosmo belong to this family.
+Nadkarni-Ghosh and Refregier (2017) also discuss BDF methods in the context of the Einstein-Boltzmann equations.
+  • Explicit Singly-Diagonal Implicit Runge-Kutta (ESDIRK) methods (e.g.
+Jørgensen et al., 2018) are a class of implicit Runge-Kutta methods designed to preserve most stability properties of Fully Implicit (FIRK) methods at a cheaper computational cost.
+An $s$-stage FIRK method has excellent stability, but a dense Butcher tableau $a_{ij}$ and must solve a system of $n\times s$ equations at every step.
+D in ESDIRK means that $a_{ij}$ has only lower triangular and diagonal nonzeros, decoupling the system into $s$ systems of $n$ equations that are faster to solve sequentially.
+S means that all diagonal nonzeros $a_{ii}=\gamma$ are equal, so all stages share the same $\boldsymbol{W}$-matrix and only one must be factorized and can be reused as in BDF methods.
+E means that $a_{11}=0$, so the first stage is explicit and cheap.
+Unlike BDF methods, they can remain L-stable without losing order to stiffness.
+We find the 4th order KenCarp4 method (Kennedy and Carpenter, 2003) to perform well with much fewer time steps.
+Bolt.jl also defaults to this solver.
+DISCO-EB uses Kvaerno5, but we find it slower.
+  • Rosenbrock methods: Rosenbrock methods (e.g.
+Lang, 2020) linearize $\boldsymbol{f}\approx\boldsymbol{J}\boldsymbol{u}$ and effectively replace Newton’s method with one linear solve.
+Their major upside is being free from convergence on a nonlinear problem.
+This suits the perturbation ODEs, as they are already linear.
+The downside is that they can need a new $\boldsymbol{J}$ at every step, which is slow to evaluate numerically in many problems.
+But generating the Jacobian analytically eliminates this issue.
+A subclass called ?Rosenbrock-W?
+methods can reuse Jacobians, but this strategy is not yet used by OrdinaryDiffEq.jl.
+We find the 5th order L-stable Rodas5P method (Steinebach, 2023) to be both most efficient and stable on both the background and perturbations, and have not found it used by other Boltzmann codes.
+Performance of these methods is compared in section 4.
+Traditional codes use approximations to reduce stiffness and increase performance by evolving fewer ODE states when possible.
+Earlier works find marginal speedups from the TCA and UFA compared to using implicit solvers, but the RSA can significantly speed up models with high $l_{\text{max}}$ (Lesgourgues and Tram, 2011; Moser et al., 2022; Hahn et al., 2024).
+However, approximation-based codes must account for the structure of the equations to change, adding overhead to ODE solvers and restructuring sparse Jacobians.
+Approximation-free solvers can challenge this by optimizing around constant structure.
+Numerical codes approximate $\boldsymbol{J}$ with finite differences using $O(n)$ evaluations of $\boldsymbol{f}$.
+This is wasteful, as $\boldsymbol{f}=\boldsymbol{J}\boldsymbol{u}$ for linear perturbations and computing the nonzeros of $\boldsymbol{J}$ requires fewer operations than $\boldsymbol{f}$!
+Symbolic codes can take advantage of such properties.
+Approximations also complicate the code and put more load on modelers to derive, implement and validate them, and repeat the process for modifications that reintroduce stiffness or invalidate approximations.
+SymBoltz is tuned for performance without invoking approximations.
+The approximation-free structure is a major simplification and pairs well with a symbolic high-level equation-oriented interface.
+In the long run, approximation schemes can of course be explored as a secondary and optional way to maximize speed.
 2.3 Differentiability
 SymBoltz is compatible with automatic differentiation.
-One can obtain the derivative of any output quantity, such as all ODE variables and derived spectra, with respect to any combination of input parameters, like the reduced Hubble parameter $h$ or cold dark matter density parameter $\Omega_{c0}$.
-Automatic differentiation is also used several places internally, such as for computing the Jacobian for implicit ODE solvers (as an alternative to the fully symbolic approach) and in the shooting method (which is a nonlinear equation solver that wraps around the ODE solver).
-Currently SymBoltz is only well-tested with forward-mode dual numbers through ForwardDiff.jl.
-In particular, support for reverse-mode is very attractive for scalar loss applications (e.g.
-likelihoods), but left for future work (see section 4.3).
+It can compute derivatives of any output quantity (e.g.
+ODE variables, $P(k)$ or $C_{l}$) with respect to any input parameters (e.g.
+$h$ or $\Omega_{c0}$).
+For example, it can be used to learn the output’s sensitivity to the input, perform Fisher forecasts or compute the Jacobian in the shooting method when parametrizing models with integrated variables.
+It can also compute the ODE Jacobian, but we do this symbolically to get it analytically with its sparsity pattern.
+Currently, SymBoltz and Bolt.jl both work only with forward-mode dual numbers through ForwardDiff.jl [^13] https://github.com/JuliaDiff/ForwardDiff.jl.
+This is appropriate for applications with more outputs than inputs, like differentiating $O(100)$ values of $P(k)$ or $C_{l}$ as functions of $O(10)$ parameters.
+Reverse-mode is ideal with fewer outputs and attractive for MCMCs or training emulators where output is compressed to a scalar loss.
+Fast reverse-mode gradients could accelerate sampling in large parameter spaces from upcoming surveys and challenge use of emulators.
+Notably, DISCO-EB supports reverse-mode, but has not yet demonstrated it for this purpose.
+However, reverse-mode in SymBoltz is left for future work.
+3 Examples
+The best way to illustrate the features in section 2 is perhaps through some examples. The code is as of SymBoltz version 0.14.1 and available in a notebook in the project repository.
+3.1 Basic usage workflow
 Figure 2: Included plot recipes in SymBoltz make it trivial to visualize any symbolic variable or expression thereof from a solution of the Einstein-Boltzmann equations.
@@
 Wavenumbers $k$ are in units of $H_{0}/c$.
-3 Examples
-The best way to illustrate the features in section 2 is perhaps through some examples. The full code is as of SymBoltz version 0.8.1 and available in a notebook in the project repository.
-3.1 Basic usage workflow
 Most usage follows a model–problem–solution workflow: {codebox}
-using SymBoltzM = ΛCDM(lmax = 10)pars = Dict( M.γ.T₀ => 2.7, M.b.Ω₀ => 0.05, M.b.rec.Yp => 0.25, M.ν.Neff => 3.0, M.c.Ω₀ => 0.27, M.h.m_eV => 0.06, M.I.ln_As1e10 => 3.0, M.I.ns => 0.96, M.g.h => 0.7)prob = CosmologyProblem(M, pars)ks = [4, 40, 400, 4000] # k / (H₀/c)sol = solve(prob, ks)The model–problem–solution split achieves three distinct goals.
-First, a symbolic representation of the $\mathrm{\Lambda CDM}$ model M is created.
-This is a standalone object designed to be interactively inspected and modified independently of numerics (see section 3.2).
-It contains every variable, parameter and equation of the cosmological model structured as a graph of submodels for each logically distinct (physical) component: the metric $g$, the gravitational theory $G$, photons $\gamma$, massless neutrinos $\nu$, massive neutrinos $h$, cold dark matter $c$, baryons $b$, the cosmological constant $\Lambda$ and inflation $I$.
-For example, equations(M) shows all model equations; equations(M.G) shows only the gravitational ones; M.g.a gives a handle to the scale factor variable $a(\tau)$ that ?belongs?
-to the metric $g$; M.b.Ω₀ is the density parameter $\Omega_{b0}$ in the baryon component $b$; and parameters(M) lists parameters of the model that the user may set.
-Everything displays with -compatibility in notebooks to encourage interactive use.
-Second, the symbolic model is compiled to a numerical problem prob with parameters pars.
-This is an expensive operation where the ?magic?
-in section 2.1 happens: equations are checked for consistency and split into background and perturbation stages; observed and unknown variables are distinguished; fast ODE code is generated; background unknowns in the perturbations are replaced by splines; the Jacobian matrix is generated in numerical/analytical and dense/sparse form; and initial values are computed.
-One can customize whether stages should be separated, how the Jacobian should be generated, and so on.
+using SymBoltzM = ΛCDM(lmax = 16)p = Dict( M.γ.T₀ => 2.7, M.b.Ω₀ => 0.05, M.b.YHe => 0.25, M.ν.Neff => 3.0, M.c.Ω₀ => 0.27, M.h.m_eV => 0.02, M.I.ln_As1e10 => 3.0, M.I.ns => 0.96, M.g.h => 0.7)prob = CosmologyProblem(M, p; jac=true, sparse=true)ks = [4, 40, 400, 4000] # k / (H₀/c)sol = solve(prob, ks)The model–problem–solution split achieves three distinct goals.
+First, a purely symbolic representation M of the $\mathrm{\Lambda CDM}$ model is created.
+This is a standalone object designed to be interactively inspected and modified.
+It contains every variable, parameter and equation of the cosmological model structured as one submodel per physical component: the metric $g$, the gravitational theory $G$, photons $\gamma$, massless neutrinos $\nu$, massive neutrinos $h$, cold dark matter $c$, baryons $b$, the cosmological constant $\Lambda$ and inflation $I$.
+For example, equations(M) shows all equations; equations(M.G) only the gravitational ones; M.g.a refers to the scale factor variable $a(\tau)$ that ?belongs?
+to the metric $g$; M.b.Ω₀ is the baryon density parameter $\Omega_{b0}$; and parameters(M) lists parameters that can be set.
+This displays with -compatibility in notebooks and transparently shows the equations being solved.
+Second, the symbolic model is compiled to a numerical problem prob with parameters p.
+This is an expensive step where the processing in section 2.1 happens: equations are checked for consistency and split into background/perturbation stages; observed/unknown variables are distinguished; fast ODE code is generated; background unknowns in the perturbations are replaced by splines; the Jacobian matrix is generated in numerical/analytical and dense/sparse form; and initial values are computed.
+The compilation is customized with optional arguments like jac and sparse.
 This is a separate step because it performs final transformations on the model M once the user has finished modifying and committed to it.
-Third, solve(prob, ks) solves the background and perturbations for the requested wavenumbers ks.
+Third, solve(prob, ks) solves the background and perturbations for the wavenumbers ks.
 Omitting ks solves only the background.
 The resulting solution object sol provides convenient access to all model variables.
-Internally, it stores the values of all ODE unknowns at the time steps taken by the (adaptive) solvers and for every requested wavenumber.
-However, it can be queried with any time, wavenumber and symbolic expression, and will automatically compute it from the unknowns and interpolate between stored times and wavenumbers (see sections 2.1.2, 2.1.3 and 2.1.4).
-For example, calling sol(1.0, 2.0, g.Φ+g.Ψ) will compute $\Phi+\Psi$ at $k=1\,H_{0}/c$ and $\tau=2\,H_{0}^{-1}$ by expanding it in terms of unknowns according to the equations in section A.2.
-It is also possible to compute arbitrary expressions over grids of $k$ and $\tau$.
-The convenient solution interface is also integrated into plot recipes for effortless visualization, as shown in fig.
-2.
-Other Boltzmann solvers save only a hardcoded set of variables, such as only the unknowns.
-Observed variables must generally be recomputed manually, even though they are already expressed somewhere in the code.
-This is cumbersome, introduces more potential for user error and adds unnecessary friction in the modeling process.
-SymBoltz has higher ambitions than simply spitting out a table with some selected variables, and is designed to interactively make modifications to the model with minimal changes and easily inspect its impact on any output.
+Internally, it stores values of all ODE unknowns at the time steps taken by the solvers and for every requested wavenumber.
+But it can be queried with any time, wavenumber and symbolic expression, and will automatically compute it from the unknowns and interpolate between stored times and wavenumbers, as explained in section 2.1.4.
+For example, calling sol(1.0, 2.0, g.Φ+g.Ψ) computes $\Phi+\Psi$ at $k=1\,H_{0}/c$ and $\tau=2\,H_{0}^{-1}$ by expanding it in terms of unknowns using the equations in section A.2.
+This interface was used to plot the evolution of several variables in fig.
+2 with little code.
+Most Boltzmann solvers save only a fixed set of variables, such as only the unknowns.
+Recovering observed variables requires manual effort.
+This is cumbersome, opens for user error and adds friction in the modeling process.
+SymBoltz is designed to provide easy access to all variables defined by the model.
+3.2 Modifying models
+Suppose we want to replace the cosmological constant $\Lambda$ with another dark energy model, such as dynamical $w_{0}w_{a}$ dark energy (Chevallier and Polarski, 2001; Linder, 2003) with equation of state
+$w(\tau)=w_{0}+w_{a}(1-a(\tau)).$ (6)
+In this case, it is possible to solve the continuity equation
+$\rho^{\prime}(\tau)=-3\mathscr{H}(\tau)\rho(\tau)(1+w(\tau))$ (7a)
+analytically with the ansatz $\rho(\tau)\propto a(\tau)^{m}\exp(na(\tau))$ to get
+$\rho(\tau)=\rho(\tau_{0})a(\tau)^{-3(1+w_{0}+w_{a})}\exp(-3w_{a}(1-a(\tau))).$ (7b)
+Perturbations are given by de Putter et al. (2010). To implement this species in SymBoltz, write down all related variables, parameters, equations and initial conditions in one place: {codebox}
+g, τ, k = M.g, M.τ, M.ka, ℋ, Φ, Ψ = g.a, g.ℋ, g.Φ, g.ΨD = Differential(τ)@parameters w₀ wₐ cₛ² Ω₀ ρ₀@variables ρ(τ) P(τ) w(τ) cₐ²(τ) δ(τ,k) θ(τ,k) σ(τ,k)eqs = [ w ~ w₀ + wₐ*(1-a) ρ₀ ~ 3*Ω₀ / (8*Num(π)) ρ ~ ρ₀ * a^(-3(1+w₀+wₐ)) * exp(-3wₐ*(1-a)) P ~ w * ρ cₐ² ~ w - 1/(3ℋ) * D(w)/(1+w) D(δ) ~ 3ℋ*(w-cₛ²)*δ - (1+w) * ( (1+9(ℋ/k)^2*(cₛ²-cₐ²))*θ - 3*D(Φ)) D(θ) ~ (3cₛ²-1)*ℋ*θ + k^2*cₛ²*δ/(1+w) + k^2*Ψ σ ~ 0]initialization_eqs = [ δ ~ -3//2 * (1+w) * Ψ θ ~ 1//2 * (k^2*τ) * Ψ]X = System(eqs, τ; initialization_eqs, name = :X)Everything is packed down in the $w_{0}w_{a}$ component X.
+Unicode code symbols are encouraged to maximize similarity with equations (e.g.
+Ω₀ over Omega_0), and SymBoltz will automatically render them in -compatible environments (e.g.
+notebooks).
+A $w_{0}w_{a}\text{CDM}$ model is now built by replacing the cosmological constant species Λ in $\Lambda\text{CDM}$ by the $w_{0}w_{a}$ species X: {codebox}
+M = ΛCDM(Λ = X, lmax = 16, name = :w₀wₐCDM)push!(p, X.w₀ => -0.9, X.wₐ => 0.2, X.cₛ² => 1.0)prob = CosmologyProblem(M, p; jac=true, sparse=true)
+This is all the user has to do. The modification simply consists of writing down the equations verbatim. It could hardly be more compact and to the point. Under the hood, SymBoltz will:
+  • create input hooks for setting the parameters $w_{0}$, $w_{a}$, $c_{s}^{2}$, $\Omega_{0}$,
+  • move $(\tau)$-dependent functions to the background,
+  • move $(\tau,k)$-dependent functions to the perturbations,
+  • expand $\texttt{D(w)}=\mathrm{d}w/\mathrm{d}\tau$ and $\texttt{D(Φ)}=\mathrm{d}\Phi/\mathrm{d}\tau$ from $w$ and $\Phi$,
+  • compute needed background variables in the perturbations,
+  • spline $\rho(\tau)$ in the perturbations (if we use (7a) over (7b)),
+  • source gravity with energy-momentum contributions,
+  • assign $\Omega_{X0}=1-\sum_{s\neq{X}}\Omega_{s0}$ (if gravity is GR) by default,
+  • eliminate common subexpressions like x=1+wx\!=\!1+w and y=1/xy\!=\!1/x,
+  • generate ODE functions and state indices for $\delta^{\prime}$ and $\theta^{\prime}$,
+  • generate sparse analytical Jacobian entries for $\delta^{\prime}$ and $\theta^{\prime}$,
+  • interpolate and output any variable from the solution, both for unknowns (i.e. $\delta$ and $\theta$) and observeds (e.g. $w$ and $c_{a}^{2}$).
+Most of these steps require manual intervention in a numerical code.
+For example, in CLASS one would have to read new parameters in input.c; declare new background and perturbation variables in background.h and perturbations.h; solve background equations, save desired output and source gravity in background.c; and recompute or look up background variables, solve perturbation equations, save desired output and source gravity in perturbations.c.
+The changes are scattered across the code, so it grows in complexity as more models are added.
+SymBoltz opens for a workflow that implements and analyzes a modified model with less effort all in one notebook, for example.
+Of course, both codes already include a $w_{0}w_{a}$ species, but this is how they are implemented. We proceed with the $w_{0}w_{a}\text{CDM}$ model.
 Figure 3: Matter and CMB ($T$, TE and $E$) power spectra computed by SymBoltz (SB; colored lines) compared to CLASS (CL; grey dashes) with relative errors $\text{rel.err.}=P_{k}^{\text{SB}}/P_{k}^{\text{CL}}-1$ and $\text{rel.err.}=C_{l}^{\text{SB}}/C_{l}^{\text{CL}}-1$ for the $\Lambda\text{CDM}$ model.
 CLASS uses the precision parameters in appendix B.
-3.2 Modifying models
-Suppose we want to replace the cosmological constant species $\Lambda$ in the model M with another dark energy species.
-A well-known example is dynamical $w_{0}w_{a}$ dark energy (Chevallier & Polarski, 2001; Linder, 2003) with equation of state [^10] SymBoltz already includes $w_{0}w_{a}$ as an available species.
-$w(\tau)=w_{0}+w_{a}(1-a(\tau)).$ (6)
-In this case, the continuity equation
-$\rho^{\prime}(\tau)=-3\mathscr{H}(\tau)\rho(\tau)(1+w(\tau))$ (7)
-admits the analytical solution (by an ansatz of the same form)
-$\rho(\tau)=\rho(\tau_{0})a(\tau)^{-3(1+w_{0}+w_{a})}\exp(-3w_{a}(1-a(\tau))).$ (8)
-To implement this species including perturbations following de Putter et al. (2010), simply write down the symbolic variables, parameters, equations and initial conditions: {codebox}
-g, τ, k = M.g, M.τ, M.ka, ℋ, Φ, Ψ = g.a, g.ℋ, g.Φ, g.ΨD = Differential(τ)@parameters w₀ wₐ cₛ² Ω₀ ρ₀@variables ρ(τ) P(τ) w(τ) cₐ²(τ) δ(τ,k) θ(τ,k) σ(τ,k)eqs = [ w ~ w₀ + wₐ*(1-a) ρ₀ ~ 3*Ω₀ / (8*Num(π)) ρ ~ ρ₀ * a^(-3(1+w₀+wₐ)) * exp(-3wₐ*(1-a)) P ~ w * ρ cₐ² ~ w - 1/(3ℋ) * D(w)/(1+w) D(δ) ~ 3ℋ*(w-cₛ²)*δ - (1+w) * ( (1+9(ℋ/k)^2*(cₛ²-cₐ²))*θ + 3*D(Φ)) D(θ) ~ (3cₛ²-1)*ℋ*θ + k^2*cₛ²*δ/(1+w) + k^2*Ψ σ ~ 0]initialization_eqs = [ δ ~ -3//2 * (1+w) * Ψ θ ~ 1//2 * k^2*τ * Ψ]X = System(eqs, τ; initialization_eqs, name = :X)Everything is packed down into the $w_{0}w_{a}$ component X.
-Note that SymBoltz encourages Unicode symbols to maximize similarity between equations and code (e.g.
-Ω₀ over Omega_0) and to easily display in compatible environments (e.g.
-notebooks).
-This is all the user has to do!
-Variables are automatically split into and carried across the background and perturbation stages, hooks for setting input parameters and getting arbitrary output variables are automatically available, and so on.
-The modification simply consists of writing down the equations verbatim.
-It could not have been more compact and to the point.
-A similar modification to CLASS is more involved, for example.
-It would require at least: reading new parameters in input.c; declaring new background and perturbation variables in background.h and perturbations.h; solving background equations, storing desired output and coupling them to gravity in background.c; and recomputing or looking up background variables, solving perturbation equations and storing desired output and coupling them to gravity in perturbations.c.
-The changes should be if-guarded correctly to ensure the code works both with and without $w_{0}w_{a}$.
-A full $w_{0}w_{a}\text{CDM}$ model and problem can now be built by replacing the cosmological constant species Λ in $\Lambda\text{CDM}$ by the $w_{0}w_{a}$ species X: {codebox}
-M = ΛCDM(Λ = X, name = :w₀wₐCDM)pars[M.X.w₀] = -0.9pars[M.X.wₐ] = 0.2pars[M.X.cₛ²] = 1.0prob = CosmologyProblem(M, pars)We proceed with this new model and problem.
 3.3 Computing power spectra
+Table 1: Time to compute the power spectra in fig. 3.
+Code (approximations) $P(k)$ $C_{l}$
+SymBoltz (no approximations) 0.3 s 3.1 s
+CLASS (only tight-coupling approximation) 5.2 s 8.7 s
+CLASS (all approximations) 0.5 s 1.6 s
 SymBoltz can compute the matter power spectrum $P(k)$ and the angular CMB power spectra $C_{l}^{\text{TT}}$, $C_{l}^{\text{TE}}$ and $C_{l}^{\text{EE}}$: {codebox}
-using Unitful, UnitfulAstro # for Mpc unitks = 10 .^ range(-4, 1, length=200) / u"Mpc"Ps = spectrum_matter(prob, ks)ls = 10:10:2000Cls = spectrum_cmb([:TT, :TE, :EE], prob, ls)Calling spectrum_matter and spectrum_cmb with prob will automatically select a grid of $k$ to solve the perturbations for, and interpolate between them when computing the spectra.
-This is a common interpolation trick in Boltzmann solvers to integrate fewer perturbation modes.
-The functions can be called with the solution sol instead, but will then interpolate only between the $k$ that sol was solved for, leaving it to the user to ensure sufficient sampling density.
-Figure 3 shows that the spectra agree with CLASS to 11‰-11% or better for the chosen parameters.
+ks = 10 .^ range(-1, 4, length = 100)Ps = spectrum_matter(prob, ks)ls = vcat([2,3,5,10], 20:20:2500)jl = SphericalBesselCache(ls)Cls = spectrum_cmb([:TT, :EE, :TE], prob, jl)The spectrum_matter and spectrum_cmb functions select a coarse $k$-grid to solve the perturbations for and interpolates to a finer $k$-grid.
+This is a common trick to integrate fewer $k$-modes.
+Precision parameters that affect the output are given with optional keyword arguments.
+Output spectra and their computation time with standard precision settings are shown in figs.
+3 and 1.
+It shows that SymBoltz and CLASS agree to around $0.1\%$ or better.
+The $P(k)$ agree to $0.1\%$ for all $k$ and $0.01\%$ for linear $k$.
+The $C_{l}$ agree to $0.1\%$ for most $l$, but slightly worse for cosmic variance-dominated $l$ and very large $l$ in $C_{l}^{\text{TE}}$.
+This is sufficient for most current data (e.g.
+Bolliet et al., 2024).
+It is similar to CAMB vs.
+CLASS with standard precision, although their agreement is pushed to $0.01\%$ with high precision (Lesgourgues, 2011b).
+SymBoltz computes $P(k)$ quickly, but the extra steps needed for $C_{l}$ are not yet as optimized as they are in CLASS. A more thorough performance comparison for $P(k)$ is given in section 4.
 3.4 Differentiable Fisher forecasting
-Fisher forecasting is a powerful technique for predicting how strong parameter constraints that can be placed by data with given errors.
-It requires accurate derivatives and is a nice application for automatic differentiation.
-Near a peak $\boldsymbol{\theta}_{0}$, where derivatives vanish, a log-likelihood function of parameters $\boldsymbol{\theta}$ is approximated by the Taylor series
-$\log L(\boldsymbol{\theta})\approx\log L(\boldsymbol{\theta}_{0})-\sum_{i,j}F_{ij}(\theta_{i}-\bar{\theta}_{i})(\theta_{j}-\bar{\theta}_{j}),$ (9)
-where $F$ is the Fisher (information) matrix with elements
-$F_{ij}=-\frac{1}{2}\frac{\partial^{2}\log L}{\partial\theta_{i}\,\partial\theta_{j}}\bigg\rvert_{\,\boldsymbol{\theta}=\boldsymbol{\theta}_{0}}.$ (10)
-Intuitively, $F$ measures how sharp the peak is, or how sensitive $\log L$ is to changes in different directions in parameter space.
-Fisher forecasting is powered by the Cramér-Rao bound, which guarantees that $F^{-1}_{ij}$ is a lower bound $\big\lvert C_{ij}\big\rvert\geq\big\lvert F^{-1}_{ij}\big\rvert$ for the covariance $C_{ij}$ between model parameters $\theta_{i}$ and $\theta_{j}$.
-The inequality is more saturated the better the likelihood approximation (9) is (i.e.
-the ?more Gaussian?
-the probability distribution is, for which the expansion is exact).
-Thus, computing $F_{ij}$ at a peak and inverting it gives the tightest parameter constraints one can hope for.
-Since $F_{ij}$ depends on derivatives of $\log L$, Fisher forecasting traditionally involves error-prone finite differences and step size tuning.
-This problem is avoided with automatic differentiation.
+Fisher forecasting is a technique to predict how strong parameter constraints that can be placed by data with given errors. It needs derivatives that can be computed with automatic differentiation.
+Near a peak $\boldsymbol{p}_{0}$, where derivatives vanish, a log-likelihood $\log L$ of parameters $\boldsymbol{p}$ is approximated by the Taylor series
+$\log L(\boldsymbol{p})\approx\log L(\boldsymbol{p}_{0})-\sum_{i,j}F_{ij}(\boldsymbol{p}_{0})(p_{i}-\bar{p}_{i})(p_{j}-\bar{p}_{j}),$ (8)
+where $\boldsymbol{F}$ is the Fisher (information) matrix with elements
+$F_{ij}(\boldsymbol{p})=-\frac{1}{2}\frac{\partial^{2}\log L(\boldsymbol{p})}{\partial p_{i}\,\partial p_{j}}.$ (9)
+Intuitively, $\boldsymbol{F}$ measures how sharp the peak is, or how sensitive $L$ is in different directions in parameter space.
+Fisher forecasting is powered by the Cramér-Rao bound $\big\lvert\smash{C_{ij}}\big\rvert\geq\big\lvert\smash{F^{-1}_{ij}}\big\rvert$ for the covariance $C_{ij}$ between parameters $p_{i}$ and $p_{j}$.
+It is an equality for a Gaussian, for which the likelihood (8) is exact.
+Thus, inverting $\boldsymbol{F}(\boldsymbol{p}_{0})$ gives the tightest parameter constraints one can hope for.
 To demonstrate differentiable Fisher forecasting with SymBoltz, we make the best possible CMB (TT) measurement: one of $\bar{C}_{l}$ over the full sky with errors only due to cosmic variance
 $\sigma_{l}=\sqrt{\frac{2}{2l+1}}\bar{C}_{l}.$ (10)
-Assuming the measurements of each $\bar{C}_{l}$ are normally distributed and uncorrelated, the log-likelihood for this experiment is $\chi^{2}$:
-$\log L(\boldsymbol{\theta})=-\frac{1}{2}\sum_{l}\left(\frac{C_{l}(\boldsymbol{\theta})-\bar{C}_{l}}{\sigma_{l}}\right)^{2}.$ (12)
-In this case, the Fisher matrix (10) becomes
-$F_{ij}=\sum_{l}\frac{\partial C_{l}}{\partial\theta_{i}}\frac{1}{\sigma_{l}^{2}}\frac{\partial C_{l}}{\partial\theta_{j}}\bigg\rvert_{\,\boldsymbol{\theta}=\boldsymbol{\theta}_{0}},$ (13)
-evaluated in some fiducial cosmology $\boldsymbol{\theta}=\boldsymbol{\theta}_{0}$.
-The derivatives $\partial C_{l}/\partial\theta_{i}$ are hard to compute and perfect candidates for automatic differentiation with ForwardDiff.jl: {codebox}
-using ForwardDiff: jacobianvary = [ M.g.h, M.c.Ω₀, M.b.Ω₀, M.b.YHe, M.ν.Neff, M.h.m_eV, M.X.w₀, M.X.wₐ, M.I.ln_As1e10, M.I.ns,]genprob = parameter_updater(prob, vary)ls, ls′ = 100:1:1000, 100:25:1000Cl(θ) = spectrum_cmb(:TT, genprob(θ), ls, ls′)θ₀ = map(par -> pars[par], vary)dCl_dθ = jacobian(Cl, θ₀)Here vary orders the subset of parameters to differentiate with respect to, and genprob is a function that generates a new problem with updated parameters θ.
-Then Cl specifies a function that computes $C_{l}(\boldsymbol{\theta})$ for ls, interpolating from a coarser grid ls′.
-This machinery just converts from SymBoltz’ parameter-to-value mapping to a function of an array of parameter values, as jacobian requires a pure mathematical function $\boldsymbol{f}:\mathbb{R}^{a}\rightarrow\mathbb{R}^{b}$ to differentiate.
-Finally $\partial C_{l}/\partial\theta_{i}$ is computed with forward-mode automatic differentiation and stored in dCl_dθ, as shown in fig.
+Assuming a $\chi^{2}$-log-likelihood
+$\log L(\boldsymbol{p})=-\frac{1}{2}\sum_{l}\bigg(\frac{C_{l}(\boldsymbol{p})-\bar{C}_{l}}{\sigma_{l}}\smash{\bigg)^{2}},$ (11)
+the Fisher matrix (9) becomes
+$F_{ij}=\sum_{l}\frac{\partial C_{l}}{\partial p_{i}}\frac{1}{\sigma_{l}^{2}}\frac{\partial C_{l}}{\partial p_{j}}.$ (12)
+The derivatives $\partial C_{l}/\partial p_{i}$ are usually found with error-prone finite differences by carefully tuning the step.
+SymBoltz avoids this problem altogether with automatic differentiation: {codebox}
+vary = [ M.g.h, M.c.Ω₀, M.b.Ω₀, M.b.YHe, M.ν.Neff, M.h.m_eV, M.X.w₀, M.X.wₐ, M.I.ln_As1e10, M.I.ns,]genprob = parameter_updater(prob, vary)jl, ls = SphericalBesselCache(100:25:1000), 100:1000Cl(p) = spectrum_cmb(:TT, genprob(p), jl, ls)p₀ = map(par -> p[par], vary)dCl_dp = ForwardDiff.jacobian(Cl, p₀)Here vary orders the parameters to differentiate with respect to, and genprob generates a new problem with updated parameters p.
+This prepares a vector-to-vector function Cl(p) that is differentiated by ForwardDiff.jacobian to compute $\partial C_{l}/\partial p_{i}$ with dual numbers, as shown in fig.
 4.
-Figure 4: Normalized derivatives $(\partial C_{l}/\partial\theta_{i})/C_{l}$ of a CMB $T$ power spectrum with respect to cosmological parameters $\theta_{i}$ from SymBoltz and automatic differentiation (AD; colored lines) versus CLASS and central finite differences (FD; gray dashes).
+Computing and inverting the Fisher matrix (12) forecasts the constraints in fig.
+5.
+They are in excellent agreement with finite difference results from CLASS, but these required significant tuning of precision parameters and step sizes.
+Differentiable Fisher forecasts have also been done with an Eisenstein-Hu transfer function fit by Campagne et al. (2023), and through an Einstein-Boltzmann solver by Hahn et al. (2024).
+Figure 4: Normalized derivatives $(\partial C_{l}/\partial p_{i})/C_{l}$ of a CMB $T$ power spectrum with respect to cosmological parameters $p_{i}$ from SymBoltz and automatic differentiation (AD; colored lines) versus CLASS and central finite differences (FD; gray dashes).
 CLASS uses the precision parameters in appendix B and finite differences with $5\%$ relative step sizes.
 Figure 5: Marginalized 68% and 95% 2D confidence ellipses for parameter constraints from a Fisher forecast on a cosmic variance-dominated CMB $T$-only survey using the derivatives in fig. 4.
-It is now trivial to compute the Fisher matrix (13).
-Inverting it forecasts the constraints in fig.
-5.
-They are in excellent agreement with finite difference results from CLASS, but these required significant tuning of precision parameters and step sizes.
 3.5 Differentiable MCMC sampling with supernova data
-Finally, we show that SymBoltz can output differentiable results for gradient-based Markov chain Monte Carlo (MCMC) samplers, like the No-U-Turn Sampler (NUTS).
-This state-of-the-art method uses likelihood gradients to move more intelligently in parameter space than the ?blind?
+Finally, we show that SymBoltz can output differentiable results for gradient-based MCMC samplers, like the No-U-Turn Sampler (NUTS).
+It uses gradients to sample parameter space more efficiently than the ?blind?
 Metropolis-Hastings algorithm.
-We predict apparent magnitudes $m(z)$ of Type Ia supernovae as a function of redshift $z$ and fit them to 1048 observations $(z_{i},m_{i})$ from the Pantheon dataset (Scolnic et al., 2018).
-Such supernovae are standard candles with constant absolute magnitude $M\approx-19.3$.
-Their apparent magnitudes are $m(z)=M+\mu(z)$, where $\mu(z)=5\lg(d_{L}(z)/10\text{ pc})$ is the distance modulus from the background-derived luminosity distance [^11] This expression is valid for any $\Omega_{k0}$ with complex $\operatorname{sinc}(x)=\sin(x)/x$, but can be split into branches for positive, negative and zero $\Omega_{k0}$.
+We predict apparent magnitudes $m(z)=M+\mu(z)$ of Type Ia supernovae at redshift $z$ and fit to 1048 Pantheon observations $(z_{i},m_{i})$ by Scolnic et al.
+(2018).
+Their standard absolute magnitude is $M\approx-19.3$, and $\mu(z)=5\lg(d_{L}(z)/10\text{ pc})$ is the distance modulus from the background-derived luminosity distance [^14] This expression is valid for any $\Omega_{k0}$ with complex $\operatorname{sinc}(x)=\sin(x)/x$, but can be split into branches for positive, negative and zero $\Omega_{k0}$.
 $d_{L}(z)=\frac{c}{H_{0}}\frac{\chi(z)}{a(z)}\operatorname{sinc}\Big(H_{0}\sqrt{-\Omega_{k0}}\,\chi(z)\Big).$ (13)
-We define the likelihood by a multivariate normal distribution of the 1048 supernovae with (constant) covariance matrix $C$ given in detail by Scolnic et al.
+We define the likelihood $L$ by a multivariate normal distribution of the 1048 supernovae with (constant) covariance matrix $C$ given in detail by Scolnic et al.
 (2018).
-To compute it, we interface SymBoltz with the probabilistic programming framework Turing.jl (Fjelde et al., 2025), which allows a high-level description of the probabilistic model equivalent to the log-likelihood
-$\log L=-\frac{1}{2}\sum_{i,j}\big(m(z_{i})-m_{i}\big)C^{-1}_{ij}\big(m(z_{j})-m_{j}\big).$ (15)
+To compute $L$, we interface SymBoltz with the probabilistic programming framework Turing.jl (Fjelde et al., 2025), which allows a high-level description of the probabilistic model equivalent to the log-likelihood
+$\log L=-\frac{1}{2}\sum_{i,j}C^{-1}_{ij}\big(m(z_{i})-m_{i}\big)\big(m(z_{j})-m_{j}\big).$ (14)
 We use the NUTS sampler in Turing.
@@
 6.
-The results are not interesting, but show that differentiable parameter inference works with SymBoltz.
-Performance improvements are needed to make this viable also for perturbation-derived spectra.
-Figure 6: Parameter inference on 1048 Type Ia supernovae from the full Pantheon supernova dataset, using 5000 MCMC samples with the gradient-based NUTS sampler in Turing.jl and differentiable theory predictions from SymBoltz.jl.
-4 Discussion of design synergies and tradeoffs
-A symbolic-numeric interface, lack of approximations and differentiability are interesting features in their own rights.
-They also constitute a self-reinforcing synergy explained in fig.
-7.
-Like any Boltzmann solver, SymBoltz makes some tradeoffs in its design.
-Symbolic-numericinterfaceFree ofapprox-imationsAutomaticdifferentiationDetect sparsity of JJCalculate $J$ robustlyEasy to specify model
-Figure 7: A symbolic-numeric interface, approximation-freeness and differentiability are three important features of SymBoltz that form a synergy.
-For example: automatic differentiation helps to calculate the ODE Jacobian robustly; which is needed by implicit solvers to integrate stiff ODEs without approximations; which makes it easier to write down models in a simple symbolic form; which can be used to detect the sparsity pattern of the Jacobian; which in turn increases the efficiency of the implicit ODE solver.
-This synergy creates a self-reinforcing design.
-4.1 Symbolic vs. numeric interface
-As explained in sections 1.1 and 2.1, modifications to most Boltzmann solvers are made by editing the low-level numerical source code directly, while changes to SymBoltz are made in a high-level symbolic interface that abstracts away internal details.
-We think three properties make the linearized Einstein-Boltzmann equations particularly attractive for symbolic abstraction.
-First, on the ?outside?, they are fundamentally one (large) system of equations with a predictable general structure.
-Second, on the ?inside?, they are complicated to solve in multiple stages due to their structure, stiffness and need for speed.
-Third, they are heavily subject to modifications as the true cosmological model is unknown.
-To simplify the modeling process, these aspects make it valuable to abstract the ?inside?
-from the ?outside?.
-Purely numerical codes generally put the programmer first.
-They are adapted to the computational structure of the problem.
-Users have full freedom to tweak problem-specific details, such as implementing complex approximation schemes.
-Without abstraction layers, the code says exactly what it does.
-This can be worthwhile for custom-tailoring codes to ?workhorse?
-cosmological models like $\mathrm{\Lambda CDM}$ for maximum performance and demanding MCMC analyses.
-However, it can result in an overwhelming monolithic design that is hard to read and modify.
-The symbolic interface in SymBoltz prioritizes the modeler.
-Its design matches the physical structure of the model.
-Manual chores are automated with compile-time logic on the equations.
-Users can more easily change equations without understanding internals provided that they can be written in a straightforward form that is compatible with the symbolic interface.
-This is not a major restriction for the Einstein-Boltzmann equations, as they have a plain and predictable structure when linearized and free of approximations.
-Nevertheless, to bridge this gap, SymBoltz lets users call arbitrary numerical functions that escape the symbolic-to-numeric compilation, if necessary.
-This abstraction layer works only if fits the structure of the underlying equations.
-A key to envision the design was to look past the computational pipeline structure of Boltzmann solvers (e.g.
-background $\rightarrow$ perturbations $\rightarrow\ldots$), and rather organize the code primarily by physical components like the metric, gravitational theory and particle species.
-Components are self-contained and joined into full cosmological models.
-As more models are added, the complexity of this modular design stays constant, while monolithic ?all-in-one?
-codes become complicated.
-This structure can be unfamiliar to those used to the layout of other codes.
-The modular structure scales best with a growing number of submodels for individual physical components.
-But SymBoltz also provides a full unstructured $\mathrm{\Lambda CDM}$ model where everything is merged into one large system, which is easier to work with if one wants to modify arbitrary small parts of the equations.
-This leaves the choice to the user, who can select between the best of both worlds.
-Mixing symbolic and numerical computing can be susceptible to performance problems.
-Numerics should be fast and symbolics is usually slow, so ?perceived performance?
-suffers if the two are interleaved.
-But SymBoltz disentangles the expensive symbolic operations in section 2.1 from numerics, so they are performed only before and after performance-critical tasks.
-We hope the design of SymBoltz provides a platform for easy exploration of alternative models, regardless of which sector of the equations one is interested in modifying.
-4.2 Approximation-freeness vs. performance
-As seen in section 1.2, traditional codes reduce stiffness by approximating the equations for explicit integrators, while SymBoltz solves the full stiff system with implicit methods.
-Without approximations, SymBoltz needs only one set of equations to solve.
-These equations are easy to write down in pure symbolic form, which pairs nicely with SymBoltz’ high-level symbolic interface.
-With approximations, traditional solvers need more complicated infrastructure to switch between versions of the same equations.
-This fits well to a low-level and fully numerical code that maximizes implementation freedom, like CLASS.
-Approximation schemes do not only alleviate stiffness, but also improve performance by reducing the ODE size.
-However, Lesgourgues & Tram (2011); Hahn et al.
-(2024) found that the TCA and UFA provide only marginal speedups compared to use of implicit solvers.
-On the other hand, the RSA can provide significant speedups with high $l_{\text{max}}$ (Lesgourgues & Tram, 2011; Moser et al., 2022).
-However, more load is put on the modeler to derive, implement and validate multiple approximations.
-This process must generally be repeated for modified models that can easily reintroduce stiffness or invalidate the approximations.
-Implicit solvers take more expensive steps than explicit ones.
-They solve a nonlinear system for the next unknowns at every time step, which gets costly as the ODE size grows (e.g.
-higher $l_{\text{max}}$).
-This is often done with Newton’s method, which iterates over linear matrix solutions $A\boldsymbol{x}=b$, where $A$ involves the ODE Jacobian $J$.
-But several factors mitigate this slowdown.
-First and foremost, implicit solvers take longer and fewer steps due to better stability properties.
-This is highly solver-dependent, however: two different implicit solvers can perform very differently on one stiff problem.
-DifferentialEquations.jl lets SymBoltz easily switch between a large suite of implicit solvers.
-The background (with stiff thermodynamics) is solved robustly with Rodas4P/Rodas5P, and the perturbations efficiently with TRBDF2/KenCarp4/Rodas5P for low/medium/high precision.
-Notably, CLASS includes and defaults to a custom implicit solver ndf15 and performs great (Lesgourgues & Tram, 2011).
-In contrast, SymBoltz can choose between a range of compatible solvers to fulfill different performance and precision requirements.
-Second, a key optimization for implicit solvers is to compute $J$ efficiently, reuse it over time steps and LU-factorize $A$ to speed up successive linear system solutions until Newton’s method no longer converges due to outdated Jacobians.
-The TRBDF2 and KenCarp4 solvers do this.
-When needed, $J$ is updated with the method in section 2.2, which is cheaper than finite differences.
-The linear matrix solver used in Newton’s method also matters, as they are among the most thoroughly optimized numerical algorithms.
-DifferentialEquations.jl interoperates with LinearSolve.jl, which lets SymBoltz easily swap the linear solver in the ODE solver.
-On our hardware we find that performance improves by $5\times$ with a recursive LU factorization algorithm from RecursiveFactorization.jl or Intel’s (proprietary) Math Kernel Library MTK.jl over Julia’s default OpenBLAS backend.
-Third, sparse matrix methods can speed up the linear solver as the system grows in size and the fraction of zeros increases.
-SymBoltz can find the exact sparsity pattern from the symbolic Jacobian in section 2.1.5, and LinearSolve.jl provides several sparse matrix methods.
-Support for this is still ongoing work.
-Although SymBoltz sacrifices some performance to get rid of approximations, these countermeasures make the impact less severe than one might fear.
-In return, the approximation-free structure is a major simplification and pairs well with the high-level interface for simple symbolic equations.
-There is still room to improve performance without resorting to approximations by tuning precision tolerances, supporting sparse matrix methods, exploring implicit-explicit (IMEX) solvers (e.g.
-Kennedy & Carpenter (2003)) and smarter sampling of $k$ and $l$, for example.
-Further performance optimizations and comparisons are left for future work.
-SymBoltz is fast enough for single runs and interactive usage, but not yet for MCMC analyses with perturbation-derived quantities.
-Of course, SymBoltz can also implement approximation schemes in the future, but its symbolic nature can make this harder than in other codes.
-As a counterweight to other codes, SymBoltz’ primary mission will be to solve full stiff equations and keep approximations optional and secondary.
-4.3 Forward-mode vs. reverse-mode automatic differentiation
-At this time SymBoltz is only well-tested with forward-mode automatic differentiation.
-However, we saw in section 1.3 that reverse-mode is more attractive for applications with more inputs than outputs.
-This is the case for popular applications with scalar likelihood or loss functions, such as MCMC parameter inference and training neural network emulators.
-Reverse-mode could be particularly powerful for parameter inference with next-generation surveys, where experimental nuisance parameters must also be sampled in large $\mathcal{O}(100)$-dimensional parameter spaces, even if eventually marginalized over.
-However, forward-mode has better characteristics when performing sensitivity analyses.
-For example, computing $\partial P(k;\boldsymbol{\theta})/\partial\theta_{i}$ or $\partial C_{l}(\boldsymbol{\theta})/\partial\theta_{i}$ (e.g.
-Fisher forecasting in section 3.4) is faster with forward-mode, since one typically wants $\mathcal{O}(100\text{-}1000)$ $k$ or $l$, but only has $\mathcal{O}(10)$ parameters.
-Of course, automatic differentiation is just an additional feature.
-One can run the code with or without automatic differentiation, or use finite differences instead.
-It does not pose any tradeoff, but robust support for both forward-mode and reverse-mode remains a future goal that would make SymBoltz more powerful for different applications.
-This situation will improve as the differentiable Julia ecosystem continues to evolve.
-5 Conclusion and future potential
-SymBoltz is a fresh Julia package for solving the linearized Einstein-Boltzmann equations.
-It relaxes all approximation switching schemes found in other codes and solves a single set of stiff equations at all times with implicit integrators, and combats the performance loss with modern and efficient implicit ODE solvers, optimized linear system solvers and (soon) sparse matrix methods.
+The differentiable computation is not yet fast enough for MCMCs with perturbation-derived spectra.
+The results here are not groundbreaking, but show that the differentiable pipeline is functional.
+This is a first step to differentiating the solver quickly, and we aim to accelerate this computation with future work.
+Differentiable Boltzmann codes have yet to demonstrate this ability, but more advanced MCMCs with a differentiable background have been done by Campagne et al. (2023), for example.
+Figure 6: Parameter inference on 1048 Type Ia supernovae from Pantheon data, using 5000 MCMC samples with the gradient-based NUTS sampler in Turing.jl and differentiable predictions from SymBoltz.
+4 Performance comparison to CLASS
+Figure 7: Work vs.
+precision for computing the matter power spectrum $P(k)$ with SymBoltz and CLASS.
+SymBoltz is approximation-free and run with several implicit ODE solvers.
+CLASS is run with approximations on/off, its implicit/explicit evolvers ndf15/rk and precision settings in appendix B, but tight coupling is always approximated.
+For each setup, $P(k)$ is computed by integrating perturbation $k$-modes with ODE solver tolerances $10^{-2}$-$10^{-9}$ that controls the adaptive stepping.
+For each tolerance, we record the best of 3 runtimes and the $L_{2}$-compressed error $\smash{\big(\sum_{i=1}^{N}\!\big(P(k_{i})/\bar{P}(k_{i})-1\big)^{2}/N\big)^{1/2}}$ relative to high-precision spectra $\bar{P}(k)$ from both codes with approximations off, tolerance $10^{-10}$ and FBDF/ndf15.
+Both codes solve a $w_{0}w_{a}\text{CDM}$ model with equal parameters, $N=114$ equal wavenumbers $k_{i}$, multipole cutoff $l_{\text{max}}=16$ for all species, default sampling of massive neutrino momenta and parallelize over $k$.
+Times are from a Linux laptop with Intel i7-12800H CPU.
+Figure 7 compares performance versus precision of the matter power spectrum $P(k)$ computed with SymBoltz and CLASS.
+This computation solves the background, thermodynamics and perturbations for several $k$ and reads off $P(k)$ at the final time.
+The work is dominated by the perturbations, so it is a relevant test of SymBoltz’ approximation-free treatment of perturbations.
+The comparison shows how approximations and the ODE solver affect integration of the Einstein-Boltzmann equations.
+Without approximations, SymBoltz with Rodas5P is more than $10\times$ faster than CLASS with its implicit ndf15 evolver at comparable precision.
+When CLASS uses approximations and explicit evolver, SymBoltz is as fast while remaining approximation-free.
+To make this comparison as fair as possible, we configured both codes to integrate identical $k$-modes (decided by CLASS) without $k$-interpolation and use the same multipole cutoff $l_{\text{max}}$ for all species.
+However, the codes sample massive neutrino momenta $q$ differently.
+This greatly impacts the number of perturbation equations (see section A.7).
+SymBoltz’ 4 default momenta gives 127 perturbation equations in total.
+CLASS’ default $\texttt{tol\_ncdm\_newtonian}=10^{-5}$ gives 11 momenta and 245 equations (before approximations).
+Increasing this tolerance to $10^{-3}$ gives 5 momentum bins, 143 equations and is around $245/143\approx 1.7$ times faster, but introduces errors beyond $1\%$ in $P(k)$.
+We therefore leave both codes with default $q$-sampling, but note that CLASS could be up to $2\times$ faster at unchanged precision with sparser sampling.
+This would shift its curves down by $\lg 2\approx 0.3$.
+Howlett et al.
+(2012) (in appendix A) also find a precise 3-4 point scheme in CAMB, and Lee et al.
+(2025) speed up CLASS using integral equations for massive neutrinos in the fork CLASSIER.
+Despite lack of approximation schemes, SymBoltz is fast due to high-order implicit solvers like Rodas5P that take accurate and long steps, fast functions for $\boldsymbol{f}$ and the analytically generated $\boldsymbol{J}$, sparse matrix methods with a constant predetermined sparsity structure and efficient sampling of massive neutrino momenta.
+SymBoltz is currently slower than CAMB and CLASS for computing $C_{l}$.
+This is subject to more optimizations than $P(k)$, such as sampling of the source function $S(\tau,k)$, interpolation in $\tau$, $k$, and $l$, evaluation of spherical Bessel functions $j_{l}(x)$ and optimized line-of-sight quadrature.
+These are very well polished by CAMB and CLASS, and unmatched by SymBoltz at this time.
+A detailed performance analysis for $C_{l}$ is therefore not meaningful now (but see table 1).
+With more work, the $C_{l}$ optimizations can be added to SymBoltz.
+The purpose of the comparison made here is to show the viability of approximation-free perturbations, which was a major milestone in designing SymBoltz.
+5 Future work
+We hope SymBoltz continues to grow and that others can build on it. Its symbolic capabilities can be extended with utilities for:
+  • separation into finer computational stages (see section 2.1.3);
+  • gauge transformation of equations (e.g. Newtonian $\leftrightarrow$ synchronous gauge), as already possible in CAMB [^15] https://camb.readthedocs.io/en/latest/symbolic.html;
+  • perturbation theory utilities for deriving initial conditions and approximation schemes for new models;
+  • transformation of differential equations with respect to conformal time $\tau$ to another (one-to-one) independent variable (e.g. to get observables as direct functions of redshift $z$);
+  • validation of dimensions in symbolic equations to catch modeling mistakes by tagging variables with physical units;
+  • symbolic tensor manipulation to aid work with modified gravity theories (as a standalone package).
+These tasks traditionally involve manual error-prone calculations, so such utilities could ease exploration of extended models.
+We particularly aim to improve performance for differentiable runs to the level of standard runs.
+An attractive milestone is fast reverse-mode gradients of scalar loss functions for gradient-based MCMCs.
+We anticipate the symbolic structure and Jacobian generation to be useful for this, as writing $\boldsymbol{u}(\tau,\boldsymbol{p})$ in the ODE (3) and differentiating it for $\boldsymbol{v}_{i}=\partial\boldsymbol{u}/\partial p_{i}$ gives $\mathrm{d}\boldsymbol{v}_{i}/\mathrm{d}\tau=\boldsymbol{J}\boldsymbol{v}_{i}+\partial\boldsymbol{f}/\partial p_{i}$, which could be constructed analytically.
+It is left for future work to generalize the code from scalar to vector and tensor perturbations, from flat to curved spacetimes, and from adiabatic to other initial conditions.
+Some additional models like quintessence dark energy and Brans-Dicke gravity are in the works, and we invite others to add extended models.
+We hope SymBoltz can foster growth of modular interoperating packages for cosmology in Julia.
+For example, applications like higher-order perturbation theory (e.g.
+CLASS-PT by Chudaykin et al.
+(2020) and PyBird by D’Amico et al.
+(2021)), nonlinear boosting (e.g.
+HALOFIT by Takahashi et al.
+(2012) and HMCODE by Mead et al.
+(2021)), CMB lensing (e.g.
+CMBLensing.jl by Millea et al.
+(2020)) and nonlinear $N$-body simulations all rely on output from Boltzmann solvers.
+Supporting automatic differentiation across language barriers is hard, and an ecosystem of such packages in one language with support for automatic differentiation, like Julia or Python with JAX, could create a modular and differentiable cosmological modeling toolbox.
+A recent example of this is DISCO-DJ’s coupling of differentiable Boltzmann and $N$-body codes for end-to-end modeling from linear perturbations to non-linear structure formation (Hahn et al., 2024; List et al., 2025).
+6 Conclusion
+SymBoltz is a fresh Julia package for solving the linear Einstein-Boltzmann equations.
+It features a symbolic-numeric interface where models are built from symbolic equations and compiled to efficient numerical code.
+This lets modelers prototype new models with few lines code at a high level.
+It relaxes all approximation schemes found in older codes by solving a single set of stiff equations at all times with implicit integrators.
+This simplifies the equations and remains performant due to high-order implicit solvers with fast generated code for ODEs and analytical and sparse Jacobians.
 It is differentiable, so one can get accurate derivatives of any output quantity with respect to any input parameter.
-This enables modelers to rapidly prototype new models by straightforwardly writing down new variables and equations.
-Version 0.8.1 features the metric in the conformal Newtonian gauge, General Relativity, cold dark matter, photons, baryons and RECFAST recombination, massless and massive neutrinos, the cosmological constant and $w_{0}w_{a}$ dark energy, and computes luminosity distances and matter and CMB spectra.
-It also has some rudimentary models for Brans-Dicke gravity, quintessence dark energy and curved geometry, but these are not complete yet.
-We think the modular design makes it very easy to add new models and compute other quantities.
-Forward-mode automatic differentiation is well-tested, while support for reverse-mode is a future goal. This would make SymBoltz very powerful for scalar loss applications like MCMCs.
-With some more work, SymBoltz could grow into a fully integrated symbolic-numeric and differentiable cosmological modeling environment.
-For example, numerical code generation could be extended from linear equations to non-linear $N$-body simulations in a consistent and unified framework.
-This could alleviate the $N$-language problem ($N\gtrsim 2$) prevalent in cosmological modeling, which glues together convenient high-level languages like Python with performant low-level languages like C and Fortran, and sometimes symbolic work in a CAS like Mathematica.
-SymBoltz shows that it is possible to create a symbolic-numeric, approximation-free and differentiable Boltzmann solver with modern numerical techniques.
-It excels at model simplicity and ease of modification.
-More work is needed to make it as feature-complete and fast as CAMB and CLASS that have been refined over many years.
-A major goal is to compute values and derivatives of perturbation-derived spectra fast enough for use with gradient-based MCMC samplers.
-Implicit ODE solvers, nonlinear and linear equation solvers, sparse matrix methods and differentiability are all demanding numerical techniques.
-Making them work robustly and efficiently together is challenging, but the current state shows promise.
-SymBoltz is built on (and has contributed to) several evolving Julia packages, and will continue to grow both on its own and with improvements to its dependencies.
+Output power spectra agree with CLASS to $0.1\%$ with standard precision and can be used fit to current data.
+Symbolic-numericinterfaceApproximation-freenessDifferentiabilityDetect sparsity of $\boldsymbol{J}$Calculate $\boldsymbol{J}$ robustlyEasy to specify model
+Figure 8: A symbolic-numeric interface, approximation-freeness and differentiability are three main features of SymBoltz that form a synergy.
+Differentiability computes the ODE Jacobian $\boldsymbol{J}$ accurately and efficiently; which is used by implicit solvers to integrate stiff ODEs without approximations; which makes it easier to write down equations in simple symbolic form; which is used to find the sparsity pattern of $\boldsymbol{J}$; which speeds up $\boldsymbol{J}$ and the implicit solver.
+This is a self-reinforcing design.
+Other recent codes like PyCosmo, Bolt.jl and DISCO-EB also incorporate some of these features.
+SymBoltz combines them and reinforces its design with a synergy between them, as explained in fig.
+8.
+Together, these new codes offer a simpler alternative in high-level languages that complements traditional codes like CAMB and CLASS that are historically highly tuned around approximations in low-level languages.
+However, new codes need more work to reach the level of features and maturity of CAMB and CLASS that have grown over many years.
+SymBoltz version 0.14.1 includes the flat Newtonian gauge metric, General Relativity, cold dark matter, photons, baryons and RECFAST recombination, massless and massive neutrinos, the cosmological constant and $w_{0}w_{a}$ dark energy with scalar perturbations, distances and matter and CMB spectra.
+Forward-mode automatic differentiation works for single-run purposes like Fisher forecasting.
+Fast reverse-mode scalar loss gradients, particularly through the perturbations, is a future milestone that could challenge emulators in gradient-based MCMC sampling.
+No differentiable Boltzmann code is yet capable of this.
 SymBoltz is easy to install from https://github.com/hersle/SymBoltz.jl.
-Documentation is available there, and the code is tested with continuous integration to ensure it remains correct and up-to-date (see appendix C).
-Anyone is encouraged to ask questions, give feedback, open issues and contribute pull requests in the repository!
-We hope SymBoltz offers valuable competition and new ideas on the Boltzmann solver market.
+Documentation [^16] https://hersle.github.io/SymBoltz.jl/ is available, and the code is tested with continuous integration to remain correct and up-to-date (see appendix C).
+All questions and contributions are welcome in the repository!
+We hope SymBoltz proves useful and that it offers valuable ideas on the Boltzmann solver market.
 Acknowledgments
-I thank Aayush Sabharwal and Christopher Rackauckas for developing and maintaining ModelingToolkit.jl and DifferentialEquations.jl and answering questions.
+I thank the Julia community for creating an open ecosystem of scientific packages underlying SymBoltz.
+I thank Aayush Sabharwal and Christopher Rackauckas for developing ModelingToolkit.jl and OrdinaryDiffEq.jl and answering questions.
 I thank Hans A.
-Winther for helpful suggestions, testing the code and giving feedback on it and this manuscript.
-I thank Julien Lesgourgues, Thomas Tram and others for developing CLASS, which has inspired SymBoltz.
+Winther for helpful suggestions and feedback on the code and drafts of this paper.
+I thank Julien Lesgourgues, Thomas Tram and others for developing and documenting CLASS.
 Appendix A List of equations and practical implementation details
@@
 When $G$, $c$ or and $H_{0}$ appear explicitly in the equations, they are in SI units and used only to convert from SI units into the dimensionless units.
-The equations closely follow the conventions in the seminal paper by Ma & Bertschinger (1995) and very closely matches the source code of SymBoltz with Unicode characters.
+The equations closely follow the conventions in the seminal paper by Ma and Bertschinger (1995) and very closely matches the source code of SymBoltz with Unicode characters.
 It is far outside the scope of this paper to derive the equations and explain the meaning of every variable.
 The independent variable is conformal time $\tau$, and all derivatives $^{\prime}=\mathrm{d}/\mathrm{d}\tau$ are with respect to it (in units of $1/H_{0}$).
+This is perhaps the most common parametrization in the literature, and is natural because most equations are autonomous with respect to $\tau$ (i.e.
+$f(\tau,\boldsymbol{u})\rightarrow f(\boldsymbol{u})$, except some multipole truncation schemes that use $1/(k\tau)$, but these are unphysical).
 By default, integration starts from the early time $\tau=\tau_{i}=10^{-6}$, and is terminated when the scale factor crosses $a=1$, and the corresponding time $\tau=\tau_{0}$ is labeled today.
@@
 This significantly complicates their behavior.
-SymBoltz currently implements equations from RECFAST [^12] https://www.astro.ubc.ca/people/scott/recfast.html version 1.5.2 (Seager et al., 1999, 2000; Wong et al., 2008; Scott & Moss, 2009):
+SymBoltz currently implements equations from RECFAST [^17] https://www.astro.ubc.ca/people/scott/recfast.html version 1.5.2 (Seager et al., 1999; Wong et al., 2008; Scott and Moss, 2009):
 $w=0,\qquad P=0,\qquad\rho=\frac{\rho_{0}}{a^{3}},\qquad{f_{\mathrm{He}}}=\frac{{Y_{\mathrm{He}}}}{\frac{m_{\mathrm{He}}}{m_{\mathrm{H}}}\big(1-{Y_{\mathrm{He}}}\big)},\qquad n_{\mathrm{H}}=\frac{(1-Y_{\mathrm{He}})\rho}{m_{\mathrm{H}}},\qquad n_{\mathrm{He}}=f_{\mathrm{He}}n_{\mathrm{H}},\qquad{c_{s}^{2}}=\frac{k_{B}}{\mu c^{2}}\bigg({T_{b}}-\frac{T_{b}^{\prime}}{3\mathscr{H}}\bigg),$
@@
 The differential equation for $T_{b}^{\prime}$ is very stiff and sensitive to $T_{b}-T_{\gamma}$, but $T_{b}\approx T_{\gamma}$ in the early universe, so we rewrite it to a more stable differential equation for $\Delta T^{\prime}=T_{b}^{\prime}-T_{\gamma}^{\prime}$ instead, initialize $\Delta T=0$ and observe $T_{b}=\Delta T+T_{\gamma}$.
-The optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\tau^{\prime})\mathrm{d}\tau^{\prime}$ is really a line-of-sine integral into the past, but is integrated together with the background ODEs by initializing $\kappa(\tau_{i})=0$ to an arbitrary value, integrating the differential equation for $\kappa^{\prime}$ and subtracting the final value of $\kappa(\tau_{0})$ (i.e.
+The optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\tau^{\prime})\mathrm{d}\tau^{\prime}$ is really a line-of-sight integral into the past, but is integrated together with the background ODEs by initializing $\kappa(\tau_{i})=0$ to an arbitrary value, integrating the differential equation for $\kappa^{\prime}$ and subtracting the final value of $\kappa(\tau_{0})$ (i.e.
 $\int_{\tau_{0}}^{\tau}=\int_{\tau_{0}}^{\tau_{i}}+\int_{\tau_{i}}^{\tau}=\int_{\tau_{i}}^{\tau}-\int_{\tau_{i}}^{\tau_{0}}$).
@@
 This is very expensive if done naively, and it is extremely important to choose a quadrature scheme that minimizes the number of sampled points.
-Fortunately, the momentum integrals have a structure that can be exploited: they are all in the form weighted form $I[g(x)]=\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)$, where $f(x)=1/(e^{x}+1)$ is the equilibrium distribution function and $g(x)$ is an arbitrary function of the dimensionless momentum $x=pc/k_{B}T$ (see Ma & Bertschinger (1995) for more details).
+Fortunately, the momentum integrals have a structure that can be exploited: they are all in the form weighted form $I[g(x)]=\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)$, where $f(x)=1/(e^{x}+1)$ is the equilibrium distribution function and $g(x)$ is an arbitrary function of the dimensionless momentum $x=pc/k_{B}T$ (see Ma and Bertschinger (1995) for more details).
 One can generally approximate $I[g(x)]\approx\sum_{i}W_{i}\,g(x_{i})$ with a weighted quadrature scheme with points $x_{i}$ and weights $W_{i}$ (more on this after the equations).
@@
 Note that this momentum quadrature strategy is generic with respect to the distribution function $f(x)$ and substitution $u(x)$, so it can easily be modified for other particle species whose distribution function cannot be integrated out.
-CAMB (Lewis, 2025) and CLASS (Lesgourgues & Tram, 2011) apply similar weighted quadrature strategies.
+CAMB (Lewis, 2025) and CLASS (Lesgourgues and Tram, 2011) apply similar weighted quadrature strategies.
 They also get away with only a handful of sampled momenta, but the precise details of the computation differ slightly.
@@
 $\Theta_{l}^{\mathrm{E}}\big(\tau_{0},k\big)=\sqrt{\frac{(l+2)!}{(l-2)!}}\int_{\tau_{i}}^{\tau_{0}}S_{E}\big(\tau,k\big)\frac{j_{l}\big(k(\tau_{0}-\tau)\big)}{\big(k(\tau_{0}-\tau)\big)^{2}}\mathrm{d}\tau\quad\text{with}\quad S_{E}=\frac{3}{16}v\Pi_{\gamma}.$
-As first suggested by Seljak & Zaldarriaga (1996), this approach enables cheap computation for any $l$ after integrating the perturbation ODEs with only a few $l\leq l_{\text{max}}$.
+As first suggested by Seljak and Zaldarriaga (1996), this approach enables cheap computation for any $l$ after integrating the perturbation ODEs with only a few $l\leq l_{\text{max}}$.
 This drastically speeds up the computation over including all $l$ in an enormous set of coupled perturbation ODEs.
@@
 Both grids range between $0.1l_{\text{min}}/\tau_{0}\leq k\leq 3l_{\text{max}}/\tau_{0}$, where $l_{\text{min}}$ and $l_{\text{max}}$ are the angular spectrum’s minimum and maximum requested multipoles.
+A.12 Adaptive source function refinement
+SymBoltz can optionally refine source functions $S(\tau,k)$ in $k$, such as when computing the matter or CMB power spectra.
+First, the source function $S_{i}(\tau)=S(\tau,k_{i})$ is found on an initial grid of wavenumbers $k_{i}$ by integrating their perturbations.
+The method then iterates over each interval $(k_{i},k_{i+1})$, integrates the perturbations for the center $k_{i+1/2}=(k_{i}+k_{i+1})/2$ and finds $S_{i+1/2}(\tau)=S(\tau,k_{i+1/2})$.
+The idea is then to compare this to a linearly interpolated source $S_{\text{int}}(\tau)=(S_{i}(\tau)+S_{i+1}(\tau))/2$ with some error measure err, for example from $\texttt{err}^{2}=\int\mathrm{d}\tau\big(S_{\text{int}}(\tau)-S_{i+1/2}(\tau)\big)^{2}$.
+If the interpolation is poor and err is greater than some absolute and relative tolerances, the left and right half-intervals $(k_{i},k_{i+1/2})$ and $(k_{i+1/2},k_{i+1})$ are recursively refined in the same way.
+This gives $S(\tau,k)$ by integrating as few kk–modes as possible.
+Instead of manually creating a non-uniform $k$-grid with many hyperparameters, the process is controlled by only one tolerance.
+It parallelizes over each interval and can also be done in $\log k$.
 Appendix B Precision parameters for CLASS
-When comparing results to CLASS in section 3, CLASS is configured with the following non-default precision parameters: {codebox}
-background_Nloga = 6000tight_coupling_trigger_tau_c_over_tau_h = 1e-2tight_coupling_trigger_tau_c_over_tau_k = 1e-3radiation_streaming_approximation = 3ur_fluid_approximation = 3ncdm_fluid_approximation = 3We also set l_max_g, l_max_pol_g, l_max_ur, l_max_ncdm to the same $l_{\text{max}}$ used by SymBoltz’ model.
-These settings disable as many approximations as possible and reduces the impact of the tight-coupling approximation, which cannot be disabled.
-Oddly, we find that the parameter background_Nloga must be decreased from the default value 40000 to make the derivatives in fig.
+When comparing results to CLASS in section 3, we use CLASS version 3.3.4 with the following precision parameters: {codebox}
+tight_coupling_approximation = 5 # compromise_CLASS; cannot disable TCAtight_coupling_trigger_tau_c_over_tau_h = 0.01 # turn off TCA earliertight_coupling_trigger_tau_c_over_tau_k = 0.001 # turn off TCA earlierradiation_streaming_approximation = 3 # turn off RSAncdm_fluid_approximation = 3 # turn off NCDMFAur_fluid_approximation = 3 # turn off UFAevolver = 1 # for implicit ndf15; 0 for explicit rktol_perturbations_integration = 1.0e-6 # less noise; varied in work-precision diagramtol_ncdm_newtonian = 1.0e-5 # default sampling of massive neutrino momentabackground_Nloga = 6000 # for stable derivativesWe run CLASS with output = mPk to compute only $P(k)$, and output = tCl, pCl to compute only $C_{l}^{\text{TT}}$, $C_{l}^{\text{EE}}$ and $C_{l}^{\text{TE}}$ with the code’s most specialized and optimized paths.
+CLASS and SymBoltz are configured to use the same Newtonian gauge, $l_{\text{max}}$ for all species, RECFAST recombination model, $\tanh$-like reionization parametrization, $w_{0}w_{a}$ dark energy equations and cosmological parameter values.
+The settings above disable as many approximations as possible and reduce the impact of the tight-coupling approximation, which cannot be disabled.
+We had to decrease the parameter background_Nloga from the default value 40000 to make the derivatives in fig.
 4 stable.
 This parameter controls the number of points used for splining background functions in the perturbations.
-The default value of this parameter was changed from 3000 to 40000 in 2023, but we suspect that the increased density in points makes the splines susceptible to oscillations from numerical noise.
-These settings are important for good agreement between the Fisher forecasts in section 3.4.
-We used CLASS version 3.3.1.
+Its default value was changed from 3000 to 40000 in 2023, but we suspect that the increased density in points makes the splines oscillate from numerical noise.
+Likewise, the tolerance for the perturbations ODEs is set one notch smaller than the default $10^{-5}$ to make the output stable.
+We use CLASS’ default sampling of massive neutrino momenta, as discussed in section 4.
+When running CLASS with all approximations on in tables 1 and 7, the first block of parameters above are not used.
+In the work-precision comparison in fig.
+7, we also vary tol_perturbations_integration and evolver as described in the figure.
 Appendix C Testing and comparison to CLASS

### v3

@@
-SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver
-[ Scale = MatchLowercase, Path = fonts/, Extension = .ttf , UprightFont = *-Regular, ] 11institutetext: Institute of Theoretical Astrophysics, University of Oslo, P.O.Box 1029 Blindern, N-0315 Oslo, Norway
-11email: herman.sletmoen@astro.uio.no
-SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver
-Herman Sletmoen
+SymBoltz.jl: A symbolic-numeric, approximation-free, and differentiable linear Einstein–Boltzmann solver
+11institutetext: Institute of Theoretical Astrophysics, University of Oslo, PO Box 1029 Blindern, 0315 Oslo, Norway
+SymBoltz.jl: A symbolic-numeric, approximation-free, and differentiable linear Einstein–Boltzmann solver
+Herman Sletmoen Email: herman.sletmoen@astro.uio.no
 (Received $X$ / Accepted $X$)
-SymBoltz is a new Julia package that solves the linear Einstein-Boltzmann equations.
-It features a symbolic-numeric interface for specifying equations, is free of approximation switching schemes and is compatible with automatic differentiation.
-Cosmological models are built from replaceable physical components in a way that scales well in model space, or alternatively written as one compact system of equations.
-The modeler should simply write down their equations, and SymBoltz solves them and eliminates friction in the modeling process.
-Symbolic knowledge enables powerful automation of tasks, such as separating computational stages like the background and perturbations, generating the analytical Jacobian matrix and its sparsity pattern, and interpolating arbitrary variables from the solution.
+SymBoltz is a new Julia package for solving the linear Einstein–Boltzmann equations in cosmology.
+It features a symbolic-numeric interface for specifying equations, is free of approximation switching schemes, and is compatible with automatic differentiation.
+Cosmological models are built from replaceable physical components in a way that scales well to extended models, or alternatively written as one compact system of equations.
+The modeler provides their equations, and SymBoltz solves them while reducing friction in the modeling process.
+Symbolic knowledge enables powerful automation of tasks, such as separating computational stages (e.g., background and perturbations), generating analytical and sparse Jacobian matrices, and interpolating arbitrary variables from the solution.
 Implicit solvers integrate the full stiff equations at all times without approximations, which greatly simplifies the code.
-Performance remains as good as in existing approximation-based codes due to high-order implicit methods that take long time steps, fast generated code, optimal handling of the Jacobian and efficient sparse matrix methods.
-Automatic differentiation gives exact derivatives of any output with respect to any input, which is important for gradient-based Markov chain Monte Carlo methods in large parameter spaces, training of emulators, Fisher forecasting and sensitivity analysis.
+Performance remains comparable to existing approximation-based codes due to efficient high-order implicit methods, fast generated code, optimal handling of the Jacobian, and sparse matrix methods.
+Automatic differentiation gives exact derivatives of any output with respect to any input, which is important for gradient-based Markov chain Monte Carlo (MCMC) sampling in large parameter spaces, Fisher forecasting, emulator training, and sensitivity analyses.
 The main features form a synergy that reinforces the design of the code.
-Results agree with established codes to $0.1\%$ with standard precision.
+Output spectra agree with established codes up to 0.1% with standard precision levels.
 More work is needed to implement additional features and for fast reverse-mode automatic differentiation of scalar loss functions.
-SymBoltz is available at https://github.com/hersle/SymBoltz.jl with single-command installation and extensive documentation, and welcomes all contributions.
+SymBoltz is publicly available with single-command installation and extensive documentation.
+We welcome all contributions to the code from the community.
 Key Words.:
-Cosmology: theory - Methods: numerical
+methods: numerical – cosmology: theory
 […]
-Cosmology is at a crossroads (Freedman, 2017).
-Despite the enormous success of the $\mathrm{\Lambda CDM}$ model in explaining many observations, the increasing precision of modern observations are revealing tensions between theory and observations.
-These suggest that the current standard model is not the ultimate truth.
-On the theoretical side, this drives a search for the true cosmological model through modifications to $\mathrm{\Lambda CDM}$ (Bull et al., 2016).
-Efficiently exploring the space of models benefits from numerical tools that are easy to modify.
-Minimizing friction in the modeling process encourages relaxing model-dependent approximations, creating user-friendly interfaces and structuring codes with modular components that are easy to replace.
-Some of the most important such tools in the cosmological modeling toolbox are linear Einstein-Boltzmann solvers ( ?Boltzmann codes?) like CAMB (Lewis et al., 2000) and CLASS (Lesgourgues, 2011a).
-On the observational frontier, next-generation surveys like the Square Kilometer Array (Dewdney et al., 2009), Vera C.
-Rubin Observatory (LSST Science Collaboration, 2009), Dark Energy Spectroscopic Instrument (DESI Collaboration, 2016), Simon’s Observatory (The Simons Observatory Collaboration, 2019) and Euclid (Euclid Collaboration, 2025) promise more precise data.
+1 Introduction
+Cosmology is at a crossroads (Freedman, 2017), as the precision of modern observations continues to reveal tensions with theory.
+This suggests that the $\Lambda$ cold dark matter ($\mathrm{\Lambda CDM}$) model is incomplete, despite successfully explaining many observations.
+On the theoretical side, this situation drives a search for more realistic models through modifications to $\mathrm{\Lambda CDM}$ (Bull et al., 2016).
+This exploration benefits from access to numerical tools that are easy to modify, encouraging us to relax model-dependent approximations, create user-friendly interfaces, and structure codes with modular components.
+Linear Einstein–Boltzmann solvers ( ?Boltzmann codes?), such as CAMB (Lewis et al., 2000) and CLASS (Lesgourgues, 2011a), are essential tools in the cosmological modeling toolbox.
+On the observational frontier, next-generation surveys such as the Square Kilometre Array (Dewdney et al., 2009), Vera C.
+Rubin Observatory (LSST Science Collaboration, 2009), Dark Energy Spectroscopic Instrument (DESI Collaboration, 2016), Simons Observatory (The Simons Observatory Collaboration, 2019), and Euclid (Euclid Collaboration, 2025) promise more precise data.
 Setting models apart with upcoming data involves both theoretical model parameters and experimental nuisance parameters that coexist in large $O(100)$-dimensional spaces (Piras et al., 2024).
-In high dimensions, modern Markov chain Monte Carlo methods like Hamiltonian Monte Carlo and the No-U-Turn Sampler (Hoffman and Gelman, 2011) outperform the traditional Metropolis-Hastings algorithm (Hastings, 1970).
-They explore parameter space more efficiently, but need both the likelihood and its gradient with respect to parameters.
-Emulators have become popular to get this from existing non-differentiable codes by training differentiable neural networks to reproduce their output (e.g Bolliet et al., 2024; Bonici et al., 2024).
-Differentiability is also used in forward-modeling field-level inference with simulations and initial conditions from Boltzmann solvers, for example (Seljak et al., 2017).
+In these high dimensions, modern Markov chain Monte Carlo (MCMC) methods, such as Hamiltonian Monte Carlo and the No-U-Turn Sampler (Hoffman and Gelman, 2014), outperform the traditional Metropolis–Hastings algorithm (Hastings, 1970).
+They use both the likelihood and its gradient to explore parameter space faster.
+Emulators have become a popular proxy to get gradients from nondifferentiable codes by training differentiable neural networks to reproduce their output (e.g., Bolliet et al., 2024; Bonici et al., 2024).
+Differentiability is also used in forward-modeling field-level inference with simulations and initial conditions from Boltzmann solvers (Seljak et al., 2017).
 Automatic differentiation can compute derivatives more accurately and faster than approximate and brute-force finite differences.
-A differentiable Boltzmann solver is a cornerstone in the modern cosmological modeling toolbox.
-SymBoltz is a new Boltzmann solver that aims to meet these needs, featuring a convenient symbolic-numeric interface, approximation-freeness and differentiability.
+This makes directly differentiable Boltzmann solvers a key addition to the modern cosmological modeling toolkit.
+SymBoltz.jl [^1] https://github.com/hersle/SymBoltz.jl is a new Boltzmann solver that aims to meet these needs, featuring a convenient symbolic-numeric interface, approximation-freeness, and differentiability.
 At its core, a Boltzmann code solves the Einstein equations for a gravitational theory coupled to some particle species described by Boltzmann equations up to first perturbative order around a homogeneous and isotropic universe.
-For example, it predicts cosmic microwave background (CMB), baryon acoustic oscillation (BAO) or supernova (SN) observations, or generates initial conditions for non-linear $N$-body simulations of large-scale structure.
+For example, it predicts cosmic microwave background (CMB), baryon acoustic oscillation (BAO), or supernova (SN) observations, and generates initial conditions for nonlinear $N$-body simulations of large-scale structure.
 This article is structured as follows.
-Section 1 motivates SymBoltz by reviewing the historical development of Boltzmann codes and methods for computing derivatives.
-Section 2 describes the structure and main features of SymBoltz and how it differs from other codes.
-Section 3 shows example usage.
-Section 4 compares performance to an existing code.
-Section 5 describes possibilities for future work.
+Section 1 motivates SymBoltz by reviewing the history of Boltzmann codes and gradient methods.
+Section 2 presents its structure and main features.
+Section 3 shows example use.
+Section 4 compares performance to CLASS.
+Section 5 outlines paths for future work.
 Section 6 concludes with the current state of the code.
-Appendices A, B and C list equations, implementation details, comparison settings and testing methods.
-1 History and motivation
-Peebles and Yu (1970) were first to numerically integrate a comprehensive set of linear Einstein-Boltzmann equations.
-Their work was refined over several years, and Ma and Bertschinger (1995) established the groundwork for modern Boltzmann solvers with their code COSMICS.
-Seljak and Zaldarriaga (1996) soon realized that one can integrate photon multipoles by parts to reduce many differential equations to one integral solution and released CMBFAST.
-Their line-of-sight integration method has become a standard technique that greatly speeds up the calculation, but requires truncating the multipoles in the differential equations.
+Appendices A, B and C list equations and technical details.
+Historically, Peebles and Yu (1970) were the first to numerically integrate a comprehensive set of linear Einstein–Boltzmann equations.
+Their work was refined over several years, and Ma and Bertschinger (1995) laid the groundwork for modern Boltzmann solvers with the COSMICS code.
+Shortly after, Seljak and Zaldarriaga (1996) integrated the photon multipoles by parts to reduce thousands of coupled differential equations to independent integral solutions.
+This line-of-sight integration method was first implemented in their CMBFAST code and has become a standard technique that greatly speeds up the calculation, but requires truncating the multipoles in the differential equations.
 This Fortran codebase has since evolved into CAMB [^2] https://camb.info/ written by Lewis et al.
@@
 Doran (2005a) ported CMBFAST to C++ with the fork CMBEASY, which structured the code in an object-oriented fashion and improved user-friendliness, but this project is abandoned today.
-Lesgourgues (2011a); Blas et al.
+Lesgourgues (2011a) and Blas et al.
 (2011) started the second major family of Boltzmann solvers with the birth of CLASS [^3] http://class-code.net in C.
-It improved performance, user-friendliness, code flexibility, ease of modification and control over precision parameters.
-It was the first cross-check of CAMB from an independent branch and boosted the scientific accuracy of the Planck mission.
-Since then a healthy arms race have fueled refinements to CAMB and CLASS.
+It improved performance, user-friendliness, code flexibility, ease of modification, and control over precision parameters.
+It was the first independent cross-check of CAMB and boosted the scientific accuracy of the Planck mission.
+Since then a healthy arms race has fueled refinements to CAMB and CLASS.
 Both codes have spawned many forks for studying alternative models and performing custom calculations.
-Today they are very well-made, efficient and reliable tools.
+Today they are very well-made, efficient, and reliable tools.
 Recently, several alternative solvers have appeared on the market.
-They target new numerical techniques, lack of approximations, differentiability, GPU parallelization, interactivity and symbolic computation.
+They target new numerical techniques, lack of approximations, differentiability, GPU parallelization, interactivity, and symbolic computation.
 PyCosmo [^4] https://pypi.org/project/PyCosmo/ by Refregier et al.
-(2018) in Python was the first symbolic-numeric Boltzmann code that generates fast C++ code from user-provided symbolic equations, optimizes the sparse and analytical Jacobian matrix and avoids approximation schemes.
+(2018) in Python was the first symbolic-numeric Boltzmann code that generates fast C++ code from user-provided symbolic equations, optimizes the sparse and analytical Jacobian matrix, and avoids approximation schemes.
 Bolt.jl [^5] https://github.com/xzackli/Bolt.jl by Li et al.
-(2023) in Julia is also approximation-free, supports forward-mode automatic differentiation and uses similar differential equations solvers as SymBoltz.
+(2023) in Julia is also approximation-free, supports forward-mode automatic differentiation, and uses similar differential equations solvers as SymBoltz.
 DISCO-EB [^6] https://github.com/ohahn/DISCO-EB by Hahn et al.
 (2024) in the JAX framework in Python is also differentiable and relaxes approximation schemes to avoid overhead from switching equations on GPUs.
-SymBoltz.jl [^6] https://github.com/hersle/SymBoltz.jl is inspired by all these developments.
-1.1 Structure of traditional Boltzmann solvers
-The full Einstein-Boltzmann equations are partial differential equations that linearize to ordinary differential equations.
-In principle, the core task of a Boltzmann solver is thus very simple: to solve these ordinary differential equations (ODEs) ${\mathrm{d}\boldsymbol{u}}/{\mathrm{d}\tau}=\boldsymbol{f}(\tau,\boldsymbol{u},\boldsymbol{p})$ for some initial conditions $\boldsymbol{u}(\tau_{0})$ and parameters $\boldsymbol{p}$, including several perturbation wavenumbers, and compute desired quantities from their solution, like power spectra.
-In practice, however, several properties complicate this task.
-First, the equations separate into computational stages that benefit substantially from being solved sequentially for performance and stability, such as the background and perturbations stages.
+SymBoltz.jl [^7] https://github.com/hersle/SymBoltz.jl is another Boltzmann code that brings together several of these developments.
+The core task of an Einstein–Boltzmann solver is to solve the Einstein–Boltzmann equations for some cosmological model.
+They are partial differential equations that linearize to ordinary differential equations (ODEs) ${\mathrm{d}\boldsymbol{u}}/{\mathrm{d}\tau}=\boldsymbol{f}(\boldsymbol{u},\boldsymbol{p},\tau)$ for some initial conditions $\boldsymbol{u}(\tau_{i})$ and parameters $\boldsymbol{p}$, including several perturbation wavenumbers.
+Any output from Einstein–Boltzmann codes is derived from the solution of these ODEs, such as the matter and CMB power spectra.
+However, several properties of the equations complicate this task.
+First, the equations separate into computational stages that benefit substantially from being solved sequentially for performance and stability, such as the background, perturbations, and line-of-sight integration stages.
 It is common to solve each stage with interpolated input from the previous stage, and joining these can be cumbersome and fragment code.
@@
 Ideally, parts of the equations should be easily replaceable to accommodate different submodels for gravity and particle species.
-Organizing a code with a suitable modular structure that scales well in model space is hard.
-Third, wildly different timescales coexist in the equations and make them extremely stiff and intractable to solve with standard explicit ODE solvers.
-This stiffness must be massaged away in the equations or dealt with numerically in other ways.
-Fourth, the size of differential equations is very large and increases for more accurate description of relativistic species.
+It is hard to organize a code with a suitable modular structure that scales well in model space.
+Third, wildly different timescales coexist in the equations, making them extremely stiff and intractable to solve with standard explicit ODE solvers.
+This stiffness must be massaged away in the equations by approximations or dealt with numerically by implicit ODE solvers.
+Fourth, the number of differential equations is very large, particularly for accurate descriptions of relativistic species.
 Typical models with accurate treatment of photons and neutrinos need $O(100)$ equations.
-Fifth, the perturbations must be solved for many different wavenumbers $k$, and tradeoffs between performance and precision must be made.
-To overcome these challenges, most Boltzmann solvers are written in low-level high-performance languages like C, C++ and Fortran.
-They are tightly adapted to the pipeline-like computational structure of the problem (e.g.
-input $\rightarrow$ background $\rightarrow$ thermodynamics $\rightarrow$ perturbations $\rightarrow\ldots\rightarrow$ output in Lesgourgues (2011a)).
+Fifth, the perturbations must be solved for many different wavenumbers $k$, and trade-offs between performance and precision must be made.
+To overcome these challenges, most Boltzmann solvers are written in low-level high-performance languages such as C, C++, and Fortran.
+They are tightly adapted to the pipeline-like computational structure of the problem (e.g., input $\rightarrow$ background $\rightarrow$ thermodynamics $\rightarrow$ perturbations $\rightarrow\ldots\rightarrow$ output in Lesgourgues 2011a).
 This makes sense for programmers, but does not necessarily provide the simplest interface for modelers.
-Traditional codes like CAMB and CLASS have nonexistent or thin abstraction layers.
+Traditional codes such as CAMB and CLASS have nonexistent or thin abstraction layers.
 To modify them, users must work directly in the low-level numerical code and have a good understanding of its internal structure.
-For example, to implement a new species, one must often modify the code in many places: input handling for new parameters, new background equations, new thermodynamics equations, new perturbation equations, joining each of these stages, output handling, and so on.
+For example, to implement a new species, users must often modify the code in many places: input handling for new parameters, new background equations, new thermodynamics equations, new perturbation equations, joining each of these stages, output handling, and so on.
 This leads to fragmented code where changes related to one species are scattered throughout the code.
 This structure scales poorly in model space.
-As more species or gravitational theories are added, each module becomes intertwined with code from other physical components.
-Even if deactivated by if statements at runtime, their mere presence in the source code increases its complexity and reduces its readability.
-One can alleviate this problem to some extent by instead forking the code for modified models, so the main code base is not polluted.
-But this just moves the problem.
-Forking duplicates the entire code base even though only small parts are modified.
+As more species and gravitational theories are added, each module is intertwined with code from other physical components.
+Even if unused components are deactivated by ?if?
+statements at runtime, their mere presence in the source code increases its complexity, reduces its readability, and makes it harder to add more models.
+We can alleviate this problem to some extent by instead forking the code for modified models, so the main code base is not polluted.
+But this only shifts the problem.
+Forking duplicates the entire code base, even though only small parts are modified.
 Forks are often abandoned and do not receive upstream improvements.
-They are also incompatible with each other unless one merges them into one code, but then one is back to the first problem.
-The two-language problem amplifies the problem, as data analysis often happens in slow high-level languages like Python.
+They are also incompatible with each other unless we merge them into one code, which reintroduces the first problem.
+The two-language problem amplifies this issue, as data analysis often happens in slower high-level languages such as Python.
 This shapes Boltzmann solvers to rigid pipelines that must compute everything at once and avoid interception at all costs, in order to maximize performance in the low-level language before passing output back to the high-level language.
-Some features are really just post-processing of the ODE solutions, but are appended to the pipeline even if peripheral to the core task of an Einstein-Boltzmann solver – solving the Einstein-Boltzmann equations.
-Features like non-linear boosting and CMB lensing could be made as smaller and interoperating modular packages, but instead become part of ?all-in-one?
-Boltzmann codes.
-These patterns lead to big monolithic Boltzmann solvers.
-They become increasingly complex as they incorporate more models and features beyond their original core scope.
+Some features really just post-process the ODE solutions, but are appended to the pipeline even if they are peripheral to the core task of an Einstein–Boltzmann solver.
+For example, features such as nonlinear boosting and CMB lensing could be outsourced to smaller and interoperating modular packages, but they often become part of ?all-in-one?
+Boltzmann codes instead.
+These patterns lead to big monolithic Boltzmann solvers that become increasingly complex as they incorporate more models and features beyond their original core scope.
 This complexity has even driven development of specialized AI assistants for CLASS (Casas et al., 2025).
 While existing Boltzmann solvers are impressively well-engineered and such tools are only helpful, we think these are symptoms of unnecessary complexity.
-1.2 Boltzmann solver approximation schemes
-The Einstein-Boltzmann equations are infamously stiff (Nadkarni-Ghosh and Refregier, 2017).
+The Einstein–Boltzmann equations are notoriously stiff (Nadkarni-Ghosh and Refregier, 2017).
 This property of differential equations means their numerical solution is unstable with standard explicit integrators and requires tiny step sizes, making them hard to solve.
 Stiffness can arise when multiple and very different time scales appear in the same problem.
-This is very common in cosmology, where particles interact very rapidly in a universe that expands very slowly, particularly in the tightly coupled baryon-photon fluid, for example.
+This is very common in cosmology, where particles interact very rapidly in a universe that expands very slowly, particularly in the tightly coupled baryon-photon fluid.
 Stiff equations are practically impossible to integrate with explicit solvers and require special treatment.
-For a long time, Boltzmann solvers have massaged away stiffness with several approximation schemes [^7] Here ?approximations?
+For a long time, Boltzmann solvers have massaged away stiffness with several approximation schemes [^8] Here, ?approximations?
 refers to schemes that switch between different equations at different times.
-It excludes techniques like multipole truncation and line-of-sight integration, which SymBoltz also uses.
-in the equations (e.g.
-Doran (2005b); Blas et al.
-(2011); Lewis (2025)):
-  • tight-coupling approximation (TCA),
-  • ultra-relativistic fluid approximation (UFA),
-  • radiation streaming approximation (RSA),
-  • non-cold dark matter fluid approximation (NCDMFA),
+It excludes techniques such as multipole truncation and line-of-sight integration, which SymBoltz also uses.
+in the equations (e.g., Doran, 2005b; Blas et al., 2011):
+  • tight-coupling approximation (TCA);
+  • ultrarelativistic fluid approximation (UFA);
+  • radiation streaming approximation (RSA);
+  • noncold dark matter fluid approximation (NCDMFA);
   • Saha approximation.
 They involve switching from one set of equations to another when some control variable measuring the applicability of the approximation crosses a threshold.
-This can change the variables in the ODE and require reinitializing it.
-Approximations can help both speed and stability and allow the use of explicit solvers.
-However, approximations put much more strain on modelers to derive and validate several versions of the equations in different regimes.
-This process may have to be repeated when the model is modified.
-The numerics also become more complicated: timeseries from each ODE solution must be stitched together, each separate ODE system can use different tolerances, ODE integrators must be reinitialized, and so on.
+This can change the variables in the ODE and require reinitializing the integration.
+Approximations enable explicit solvers and can improve both the speed and stability of the solution.
+However, they put much more load on modelers to derive and validate several versions of the equations in different regimes.
+This process must generally be repeated with modifications to the model, as they can invalidate the approximations or reintroduce stiffness.
+They also complicate the numerics, as time series from each ODE solution must be stitched together, each separate ODE system can use different tolerances, ODE integrators must be reinitialized, and so on.
 Another way to integrate stiff equations is to use appropriate implicit solvers.
-PyCosmo first solved the full stiff Einstein-Boltzmann equations with an implicit integrator, followed by Bolt.jl and DISCO-EB.
+PyCosmo first solved the full stiff Einstein–Boltzmann equations with an implicit integrator, followed by Bolt and DISCO-EB.
 CLASS also has an implicit solver, but does not permit disabling the tight-coupling approximation, for example.
 Historically, implicit solvers may have been underutilized because they are harder to implement than explicit solvers and have a reputation for being slower, and due to iterative evolution from initial Boltzmann codes that were centered around approximations.
-However, new life has recently been breathed into the field of implicit solvers, with development of new methods combined with techniques like automatic and symbolic differentiation that make them more feasible and powerful (e.g.
-Steinebach, 2023; Ekanathan et al., 2025).
-1.3 Differentiation methods
-Derivatives are important in computing.
-Cosmological applications are no exception.
-For example, algorithms that optimize likelihoods and Markov chain Monte Carlo (MCMC) samplers for Bayesian parameter inference can take advantage of derivatives of the likelihood with respect to each parameter to intelligently step in a direction where the likelihood increases.
-In machine learning, the same applies when training neural networks emulators for cosmological observables as functions of parameters by minimizing a scalar loss.
-Some cosmologies are parametrized as boundary-value problems with the shooting method, and use nonlinear root solvers like Newton’s method that need Jacobians.
+However, new life has recently been breathed into the field of implicit methods, with the development of new solvers combined with techniques such as automatic and symbolic differentiation that make them more feasible and powerful (e.g., Steinebach, 2023; Ekanathan et al., 2025).
+Derivatives are important in scientific computing and in cosmological applications.
+For example, algorithms that optimize likelihoods and MCMC samplers for Bayesian parameter inference can take advantage of derivatives of the likelihood with respect to each parameter to intelligently step in a direction where the likelihood increases.
+In machine learning, the same applies when training neural network emulators for cosmological observables by minimizing a scalar loss function of parameters.
+Some cosmologies are parameterized as boundary-value problems with the shooting method and use nonlinear root solvers such as Newton’s method, which needs Jacobians.
 Implicit ODE solvers also use Jacobians to solve for values at the next time step.
 Fisher forecasting uses the Hessian (double derivative) of the likelihood with respect to parameters to predict how strong parameter constraints that can be placed by data with some uncertainty.
-Boltzmann solvers save time by interpolating spectra computed on coarse grids of $k$ and $l$ to finer grids, and this can be made more precise with derivatives with respect to $k$ and $l$.
+Boltzmann solvers save time by interpolating spectra computed on coarse grids of $k$ and $l$ to finer grids, which can be made more precise with derivatives with respect to $k$ and $l$.
 These are all examples where (applications of) Boltzmann solvers need derivatives.
 There are at least four ways to compute derivatives.
-Manual differentiation is human application of differentiation rules, but is limited to simple expressions and by human error.
-Symbolic differentiation automates this process with computer algebra systems, but is inherently symbolic and cannot differentiate arbitrary programs nested with control flow like conditional statements and loops that depend on numerical values.
+Manual differentiation is human application of differentiation rules, but it is limited to simple expressions and by human error.
+Symbolic differentiation automates this process with computer algebra systems, but it is inherently symbolic and cannot differentiate arbitrary programs with control flow, such as conditional statements and loops that depend on numerical values.
 Finite differentiation approximates the derivative $f^{\prime}(x)\approx(f(x+\epsilon/2)-f(x-\epsilon/2))\,/\,\epsilon$ with a small $\epsilon>0$ (here using central differences) by simply evaluating the program several times.
-This can differentiate arbitrary programs, but is approximate, introduces the step size $\epsilon$ as a hyperparameter that must be tuned for both accuracy and stability, and is a brute-force approach that requires $O(n)$ evaluations ($2n$ using central differences) to compute the gradient of an $n$-variate scalar $f$.
-Automatic differentiation can be understood by viewing any computer program as a (big) composite function
+This can differentiate arbitrary programs, but it is approximate and introduces the step size $\epsilon$ as a hyperparameter that must be tuned for accuracy and stability.
+It is a brute-force approach that requires $O(n)$ evaluations ($2n$ using central differences) to compute the gradient of a function $f$ of $n$ variables.
+Automatic differentiation (e.g., Griewank and Walther, 2008) can be understood by viewing any computer program as a (big) composite function
 $\boldsymbol{f}=\boldsymbol{f}_{N}\circ\boldsymbol{f}_{N-1}\circ\cdots\circ\boldsymbol{f}_{2}\circ\boldsymbol{f}_{1}=\boldsymbol{f}_{N}(\boldsymbol{f}_{N-1}(\cdots\boldsymbol{f}_{2}(\boldsymbol{f}_{1})))$ (1)
-of (many) elementary operations $\boldsymbol{f}_{i}:\mathbb{R}^{a_{i}}\rightarrow\mathbb{R}^{b_{i}}$ (think of $\boldsymbol{f}_{i}$ as the $i$-th line of code).
-It then numerically evaluates the chain rule
-$\boldsymbol{J}=\boldsymbol{J}_{N}\cdot\boldsymbol{J}_{N-1}\cdots\boldsymbol{J}_{2}\cdot\boldsymbol{J}_{1}.$ (2)
-through the Jacobian $(\boldsymbol{J}_{n})_{ij}=\partial f_{n,i}/\partial f_{n-1,j}$ of every elementary operation to accumulate the derivative of the entire program.
-This is numerically exact, free of precision parameters, usually requires fewer operations than finite differences, but is perhaps less intuitive, harder to implement and needs the source code of the program (1) to interpret it in the non-standard way (2).
+of elementary operations $\boldsymbol{f}_{i}:\mathbb{R}^{m_{i}}\rightarrow\mathbb{R}^{n_{i}}$ (think of $\boldsymbol{f}_{i}$ as the $i$-th line of code). Then it numerically evaluates the chain rule
+$\boldsymbol{J}=\boldsymbol{J}_{N}\cdot\boldsymbol{J}_{N-1}\cdot\cdots\cdot\boldsymbol{J}_{2}\cdot\boldsymbol{J}_{1}$ (2)
+through the Jacobian $(\boldsymbol{J}_{n})_{ij}=\partial f_{n,i}/\partial f_{n-1,j}$ of every operation to accumulate the derivative of the entire program.
+This is numerically exact and free of precision parameters, while usually requiring fewer operations than finite differences.
+But it is perhaps less intuitive, harder to implement, and needs the source code of the program (1) to interpret it in the nonstandard way (2).
 Notably, while the function (1) must be evaluated inside-to-outside (right-to-left), the chain rule (2) is an associative matrix product that can be evaluated in any order.
 This generally changes the number of operations and is a more open-ended computational problem.
-Forward-mode automatic differentiation seeds $\boldsymbol{J}_{1}=\boldsymbol{1}$ (the derivative of the input with respect to itself) and multiplies $\boldsymbol{J}_{N}(\boldsymbol{J}_{N-1}(\cdots(\boldsymbol{J}_{2}(\boldsymbol{J}_{1}))))$ by ?pushing?
+Forward-mode automatic differentiation seeds $\boldsymbol{J}_{1}=\boldsymbol{1}$ (the derivative of the input with respect to itself) and multiplies $\boldsymbol{J}_{N}(\boldsymbol{J}_{N-1}(\cdots(\boldsymbol{J}_{2}\boldsymbol{J}_{1})))$ by ?pushing?
 every column of $\boldsymbol{J}_{1}$ forward through the product in the same evaluation order as $\boldsymbol{f}$.
-Reverse-mode first computes $\boldsymbol{f}$ in a forward pass, then seeds $\boldsymbol{J}_{N}=\boldsymbol{1}$ (the derivative of the output with respect to itself) and multiplies $((((\boldsymbol{J}_{N})\boldsymbol{J}_{N-1})\cdots)\boldsymbol{J}_{2})\boldsymbol{J}_{1}$ by ?pulling?
+Reverse-mode first computes $\boldsymbol{f}$ in a forward pass, then seeds $\boldsymbol{J}_{N}=\boldsymbol{1}$ (the derivative of the output with respect to itself) and multiplies $(((\boldsymbol{J}_{N}\boldsymbol{J}_{N-1})\cdots)\boldsymbol{J}_{2})\boldsymbol{J}_{1}$ by ?pulling?
 every row of $\boldsymbol{J}_{N}$ backwards through the product.
-This usually makes forward-mode faster when $\boldsymbol{f}:\mathbb{R}^{a}\rightarrow\mathbb{R}^{b}$ has more outputs $b\gg a$, and reverse-mode better when there are more inputs $a\gg b$.
-In practice, both modes are implemented using techniques like operator overloading or source code transformation.
+This usually makes forward-mode faster when $\boldsymbol{f}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ has more outputs ($n\gg m$), and reverse-mode better when there are more inputs ($m\gg n$).
+In practice, both modes are implemented using techniques called operator overloading or source code transformation.
 The former specializes every function on a particular ?dual number?
-that computes both the value and derivative of the function (e.g.
-Revels et al., 2016).
+type that propagates both the value and derivative of the function (e.g., Revels et al., 2016).
 The latter analyzes the source code for $\boldsymbol{f}$ and transforms it into another code that computes $\boldsymbol{J}$.
-In any case, automatic differentiation requires access to the source code!
+In any case, automatic differentiation does not work on compiled binaries, but requires access to the source code to compute gradients with a different path through the instructions of the program.
 GravityBaryonsPhotonsMasslessneutrinosMassiveneutrinosCold darkmatterCosmo-logicalconstant Symbolic model Background: $\boldsymbol{f}$, $\boldsymbol{J}$Perturbations: $\boldsymbol{f}$, $\boldsymbol{J}$ Numerical problem Background: $S(\tau)$Perturbations: $S(\tau,k)$ Solution object CompileSolve
-Figure 1: SymBoltz represents cosmological models with symbolic equations grouped in physical components for the metric, gravity and particle species.
+Figure 1: SymBoltz represents cosmological models with symbolic equations grouped in physical components for the metric, gravity, and particle species.
 This is compiled to a numerical problem that splits equations into background and perturbation stages and generates fast code for ODE functions $\boldsymbol{f}$ and Jacobians $\boldsymbol{J}$.
-The problem is then solved and the result stored in a solution object that gives access to any variable $S$ in the model.
+The problem is then solved, and the result is stored in a solution object that gives access to any model variable $S$.
 2 Code architecture and main features
 SymBoltz is designed around three main features.
-In short, it has a symbolic-numeric abstraction interface where users enter high-level symbolic equations, and automatically compiles them to fast low-level numerical functions.
-It cures stiffness with implicit ODE solvers, and uses no approximation schemes to keep models simple, elegant and extensible.
-It is differentiable, so one can get accurate derivatives of any output with respect to any input.
-This provides rapid model prototyping, helpful abstractions and automates tasks one must do manually when modifying other codes.
+In brief, it has a symbolic-numeric abstraction interface where users enter high-level symbolic equations that are automatically compiled to fast low-level numerical functions.
+It cures stiffness with implicit ODE solvers instead of approximation schemes to keep models simple, elegant, and extensible.
+It is differentiable, so it is possible to get accurate derivatives of any output with respect to any input.
+This provides rapid model prototyping, helpful abstractions, and automates tasks that must be done manually when modifying other codes.
 SymBoltz encourages interactive use and pursues a modular design that lets users integrate what they need from the package into their own applications.
-The end goal is to prioritize the modeler, who should be able to just write down their equations while SymBoltz handles modeling chores.
-The code is written in the Julia programming language (Bezanson et al., 2017), with a rich ecosystem of scientific packages and an aim to resolve the two-language problem.
-The symbolic-numeric interface is built on ModelingToolkit.jl (Ma et al., 2022) and Symbolics.jl (Gowda et al., 2022).
-The compiled functions are integrated by implicit ODE solvers in OrdinaryDiffEq.jl (Rackauckas and Nie, 2017).
-Automatic differentiation works through ForwardDiff.jl (Revels et al., 2016).
-The next subsections describe SymBoltz’ three main features in more detail.
+The end goal is to prioritize the modeler, who should be able to just write down their equations, while SymBoltz automates modeling chores.
+The code is written in the Julia programming language (Bezanson et al., 2017), which has a rich ecosystem of scientific packages and aims to resolve the two-language problem.
+The symbolic-numeric interface is built on ModelingToolkit.jl [^9] https://github.com/SciML/ModelingToolkit.jl by Ma et al.
+(2022) and Symbolics.jl [^10] https://github.com/JuliaSymbolics/Symbolics.jl by Gowda et al.
+(2022).
+The compiled functions are solved by implicit ODE integrators in OrdinaryDiffEq.jl [^11] https://github.com/SciML/OrdinaryDiffEq.jl/ by Rackauckas and Nie (2017) and linear algebra methods in LinearSolve.jl [^12] https://github.com/SciML/LinearSolve.jl.
+Automatic differentiation works through ForwardDiff.jl [^13] https://github.com/JuliaDiff/ForwardDiff.jl by Revels et al.
+(2016).
+The next subsections describe SymBoltz’ three main features in depth.
 Other implementation details are given in Appendix A.
-This paper is as of SymBoltz version 0.14.1.
-Please see the package documentation for definitive up-to-date information.
+This paper describes SymBoltz version 1.0.0.
+We refer to the package documentation for definitive up-to-date information.
 2.1 Symbolic-numeric interface
@@
 1.
-Variables and equations are specified in a high-level user-friendly symbolic modeling language, then compiled down to low-level numerical code that is integrated by ODE solvers.
-With knowledge of the symbolic equations, SymBoltz can analyze their structure programmatically to ease model specification, automate related boilerplate tasks and generate fast and stable code that avoids approximations.
-We think three properties motivate some kind of symbolic abstraction layer around the Einstein-Boltzmann equations.
-First, the equations are most easily written as one large ODE, but optimally solved as several ODEs in stages like the background, perturbations and line-of-sight integration.
-Mathematically, the Jacobian matrix of this single ODE encodes which variables are independent and belong to which stages.
-It is easiest for modelers to specify equations in one common system, where the perturbations can refer to variables in the background, for example.
-The separation into computational stages can be automated.
-Second, solving each stage is hard due to stiffness, performance requirements and its dependence on previous stages.
-To overcome this, it is helpful to automatically generate ODE code that is optimal in speed, accuracy and stability, generate the ODE Jacobian needed by implicit solvers and interpolate variables from previous stages.
-Modelers should not worry about this.
-Third, the equations are often modified because the true cosmology is unknown.
-This can be made easier by organizing equations in a way that reflects the physical structure of the model, so it is easy to replace a related subset of equations with another without duplicating the entire model.
+A model defines what equations should be solved; a problem decides how they are solved; and a solution holds what was solved.
+Variables and equations are specified in a high-level user-friendly symbolic modeling language, then compiled to low-level numerical code that is integrated by ODE solvers.
+With knowledge of the symbolic equations, SymBoltz analyzes their structure programmatically to ease model specification, automates mechanical boilerplate tasks, and generates fast and stable code that avoids approximations.
+In our opinion, there are three properties that motivate such a symbolic abstraction layer around the Einstein–Boltzmann equations.
+First, the equations are most easily written as one large ODE, but optimally solved as several ODEs in stages such as the background, perturbations, and line-of-sight integration.
+Mathematically, the Jacobian matrix of the single ODE encodes which variables are independent and belong to which stages.
+It is easiest for modelers to specify equations in one common system, where the perturbations can refer to variables in the background, and then separate the computational stages automatically.
+Second, solving each stage is hard due to stiffness, performance requirements, and dependence on previous stages.
+To overcome this, it helps to automatically generate ODE code that is optimal in speed, accuracy, and stability, generate the ODE Jacobian needed by implicit solvers, and interpolate variables from previous stages.
+Modelers should not have to do this manually.
+Third, the equations are often modified, as we do not know the true cosmology.
+This can be made easier by organizing equations in a way that reflects the physical structure of the model, so it is easy to replace one related subset of equations with another without duplicating the entire model.
 This could be slow to do at the numerical level, but can be done with no runtime penalty in an intermediate symbolic representation of the equations.
-Inspired by this, SymBoltz inverts the traditional layout of Boltzmann solvers.
-While most codes are built as pipelines that follow the background, perturbations and other stages of the equations, SymBoltz is primarily structured around the physical components that they consists of.
-Related variables and equations are grouped in distinct components for the metric, gravitational theory, photons, baryons, dark matter, dark energy, neutrinos and other species (see fig.
+Inspired by this, SymBoltz inverts the traditional layout of Boltzmann codes.
+Most codes are built as pipelines following the background, perturbations, and other stages, but SymBoltz is primarily structured around the physical components that make up the equations.
+Related variables and equations are grouped in distinct components for the metric, gravity, photons, baryons, dark matter, dark energy, neutrinos, and other species (see Fig.
 1).
-This isolates everything related to one submodel to one place.
-Any set of such submodels are joined into a full cosmological model, like $\mathrm{\Lambda CDM}$.
-Interactions (e.g.
-Compton scattering or sourcing of gravity) are equations between components.
+This isolates everything related to one submodel in one place.
+Any set of such submodels is joined into a full cosmological model, such as $\mathrm{\Lambda CDM}$.
+Interactions (e.g., Compton scattering or sourcing of gravity) are equations that connect components.
 Adding new submodels is easy, and SymBoltz is largely devoted to simply building a well-organized library of them.
-SymBoltz then takes a full cosmological model, separates it into stages and generates numerical code to solve each stage.
-This preprocessing does not slow the code at runtime; to the contrary the symbolic equations give extra information like the Jacobian that enhances speed and stability.
+SymBoltz then takes a full cosmological model, separates it into stages, and generates numerical code to solve each stage.
+This preprocessing does not slow the code at runtime.
+To the contrary, the symbolic equations give extra information such as the Jacobian that enhances speed and stability.
 Modifications automatically enjoy these benefits without extra user input.
 The modular component-based structure scales well in model space.
-For example, modified gravity or dark energy models can be written down as self-contained submodels and combined with other components to create many different extended cosmological models.
-The generic symbolic framework can also be used to build reduced models with non-interacting radiation, matter and dark energy species.
+For example, modified gravity or dark energy models can be written as self-contained submodels and combined with other components to create many different extended cosmological models.
+The generic symbolic framework also handles reduced models with noninteracting radiation, matter and dark energy species.
 When exploring modified gravity theories, it can be helpful to test such toy models to understand how gravity responds to pure fluids without coupling to baryons and photons.
-In contrast, the monolithic layout of traditional codes scatter changes related to one submodel across different modules for input/output, the background, perturbations and so on.
-As more and more submodels are added, the code is increasingly intertwined and fragmented.
-They are also not flexible enough to test simpler toy models, as baryons and photons are hardcoded as fundamental to the structure of the code.
-This design fits the computational structure better than the physical one.
-It can result in an overwhelming code that becomes hard to read and modify, and whose complexity scales poorly in model space.
+In contrast, the monolithic layout of traditional codes scatters changes for one submodel across different modules for input, the background, perturbations, output, and so on.
+As more submodels are added, the code is increasingly intertwined and fragmented.
+They are also not flexible enough to test simpler toy models, as baryons and photons are inherently hardcoded.
+This design fits the computational procedure better than the physical structure.
+It can lead to an overwhelming code that is hard to read and modify.
+This complexity scales poorly in model space.
 As an alternative to component-based modeling, SymBoltz also includes a version of the default $\mathrm{\Lambda CDM}$ model with all equations packed into one big ?unstructured?
@@
 This makes it trivial to change anything in the equations, but sacrifices modularity.
-The symbolic interface makes it very compact: SymBoltz defines a full $\mathrm{\Lambda CDM}$ model in just 277 straightforward lines of code [^8] https://hersle.github.io/SymBoltz.jl/stable/unstructured/, while the equivalent code to browse in CLASS is spread over 10 files with 27721 lines [^9] Counting input.{h,c}, background.{h,c}, thermodynamics.{h,c}, perturbations.{h,c} and wrap_recfast.{h,c} with wc -l..
-PyCosmo’s symbolic interface also generates code for ODEs and sparse Jacobians, but does not focus on component-based modeling or automating tasks like stage separation, for example.
-SymBoltz’ symbolic-numeric interface simplifies solving and modifying the Einstein-Boltzmann equations.
-It aims to maximize speed, stability and convenience with minimal user input.
-The next sections go into detail on how this is done.
-Section 3.2 shows a concrete example and some of its features in action.
+The symbolic interface makes it very compact: SymBoltz defines a full $\mathrm{\Lambda CDM}$ model in only 277 lines of code, [^14] https://hersle.github.io/SymBoltz.jl/stable/LCDM/ while the equivalent code to browse in CLASS is spread over 10 files with 27721 lines.151515input.{h,c}, background.{h,c}, thermodynamics.{h,c}, perturbations.{h,c}, and wrap_recfast.{h,c} counted by wc.
+PyCosmo’s symbolic interface also generates code for ODEs and sparse Jacobians, but does not focus on component-based modeling or automating tasks such as stage separation.
+In this way, SymBoltz’ interface simplifies the solution and modification of the Einstein–Boltzmann equations.
+It aims to maximize speed, stability, and convenience with minimal user input.
+The next subsections explain how the symbolic features work in detail, and Sect.
+3 shows them in action in a concrete example.
 2.1.1 Automatic numerical code generation
-SymBoltz automatically compiles symbolic equations to numerical code for ordinary differential equations (ODEs)
-$\frac{\mathrm{d}\boldsymbol{u}}{\mathrm{d}\tau}=\boldsymbol{f}(\tau,\boldsymbol{u},\boldsymbol{p}).$ (3)
+SymBoltz automatically compiles symbolic equations to numerical code for ODEs
+$\frac{\mathrm{d}\boldsymbol{u}}{\mathrm{d}\tau}=\boldsymbol{f}(\boldsymbol{u},\boldsymbol{p},\tau).$ (3)
 The generated code is fast and prevents users unfamiliar with Julia or SymBoltz from writing slow code.
-If necessary, one can escape the standard code generation and call arbitrary numerical functions, for example to solve a nonlinear equation for the minimum of a potential or interpolate tabulated data.
-The code generation handles tasks like allocating indices for each ODE state $u_{i}$, and performs optimizations such as common subexpression elimination (see section 3.2).
-This is helpful as Boltzmann solvers tend to have big $\boldsymbol{f}$ and ODE solvers call $\boldsymbol{f}$ many times.
+If necessary, we can escape the standard code generation and call arbitrary numerical functions; for example, to solve a nonlinear equation for the minimum of a potential or interpolate tabulated data.
+The code generation handles tasks such as allocating indices for each ODE state $u_{i}$, and does optimizations such as common subexpression elimination (see Sect.
+3.2).
+This is helpful as Boltzmann solvers tend to have large $\boldsymbol{f}$, and ODE solvers evaluate $\boldsymbol{f}$ many times.
 2.1.2 Automatic handling of observed variables
@@
 variables are any functions of the unknowns.
-The Einstein-Boltzmann equations are usually written with many observed variables.
-For example, consider the metric and gravity equations in sections A.1 and A.2 sourced by some arbitrary $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$.
-Here $a(\tau)$ and $\Phi(\tau,k)$ are the only unknown (differential) variables that the ODE is solved for, while one can ?observe?
-$z(\tau)$, $\mathscr{H}(\tau)$, $H(\tau)$ and $\Psi(\tau,k)$ from the unknowns.
-In addition, observed derivatives like $\mathscr{H}^{\prime}$ or $\Psi^{\prime}$ are not always trivial to express, but SymBoltz automatically expands them using the definitions of $\mathscr{H}$ or $\Psi$.
-Of course, one can eliminate all observeds by explicitly inserting them into the equations for the unknowns.
-But this reduces readability, as observed variables are helpful intermediate definitions that break up the equations, and one may want to extract them from the solution as well.
-Furthermore, modified models can change the sets of unknown and observed variables (e.g.
-modified gravity can change the constraint equation for $\Psi$ into a differential equation).
-It is easier to compose models when variables are not hardcoded as either unknown or observed.
-SymBoltz reads in a full system of equations like that defined by appendix A and separates variables into unknowns and observeds.
-It then generates numerical code for solving the ODE for the unknowns (see section 2.1.1), but can automatically recompute any observed variables from the solution of the unknowns.
+The Einstein–Boltzmann equations are usually written with many observed variables.
+For example, consider the metric and gravity equations in Appendices A.1 and A.2 sourced by some known $\rho(\tau)$, $\delta\rho(\tau,k)$, and $\Pi(\tau,k)$.
+Here, $a(\tau)$ and $\Phi(\tau,k)$ are the only unknown (differential) variables that the ODE is integrated for, whereas $z(\tau)$, $\mathscr{H}(\tau)$, $H(\tau)$, and $\Psi(\tau,k)$ are all observed (algebraically) from the unknowns.
+Furthermore, it is not always straightforward to express observed derivatives such as $\mathscr{H}^{\prime}$ or $\Psi^{\prime}$, but SymBoltz automatically expands them using the definitions of $\mathscr{H}$ or $\Psi$.
+Of course, we can eliminate all observeds by explicitly inserting them into the equations for the unknowns.
+However, this reduces readability, as observed variables are helpful intermediate definitions that break up the equations, and we may want to extract them from the solution as well.
+Furthermore, modified models can change the sets of unknown and observed variables (e.g., modified gravity can change the constraint equation for $\Psi$ into a differential equation).
+It is easier to build models when variables are not hardcoded as either unknown or observed.
+SymBoltz reads a full system of equations as input, such as that defined by the entirety of Appendix A, and automatically separates unknown and observed variables.
+After solving the ODE for its unknowns, SymBoltz can automatically recompute any observed variable from its expansion in unknowns.
 The bottom line is that users can easily use any variable anywhere by just referring to it, whether it is unknown or observed.
-In other solvers one may have to look up the expression for observed variables and recompute them manually.
-This is tedious, error-prone and can tempt users to make simplifying assumptions (e.g.
-approximating $\Psi\approx\Phi$ without anisotropic stress).
+In other solvers it can be necessary to look up observed expressions and recompute them manually.
+This is tedious, error-prone, and can tempt users to unnecessarily make simplifying assumptions (e.g., approximating $\Psi\approx\Phi$ without anisotropic stress).
 2.1.3 Automatic stage separation and splining of unknowns
-In principle, one can solve the Einstein-Boltzmann equations by integrating the entire system (e.g.
-background and perturbations) at once.
-However, the system can be broken down into sequential computational stages that each depend only on those before it.
-To alleviate stiffness in each stage, separate concerns by solving every equation only once (e.g.
-avoid recomputing the background for every perturbation mode) and to improve performance, all Boltzmann codes solve the system stage-by-stage and spline variables from one stage as input to the next.
-To illustrate, again consider the general relativistic equations in section A.2 sourced by some known $\rho(\tau)$, $\delta\rho(\tau,k)$ and $\Pi(\tau,k)$.
-Clearly $\Phi(\tau,k)$ and $\Psi(\tau,k)$ depend on $a(\tau)$, but $a(\tau)$ does not depend on $\Phi(\tau,k)$ or $\Psi(\tau,k)$, reflecting the perturbative nature of the problem.
-One can first solve for only $a(\tau)$ in the ?background?, then spline and look up $a(\tau)$ to solve for $\Phi(\tau,k)$ and $\Psi(\tau,k)$ in the ?perturbations?, instead of solving for all three together and repeatedly solve for $a(\tau)$ for every $k$.
-SymBoltz uses the same stage separation strategy as any other Boltzmann code, but automates it with its knowledge of the symbolic equations.
+In principle, the entire Einstein–Boltzmann system (i.e., background and perturbations) can be integrated all at once.
+However, it can be broken down into sequential computational stages that each depend only on those before it.
+To alleviate stiffness in each stage, avoid recomputing the background for every perturbation mode, and to improve performance by integrating smaller ODEs, all Boltzmann codes solve the system stage-by-stage and spline variables from one stage as input to the next.
+To illustrate this case, again consider the general relativistic equations in Appendix A.2 sourced by some known $\rho(\tau)$, $\delta\rho(\tau,k)$, and $\Pi(\tau,k)$.
+Clearly, $\Phi(\tau,k)$ and $\Psi(\tau,k)$ depend on $a(\tau)$, but $a(\tau)$ does not depend on $\Phi(\tau,k)$ or $\Psi(\tau,k)$, reflecting the perturbative nature of the problem.
+We can first solve for only $a(\tau)$ in the ?background,?
+then spline and look up $a(\tau)$ to solve for $\Phi(\tau,k)$ and $\Psi(\tau,k)$ in the ?perturbations,?
+instead of solving for all three together and repeatedly integrate $a(\tau)$ for every $k$.
+SymBoltz uses the same stage separation strategy as other Boltzmann codes, but automates it with knowledge of the symbolic equations.
 First, all equations are split into background and perturbation stages.
-Cubic Hermite splines are then constructed for all background unknowns (like $a(\tau)$) and passed on to the perturbations.
-This spline type is optimal for interpolating ODEs, as Hermite splines take both $\boldsymbol{u}(\tau)$ and $\boldsymbol{u}^{\prime}(\tau)$ into account for better accuracy, and $\boldsymbol{u}^{\prime}(\tau)$ is known analytically from $\boldsymbol{u}(\tau)$ and definition (3).
-In contrast, observed variables (like $\mathscr{H}(\tau)$) are computed from the (splined) unknowns because their derivatives are not known directly, and splining them is less accurate.
-This makes any background variable available in the perturbations ?for free?.
-All stages of the Einstein-Boltzmann equations are solved as if it is written as one combined system of equations, while splining and stage separation are done automatically under the hood.
-The separation into background and perturbation stages is guaranteed and just reflects the perturbative structure of the linearized Einstein-Boltzmann equations, where each order depends only on those before it.
-However, they can often be broken further down: most thermodynamics (recombination) models can be separated from the background, and some variables have integral solutions (e.g.
-the optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\bar{\tau})\mathrm{d}\bar{\tau}$ or line-of-sight integration) that can be computed in isolation after solving the differential equations.
-In the future, SymBoltz’ background-perturbations separation could be extended to automatically split all smaller stages by symbolically inspecting the dependencies between variables in the equations.
+A Cubic Hermite spline is then constructed for all background unknowns (e.g., $a(\tau)$, $X_{\mathrm{H}}^{+}(\tau)$, $X_{\mathrm{He}}^{+}(\tau)$, $T_{b}(\tau)$, and $\kappa(\tau)$) and replaces the background unknowns in the perturbations.
+One vector-valued spline is used to avoid repeatedly looking up $\tau$ for every unknown.
+Hermite splines are optimal for interpolating ODEs, as they take both $\boldsymbol{u}(\tau)$ and $\boldsymbol{u}^{\prime}(\tau)$ into account for better accuracy, and $\boldsymbol{u}^{\prime}(\tau)$ is known analytically from $\boldsymbol{u}(\tau)$ through the ODE (3).
+In contrast, observed variables (e.g., $\mathscr{H}(\tau)$) are computed from the (splined) unknowns, as their derivatives are not known directly and splining them is less accurate.
+This makes any background variable available in the perturbations ?for free.?
+All stages of the Einstein–Boltzmann equations are solved as if they are one common system of equations, while stage separation is done automatically under the hood.
+The separation into background and perturbation stages is guaranteed.
+It just reflects the perturbative structure of the linear Einstein–Boltzmann equations, where each order depends only on lower orders.
+However, they can often be broken further down: most thermodynamics (recombination) models can be separated from the background, and integral solutions (e.g., the optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\bar{\tau})\mathrm{d}\bar{\tau}$ and line-of-sight integration) can be computed after solving the differential equations.
+Future work could extend SymBoltz’ background-perturbations separation to split all finer stages by symbolically inspecting the dependencies between variables in the equations.
 2.1.4 Automatic solution interpolation
-SymBoltz integrates the background and several perturbation ODEs in conformal time $\tau$ for several wavenumbers $k$.
-All the results are stored in a special solution object.
-This solution object can be queried for any variable or symbolic expression
+SymBoltz integrates the background and perturbation ODEs in conformal time $\tau$ and for several wavenumbers $k$.
+All the results are stored in a dedicated solution object.
+This object can be queried for any variable or symbolic expression
 $S(\tau)\quad\text{or}\quad S(\tau,k).$ (4)
 This looks up the background solution if $S$ is a background variable, or the perturbation solutions if it is perturbative.
-It interpolates from the times stored by the ODE solver to the requested $\tau$ using the solver’s dense interpolation.
+It interpolates from the times stored by the ODE solver to the requested $\tau$ using the solver’s dense interpolation scheme.
 If $S$ is perturbative, it also interpolates from the solved perturbation $k$-modes to the requested $k$.
 If $S$ is an unknown, it is returned directly from the ODE solution.
-If $S$ is observed, it is instead recomputed automatically from the equation that defines it as a function of the unknowns.
-The result is that the user can access any variable in the model without having to do this manually.
-The solution interpolation is also incorporated into plot recipes for visualization of any variable as a function of $\tau$ and/or $k$.
+If $S$ is observed, it is instead recomputed automatically from its expansion in terms of unknowns.
+The result is that the user can access any variable in the model without having to do any of this manually.
+The solution interpolation is also incorporated into plotting recipes that easily visualize any variable as a function of $\tau$ and $k$.
 2.1.5 Automatic Jacobian generation and sparsity detection
-Just like SymBoltz generates code for $\boldsymbol{f}$ in the ODE (3), it uses the same symbolic equations to generate its Jacobian $\boldsymbol{J}$ with entries
+Just as SymBoltz generates code for $\boldsymbol{f}$ in the ODE (3), it uses the same symbolic equations to generate its Jacobian $\boldsymbol{J}$ with entries
 $J_{ij}=\frac{\partial f_{i}}{\partial u_{j}}.$ (5)
@@
 Manually coding them is tiresome, error-prone, and must be repeated for new models.
-Numerical evaluation with finite differences is both approximate and slow.
-It can also be computed with automatic differentiation, but we find this slower and overkill compared to the analytical $\boldsymbol{J}$.
-Analytical Jacobians are as fast and stable as possible, and automatic generation lets the solver benefit from it while the modeler focuses on the high-level equations.
+Numerical evaluation with finite differences is approximate and slow.
+Automatic differentiation can compute $\boldsymbol{J}$ from $\boldsymbol{f}$, but makes it harder to exploit the sparsity of $\boldsymbol{J}$.
+Analytical Jacobians are as fast and stable as possible, and SymBoltz generates them automatically for the solver without distracting the modeler from focusing on the physical equations.
 SymBoltz stores the Jacobian in sparse form to boost performance when it has many zeros.
 From the analytical $\boldsymbol{J}$, SymBoltz precomputes its exact sparsity pattern (where $J_{ij}=0$).
-This is hard to specify manually, and numerical solvers cannot distinguish false local zeros (for some input) from true global zeros (for all inputs).
-Without approximations, SymBoltz then reuses the same sparsity structure throughout the integration, computes the nonzero $J_{ij}$ analytically and stores them directly in the sparse $\boldsymbol{J}$ with no extra cost from looking up sparse matrix indices.
-Easy, accurate, fast and sparse Jacobians are crucial for SymBoltz to remain performant without approximations! It is a major advantage with the symbolic approach, as in PyCosmo.
+This is hard to specify manually, and numerical solvers cannot distinguish false local zeros (for some inputs) from true global zeros (for all inputs).
+Without approximations, SymBoltz reuses the same sparsity structure throughout the integration, computes the nonzero $J_{ij}$ analytically, and stores them at a precomputed index in the sparse $\boldsymbol{J}$.
+This avoids converting $\boldsymbol{J}$ from dense to sparse form at runtime.
+Easy analytical and sparse Jacobians are crucial for SymBoltz to remain fast without approximations.
+This is a major advantage with the symbolic approach, as in PyCosmo.
 2.2 Approximation-freeness
-SymBoltz treats stiffness in the Einstein-Boltzmann equations with modern implicit ODE solvers that integrate the full equations at all times.
-It is therefore free of approximation schemes, such as the TCA, UFA, RSA, NCDMFA and Saha approximation.
-This is friendlier to the modeler, who now simply has to provide a single set of equations instead of deriving approximations, validating them and dealing with related chores.
-This pairs well with a high-level equation-oriented symbolic interface.
+SymBoltz treats stiffness in the Einstein–Boltzmann equations with modern implicit ODE solvers that integrate the full equations at all times.
+It is therefore free of approximation schemes, such as the TCA, UFA, RSA, NCDMFA, and Saha approximation.
+This is friendlier to the modeler, who now has to provide only one set of equations, instead of deriving, implementing, and validating approximations.
+This approach is also well-suited to the high-level equation-oriented symbolic interface.
 Implicit solvers take more expensive steps than explicit methods.
-At every step, they solve a generally nonlinear system of equations for the next unknowns $\boldsymbol{u}$.
-Using Newton’s method, it iteratively solves linear systems $\boldsymbol{W}\boldsymbol{u}=\boldsymbol{b}$ by LU-factorizing $\boldsymbol{W}=\boldsymbol{I}-\gamma h\boldsymbol{J}$, where $\gamma$ is a constant, $h$ is the time step and $\boldsymbol{J}$ is the ODE Jacobian (5).
+At every step, they solve a generally nonlinear system of equations for the next unknowns $\boldsymbol{u}$, often using Newton’s method.
+It iteratively solves linear systems $\boldsymbol{W}\boldsymbol{u}=\boldsymbol{b}$ by LU-factorizing $\boldsymbol{W}=\boldsymbol{I}-\gamma h\boldsymbol{J}$, where $\gamma$ is a constant, $h$ is the time step, and $\boldsymbol{J}$ is the ODE Jacobian (5).
 [^16] For example, the implicit Euler method $\boldsymbol{u}_{n+1}=\boldsymbol{u}_{n}+h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})$ solves $\boldsymbol{F}(\boldsymbol{u}_{n+1})=\boldsymbol{u}_{n+1}-h\boldsymbol{f}(\tau_{n+1},\boldsymbol{u}_{n+1})-\boldsymbol{u}_{n}=\boldsymbol{0}$ at every step with Newton’s method using its Jacobian $\boldsymbol{W}=\boldsymbol{\nabla}\boldsymbol{F}=\boldsymbol{I}-h\boldsymbol{J}(\tau_{n+1},\boldsymbol{u}_{n+1})$.
-LU-factorizing a dense $n\times n$ matrix from an $n$-dimensional ODE costs $O(n^{3})$ operations and bottlenecks larger ODEs (e.g.
-larger $l_{\text{max}}$).
-In return, implicit methods use more information about the ODE to take stable and long steps.
+LU-factorizing a dense $n\times n$ matrix from an $n$-dimensional ODE costs $O(n^{3})$ operations and bottlenecks larger ODEs (e.g., larger $l_{\text{max}}$).
+In return, implicit methods use more information to take long and stable steps.
 To be fast, implicit solvers need several optimizations.
@@
 A common trick is to reuse an LU-factorized $\boldsymbol{W}$ over several steps, even if it really changes.
-Newton’s method only needs an approximate $\boldsymbol{W}$ to find the root, and $\boldsymbol{W}$ is recomputed only if convergence is slow due to an outdated $\boldsymbol{J}$ or a new step size $h$.
+Newton’s method only requires an approximate $\boldsymbol{W}$ to find the root, and $\boldsymbol{W}$ is recomputed only if convergence is slow due to an outdated $\boldsymbol{J}$ or a changed step size $h$.
 This trades fewer LU-factorizations and $\boldsymbol{J}$ evaluations for more linear solve back-substitutions, which only cost $O(n^{2})$ operations.
 The next trick is to speed up the linear algebra with sparse matrices if $\boldsymbol{W}$ has many zeros, as for the perturbations.
-SymBoltz handles this by generating the analytical and sparse Jacobian (see section 2.1.5).
-SymBoltz generates code for all implicit solvers in OrdinaryDiffEq.jl [^11] https://github.com/SciML/OrdinaryDiffEq.jl/ and dense and sparse matrix methods in LinearSolve.jl [^12] https://github.com/SciML/LinearSolve.jl.
+SymBoltz handles this by generating the analytical and sparse Jacobian (see Sect.
+2.1.5).
+SymBoltz generates code for all implicit solvers in OrdinaryDiffEq, and dense and sparse matrix methods in LinearSolve.
 We can test different solvers, which is useful as performance and stability of implicit methods and linear algebra operations are problem-dependent.
-Using $O(100)$ perturbation equations with more than $95\%$ sparsity, we find RFLUFactorization fastest for dense matrices, but KLUFactorization several times faster when exploiting sparsity.
-We tested these ODE solvers:
-  • Backward Differentiation Formula (BDF) methods like FBDF and QNDF (Shampine and Reichelt, 1997) are multi-step methods that use several past points to solve for the next step.
-They do not have Runge-Kutta ?stages?
-and solve just one implicit equation per step, so they scale well to large ODEs and can reuse Jacobians aggressively.
+Using $O(100)$ perturbation equations with more than $95\%$ sparsity, we find RFLUFactorization in LinearSolve fastest for dense matrices, but KLUFactorization several times faster when exploiting sparsity.
+We tested these ODE solvers from OrdinaryDiffEq:
+  • Backward differentiation formula (BDF) methods, such as FBDF and QNDF (Shampine and Reichelt, 1997), are multi-step methods that use several past points to solve for the next.
+They do not have Runge–Kutta ?stages?
+and solve just one implicit equation per step, so they scale well to large ODEs and can reuse Jacobians heavily.
 They have variable order of 1-5, but are ?L-stable?
-only up to 2nd order.
-This makes them take small steps in stiff regimes without utilizing their high order.
+only up to second order.
+This makes them take short steps in stiff regimes without utilizing their high orders.
 We find them to be slowest.
-The solvers ndf15 in CLASS and BDF2 in PyCosmo belong to this family.
-Nadkarni-Ghosh and Refregier (2017) also discuss BDF methods in the context of the Einstein-Boltzmann equations.
-  • Explicit Singly-Diagonal Implicit Runge-Kutta (ESDIRK) methods (e.g.
-Jørgensen et al., 2018) are a class of implicit Runge-Kutta methods designed to preserve most stability properties of Fully Implicit (FIRK) methods at a cheaper computational cost.
-An $s$-stage FIRK method has excellent stability, but a dense Butcher tableau $a_{ij}$ and must solve a system of $n\times s$ equations at every step.
+The solvers ndf15 in CLASS and BDF2 in PyCosmo also belong to this family.
+Nadkarni-Ghosh and Refregier (2017) also discuss BDF methods in the context of the Einstein–Boltzmann equations.
+  • Explicit singly diagonal implicit Runge–Kutta (ESDIRK) methods (e.g., Jørgensen et al., 2018) are a class of implicit Runge–Kutta methods designed to preserve most stability properties of fully implicit (FIRK) methods at a cheaper computational cost.
+An $s$-stage FIRK method has excellent stability, but a dense Butcher tableau $a_{ij}$, and must solve a system of $n\times s$ equations at every step.
 D in ESDIRK means that $a_{ij}$ has only lower triangular and diagonal nonzeros, decoupling the system into $s$ systems of $n$ equations that are faster to solve sequentially.
-S means that all diagonal nonzeros $a_{ii}=\gamma$ are equal, so all stages share the same $\boldsymbol{W}$-matrix and only one must be factorized and can be reused as in BDF methods.
+S means that all diagonal nonzeros $a_{ii}=\gamma$ are equal, so all stages share the same $\boldsymbol{W}$-matrix, and only one must be factorized and can be reused as in BDF methods.
 E means that $a_{11}=0$, so the first stage is explicit and cheap.
 Unlike BDF methods, they can remain L-stable without losing order to stiffness.
-We find the 4th order KenCarp4 method (Kennedy and Carpenter, 2003) to perform well with much fewer time steps.
-Bolt.jl also defaults to this solver.
-DISCO-EB uses Kvaerno5, but we find it slower.
-  • Rosenbrock methods: Rosenbrock methods (e.g.
-Lang, 2020) linearize $\boldsymbol{f}\approx\boldsymbol{J}\boldsymbol{u}$ and effectively replace Newton’s method with one linear solve.
-Their major upside is being free from convergence on a nonlinear problem.
-This suits the perturbation ODEs, as they are already linear.
-The downside is that they can need a new $\boldsymbol{J}$ at every step, which is slow to evaluate numerically in many problems.
-But generating the Jacobian analytically eliminates this issue.
-A subclass called ?Rosenbrock-W?
-methods can reuse Jacobians, but this strategy is not yet used by OrdinaryDiffEq.jl.
-We find the 5th order L-stable Rodas5P method (Steinebach, 2023) to be both most efficient and stable on both the background and perturbations, and have not found it used by other Boltzmann codes.
-Performance of these methods is compared in section 4.
+We find the fourth order KenCarp4 method (Kennedy and Carpenter, 2003) to perform well with much fewer time steps.
+Bolt also defaults to it.
+DISCO-EB uses Kvaerno5, but we find it to be slower.
+  • Rosenbrock methods (e.g., Lang, 2020) linearize $\boldsymbol{f}\approx\boldsymbol{J}\boldsymbol{u}$ and effectively replace Newton’s method with one linear solve.
+Being free from nonlinear convergence issues is a major upside and suits the linear perturbation ODEs.
+The downside is that they can need a new $\boldsymbol{J}$ at every step, which is slow to evaluate in many problems.
+But generating $\boldsymbol{J}$ analytically eliminates this issue.
+The subclass of ?Rosenbrock-W?
+methods can reuse Jacobians, but this strategy is not yet used by OrdinaryDiffEq.
+We find the fifth order L-stable Rodas5P method (Steinebach, 2023) to be both most efficient and stable on both the background and perturbations, and have not found it used by other Boltzmann codes.
+Section 4 compares the performance of these methods.
 Traditional codes use approximations to reduce stiffness and increase performance by evolving fewer ODE states when possible.
 Earlier works find marginal speedups from the TCA and UFA compared to using implicit solvers, but the RSA can significantly speed up models with high $l_{\text{max}}$ (Lesgourgues and Tram, 2011; Moser et al., 2022; Hahn et al., 2024).
-However, approximation-based codes must account for the structure of the equations to change, adding overhead to ODE solvers and restructuring sparse Jacobians.
-Approximation-free solvers can challenge this by optimizing around constant structure.
+However, approximation-based codes must account for the structure of the equations to change, which adds overhead to ODE solvers and restructuring of sparse Jacobians.
+Approximation-free solvers can challenge this by optimizing around unchanged structure.
 Numerical codes approximate $\boldsymbol{J}$ with finite differences using $O(n)$ evaluations of $\boldsymbol{f}$.
-This is wasteful, as $\boldsymbol{f}=\boldsymbol{J}\boldsymbol{u}$ for linear perturbations and computing the nonzeros of $\boldsymbol{J}$ requires fewer operations than $\boldsymbol{f}$!
+This is wasteful, as $\boldsymbol{f}=\boldsymbol{J}\boldsymbol{u}$ for linear perturbations, so computing the nonzeros of $\boldsymbol{J}$ should take fewer operations than $\boldsymbol{f}$!
 Symbolic codes can take advantage of such properties.
-Approximations also complicate the code and put more load on modelers to derive, implement and validate them, and repeat the process for modifications that reintroduce stiffness or invalidate approximations.
+Approximations also complicate the code and put more load on modelers to derive, implement, and validate them, and repeat the process as modifications reintroduce stiffness or invalidate approximations.
 SymBoltz is tuned for performance without invoking approximations.
@@
 SymBoltz is compatible with automatic differentiation.
-It can compute derivatives of any output quantity (e.g.
-ODE variables, $P(k)$ or $C_{l}$) with respect to any input parameters (e.g.
-$h$ or $\Omega_{c0}$).
-For example, it can be used to learn the output’s sensitivity to the input, perform Fisher forecasts or compute the Jacobian in the shooting method when parametrizing models with integrated variables.
-It can also compute the ODE Jacobian, but we do this symbolically to get it analytically with its sparsity pattern.
-Currently, SymBoltz and Bolt.jl both work only with forward-mode dual numbers through ForwardDiff.jl [^13] https://github.com/JuliaDiff/ForwardDiff.jl.
-This is appropriate for applications with more outputs than inputs, like differentiating $O(100)$ values of $P(k)$ or $C_{l}$ as functions of $O(10)$ parameters.
+It can compute derivatives of any output quantity (e.g., ODE variables, $P(k)$, or $C_{l}$) with respect to any input parameters (e.g., $h$ or $\Omega_{c0}$).
+For example, it can be used to learn the sensitivity of the output to the input, perform Fisher forecasts, or compute the Jacobian in the shooting method for boundary-value parameterized models.
+It can also compute the ODE Jacobian, but SymBoltz does this symbolically to get it analytically with its sparsity pattern.
+Currently, SymBoltz and Bolt both work only with forward-mode dual numbers through ForwardDiff.
+This is appropriate for applications with more outputs than inputs, such as differentiating $O(100)$ values of $P(k)$ or $C_{l}$ as functions of $O(10)$ parameters.
 Reverse-mode is ideal with fewer outputs and attractive for MCMCs or training emulators where output is compressed to a scalar loss.
-Fast reverse-mode gradients could accelerate sampling in large parameter spaces from upcoming surveys and challenge use of emulators.
+Fast reverse-mode gradients could accelerate sampling in large parameter spaces from upcoming surveys and challenge the use of emulators.
 Notably, DISCO-EB supports reverse-mode, but has not yet demonstrated it for this purpose.
@@
 3 Examples
-The best way to illustrate the features in section 2 is perhaps through some examples. The code is as of SymBoltz version 0.14.1 and available in a notebook in the project repository.
+The features in Sect. 2 are best illustrated through some examples. This code is as of SymBoltz version 1.0.0 and available in a notebook in the project repository.
 3.1 Basic usage workflow
-Figure 2: Included plot recipes in SymBoltz make it trivial to visualize any symbolic variable or expression thereof from a solution of the Einstein-Boltzmann equations.
+Figure 2: SymBoltz includes plotting recipes that make it easy to visualize any symbolic variable or expression thereof from a solution of the Einstein–Boltzmann equations.
 This plot was made with one short line of code per subplot.
 Wavenumbers $k$ are in units of $H_{0}/c$.
-Most usage follows a model–problem–solution workflow: {codebox}
-using SymBoltzM = ΛCDM(lmax = 16)p = Dict( M.γ.T₀ => 2.7, M.b.Ω₀ => 0.05, M.b.YHe => 0.25, M.ν.Neff => 3.0, M.c.Ω₀ => 0.27, M.h.m_eV => 0.02, M.I.ln_As1e10 => 3.0, M.I.ns => 0.96, M.g.h => 0.7)prob = CosmologyProblem(M, p; jac=true, sparse=true)ks = [4, 40, 400, 4000] # k / (H₀/c)sol = solve(prob, ks)The model–problem–solution split achieves three distinct goals.
-First, a purely symbolic representation M of the $\mathrm{\Lambda CDM}$ model is created.
+Most usage follows a model-problem-solution workflow: {codebox}
+using SymBoltzM = ΛCDM(lmax = 16)p = Dict( M.γ.T₀ => 2.7, M.b.Ω₀ => 0.05, M.b.YHe => 0.25, M.ν.Neff => 3.0, M.c.Ω₀ => 0.27, M.h.m_eV => 0.02, M.I.ln_As1e10 => 3.0, M.I.ns => 0.96, M.g.h => 0.7)prob = CosmologyProblem(M, p; jac=true, sparse=true)ks = [4, 40, 400, 4000] # k / (H₀/c)sol = solve(prob, ks)The model-problem-solution split achieves three distinct goals.
+First, a symbolic representation M of the $\mathrm{\Lambda CDM}$ model is created.
 This is a standalone object designed to be interactively inspected and modified.
-It contains every variable, parameter and equation of the cosmological model structured as one submodel per physical component: the metric $g$, the gravitational theory $G$, photons $\gamma$, massless neutrinos $\nu$, massive neutrinos $h$, cold dark matter $c$, baryons $b$, the cosmological constant $\Lambda$ and inflation $I$.
-For example, equations(M) shows all equations; equations(M.G) only the gravitational ones; M.g.a refers to the scale factor variable $a(\tau)$ that ?belongs?
-to the metric $g$; M.b.Ω₀ is the baryon density parameter $\Omega_{b0}$; and parameters(M) lists parameters that can be set.
-This displays with -compatibility in notebooks and transparently shows the equations being solved.
+It contains every variable, parameter, and equation of the model structured as one submodel per physical component: the metric $g$, gravitational theory $G$, photons $\gamma$, massless neutrinos $\nu$, massive neutrinos $h$, cold dark matter $c$, baryons $b$, the cosmological constant $\Lambda$, and inflation $I$.
+For example, equations(M) shows all its equations; equations(M.G) only the gravitational ones; M.g.a refers to the scale factor variable $a(\tau)$ that ?belongs?
+to the metric $g$; M.b.Ω₀ is the baryon density parameter $\Omega_{b0}$; and parameters(M) shows all parameters that can be set.
+Everything displays with -compatibility in notebooks and transparently shows the contents of the model.
 Second, the symbolic model is compiled to a numerical problem prob with parameters p.
-This is an expensive step where the processing in section 2.1 happens: equations are checked for consistency and split into background/perturbation stages; observed/unknown variables are distinguished; fast ODE code is generated; background unknowns in the perturbations are replaced by splines; the Jacobian matrix is generated in numerical/analytical and dense/sparse form; and initial values are computed.
-The compilation is customized with optional arguments like jac and sparse.
-This is a separate step because it performs final transformations on the model M once the user has finished modifying and committed to it.
+This expensive step does the processing in Sect.
+2.1: it checks equations for consistency; splits them into background and perturbation stages; distinguishes observed and unknown variables; generates fast code for ODEs and Jacobians in numerical or analytical and dense or sparse form; splines background unknowns in the perturbations; and finds initial conditions.
+Optional keyword arguments customize the compilation and control precisely how the problem is solved.
+This is a separate step because it does final transformations on the model M once the user has finished modifying and committed to it.
 Third, solve(prob, ks) solves the background and perturbations for the wavenumbers ks.
@@
 The resulting solution object sol provides convenient access to all model variables.
-Internally, it stores values of all ODE unknowns at the time steps taken by the solvers and for every requested wavenumber.
-But it can be queried with any time, wavenumber and symbolic expression, and will automatically compute it from the unknowns and interpolate between stored times and wavenumbers, as explained in section 2.1.4.
-For example, calling sol(1.0, 2.0, g.Φ+g.Ψ) computes $\Phi+\Psi$ at $k=1\,H_{0}/c$ and $\tau=2\,H_{0}^{-1}$ by expanding it in terms of unknowns using the equations in section A.2.
-This interface was used to plot the evolution of several variables in fig.
-2 with little code.
+Internally, it stores values of all ODE unknowns at the time steps taken by the solvers for every requested wavenumber.
+However, it can be queried with any time, wavenumber, and symbolic expression, and will automatically compute it from the unknowns and interpolate between stored times and wavenumbers, as explained in Sect.
+2.1.4.
+For example, sol(g.Φ+g.Ψ, 1.0, 2.0) computes $\Phi+\Psi$ at $\tau=1\,H_{0}^{-1}$ and $k=2\,H_{0}/c$ by expanding $\Psi$ in terms of $\Phi$ and other unknowns.
+This interface is also used to plot the evolution of several variables in Fig.
+2 with very compact code.
 Most Boltzmann solvers save only a fixed set of variables, such as only the unknowns.
 Recovering observed variables requires manual effort.
-This is cumbersome, opens for user error and adds friction in the modeling process.
+This is cumbersome, open to user error, and adds friction in the modeling process.
 SymBoltz is designed to provide easy access to all variables defined by the model.
@@
 $\rho(\tau)=\rho(\tau_{0})a(\tau)^{-3(1+w_{0}+w_{a})}\exp(-3w_{a}(1-a(\tau))).$ (7b)
-Perturbations are given by de Putter et al. (2010). To implement this species in SymBoltz, write down all related variables, parameters, equations and initial conditions in one place: {codebox}
-g, τ, k = M.g, M.τ, M.ka, ℋ, Φ, Ψ = g.a, g.ℋ, g.Φ, g.ΨD = Differential(τ)@parameters w₀ wₐ cₛ² Ω₀ ρ₀@variables ρ(τ) P(τ) w(τ) cₐ²(τ) δ(τ,k) θ(τ,k) σ(τ,k)eqs = [ w ~ w₀ + wₐ*(1-a) ρ₀ ~ 3*Ω₀ / (8*Num(π)) ρ ~ ρ₀ * a^(-3(1+w₀+wₐ)) * exp(-3wₐ*(1-a)) P ~ w * ρ cₐ² ~ w - 1/(3ℋ) * D(w)/(1+w) D(δ) ~ 3ℋ*(w-cₛ²)*δ - (1+w) * ( (1+9(ℋ/k)^2*(cₛ²-cₐ²))*θ - 3*D(Φ)) D(θ) ~ (3cₛ²-1)*ℋ*θ + k^2*cₛ²*δ/(1+w) + k^2*Ψ σ ~ 0]initialization_eqs = [ δ ~ -3//2 * (1+w) * Ψ θ ~ 1//2 * (k^2*τ) * Ψ]X = System(eqs, τ; initialization_eqs, name = :X)Everything is packed down in the $w_{0}w_{a}$ component X.
-Unicode code symbols are encouraged to maximize similarity with equations (e.g.
-Ω₀ over Omega_0), and SymBoltz will automatically render them in -compatible environments (e.g.
-notebooks).
-A $w_{0}w_{a}\text{CDM}$ model is now built by replacing the cosmological constant species Λ in $\Lambda\text{CDM}$ by the $w_{0}w_{a}$ species X: {codebox}
+Perturbations are given by de Putter et al. (2010). To implement this species in SymBoltz, we write down all related variables, parameters, equations, and initial conditions in one place: {codebox}
+g, τ, k = M.g, M.τ, M.ka, ℋ, Φ, Ψ = g.a, g.ℋ, g.Φ, g.ΨD = Differential(τ)@parameters w₀ wₐ cₛ² Ω₀ ρ₀=3Ω₀/8π@variables ρ(τ) P(τ) w(τ) cₐ²(τ) δ(τ,k) θ(τ,k) σ(τ,k)eqs = [ w ~ w₀ + wₐ*(1-a) ρ ~ ρ₀ * a^(-3(1+w₀+wₐ)) * exp(-3wₐ*(1-a)) P ~ w * ρ cₐ² ~ w - 1/(3ℋ) * D(w)/(1+w) D(δ) ~ 3ℋ*(w-cₛ²)*δ - (1+w) * ( (1+9(ℋ/k)^2*(cₛ²-cₐ²))*θ - 3*D(Φ)) D(θ) ~ (3cₛ²-1)*ℋ*θ + k^2*cₛ²*δ/(1+w) + k^2*Ψ σ ~ 0]initialization_eqs = [ δ ~ -3//2 * (1+w) * Ψ θ ~ 1//2 * (k^2*τ) * Ψ]X = System(eqs, τ; initialization_eqs, name=:X)
+The last line packs everything into a $w_{0}w_{a}$ component X.
+Unicode symbols are encouraged to maximize similarity with equations (e.g., $\Omega_{0}$ over Omega_0), and SymBoltz automatically renders them in -compatible environments (e.g., notebooks).
+A full $w_{0}w_{a}\text{CDM}$ model is now built by replacing the cosmological constant species $\Lambda$ in $\Lambda\text{CDM}$ by the $w_{0}w_{a}$ species X: {codebox}
 M = ΛCDM(Λ=X, lmax=16, name=:w₀wₐCDM)push!(p, X.w₀ => -0.9, X.wₐ => 0.2, X.cₛ² => 1.0)prob = CosmologyProblem(M, p; jac=true, sparse=true)
-This is all the user has to do. The modification simply consists of writing down the equations verbatim. It could hardly be more compact and to the point. Under the hood, SymBoltz will:
-  • create input hooks for setting the parameters $w_{0}$, $w_{a}$, $c_{s}^{2}$, $\Omega_{0}$,
-  • move $(\tau)$-dependent functions to the background,
-  • move $(\tau,k)$-dependent functions to the perturbations,
-  • expand $\texttt{D(w)}=\mathrm{d}w/\mathrm{d}\tau$ and $\texttt{D(Φ)}=\mathrm{d}\Phi/\mathrm{d}\tau$ from $w$ and $\Phi$,
-  • compute needed background variables in the perturbations,
-  • spline $\rho(\tau)$ in the perturbations (if we use (7a) over (7b)),
-  • source gravity with energy-momentum contributions,
-  • assign $\Omega_{X0}=1-\sum_{s\neq{X}}\Omega_{s0}$ (if gravity is GR) by default,
-  • eliminate common subexpressions like x=1+wx\!=\!1+w and y=1/xy\!=\!1/x,
-  • generate ODE functions and state indices for $\delta^{\prime}$ and $\theta^{\prime}$,
-  • generate sparse analytical Jacobian entries for $\delta^{\prime}$ and $\theta^{\prime}$,
-  • interpolate and output any variable from the solution, both for unknowns (i.e. $\delta$ and $\theta$) and observeds (e.g. $w$ and $c_{a}^{2}$).
-Most of these steps require manual intervention in a numerical code.
-For example, in CLASS one would have to read new parameters in input.c; declare new background and perturbation variables in background.h and perturbations.h; solve background equations, save desired output and source gravity in background.c; and recompute or look up background variables, solve perturbation equations, save desired output and source gravity in perturbations.c.
+The modification consists of writing down the equations compactly and directly. This is all the user must do, and under the hood SymBoltz will:
+  • create input hooks for setting the parameters $w_{0}$, $w_{a}$, $c_{s}^{2}$, $\Omega_{0}$;
+  • move $(\tau)$-dependent functions to the background;
+  • move $(\tau,k)$-dependent functions to the perturbations;
+  • expand $\texttt{D(w)}=\mathrm{d}w/\mathrm{d}\tau$ and $\texttt{D($\Phi$)}=\mathrm{d}\Phi/\mathrm{d}\tau$ from $w$ and $\Phi$;
+  • compute needed background variables in the perturbations;
+  • spline $\rho(\tau)$ in the perturbations (if we use Eq. (7a) over (7b));
+  • source gravity with energy-momentum contributions;
+  • assign $\Omega_{X0}=1-\sum_{s\neq{X}}\Omega_{s0}$ by default (if gravity is GR);
+  • eliminate common subexpressions (e.g., $x\hskip-2.0pt=\hskip-2.0pt1\hskip-1.0pt+\hskip-1.0ptw$ and $y\hskip-2.0pt=\hskip-2.0pt1\hskip-1.0pt/\hskip-1.0ptx$);
+  • generate ODE functions and state indices for $\delta^{\prime}$ and $\theta^{\prime}$;
+  • generate analytical and sparse Jacobian entries for $\delta^{\prime}$ and $\theta^{\prime}$;
+  • interpolate and output any variable from the solution, both for unknowns (i.e., $\delta$ and $\theta$) and observeds (e.g., $w$ and $c_{a}^{2}$).
+More manual work is required in a purely numerical code.
+For example, in CLASS, we would have to read new parameters in input.c; declare new background and perturbation variables in background.h and perturbations.h; solve background equations, save desired output and source gravity in background.c; and recompute or look up background variables, solve perturbation equations, save desired output, and source gravity in perturbations.c.
 The changes are scattered across the code, so it grows in complexity as more models are added.
-SymBoltz opens for a workflow that implements and analyzes a modified model with less effort all in one notebook, for example.
-Of course, both codes already include a $w_{0}w_{a}$ species, but this is how they are implemented. We proceed with the $w_{0}w_{a}\text{CDM}$ model.
-Figure 3: Matter and CMB ($T$, TE and $E$) power spectra computed by SymBoltz (SB; colored lines) compared to CLASS (CL; grey dashes) with relative errors $\text{rel.err.}=P_{k}^{\text{SB}}/P_{k}^{\text{CL}}-1$ and $\text{rel.err.}=C_{l}^{\text{SB}}/C_{l}^{\text{CL}}-1$ for the $\Lambda\text{CDM}$ model.
+SymBoltz enables a workflow that implements and analyzes a modified model with less effort all in one notebook, for example.
+Of course, both codes already include a $w_{0}w_{a}$ species, but this is how they are implemented.
+We proceed with the $w_{0}w_{a}\text{CDM}$ model to the next section.
+Figure 3: Matter and CMB ($T$, $E$, and TE) power spectra computed by SymBoltz (colored lines) compared to CLASS (gray dashes) with relative errors $P(k)_{\text{SymBoltz}}/P(k)_{\text{CLASS}}-1$ and $C_{l,\text{SymBoltz}}/C_{l,\text{CLASS}}-1$ for the $w_{0}w_{a}\text{CDM}$ model.
 CLASS uses the precision parameters in Appendix B.
@@
 CLASS (all approximations) 0.5 s 1.6 s
-SymBoltz can compute the matter power spectrum $P(k)$ and the angular CMB power spectra $C_{l}^{\text{TT}}$, $C_{l}^{\text{TE}}$ and $C_{l}^{\text{EE}}$: {codebox}
-ks = 10 .^ range(-1, 4, length = 100)Ps = spectrum_matter(prob, ks)ls = vcat([2,3,5,10], 20:20:2500)jl = SphericalBesselCache(ls)Cls = spectrum_cmb([:TT, :EE, :TE], prob, jl)The spectrum_matter and spectrum_cmb functions select a coarse $k$-grid to solve the perturbations for and interpolates to a finer $k$-grid.
+SymBoltz can compute the matter power spectrum $P(k)$, and the angular CMB power spectra $C_{l}^{\text{TT}}$, $C_{l}^{\text{EE}}$, and $C_{l}^{\text{TE}}$: {codebox}
+ks = 10 .^ range(-1, 4, length = 100)Ps = spectrum_matter(prob, ks)ls = vcat([2,3,5,10], 20:20:2500)jl = SphericalBesselCache(ls)Cls = spectrum_cmb([:TT, :EE, :TE], prob, jl)The spectrum_matter and spectrum_cmb functions select a coarse $k$-grid to solve the perturbations for and interpolate to a finer $k$-grid.
 This is a common trick to integrate fewer $k$-modes.
 Precision parameters that affect the output are given with optional keyword arguments.
-Output spectra and their computation time with standard precision settings are shown in figs.
-3 and 1.
+Output spectra and their computation time with standard precision settings are shown in Fig.
+3 and Table 1.
 It shows that SymBoltz and CLASS agree to around $0.1\%$ or better.
 The $P(k)$ agree to $0.1\%$ for all $k$ and $0.01\%$ for linear $k$.
-The $C_{l}$ agree to $0.1\%$ for most $l$, but slightly worse for cosmic variance-dominated $l$ and very large $l$ in $C_{l}^{\text{TE}}$.
-This is sufficient for most current data (e.g.
-Bolliet et al., 2024).
-It is similar to CAMB vs.
-CLASS with standard precision, although their agreement is pushed to $0.01\%$ with high precision (Lesgourgues, 2011b).
-SymBoltz computes $P(k)$ quickly, but the extra steps needed for $C_{l}$ are not yet as optimized as they are in CLASS. A more thorough performance comparison for $P(k)$ is given in section 4.
+The $C_{l}$ agree to $0.1\%$ for most $l$, although slightly worse for cosmic variance-dominated $l$ and very large $l$ in $C_{l}^{\text{TE}}$.
+This is sufficient for most current data (e.g., Bolliet et al., 2024).
+It is similar to CAMB versus CLASS with standard precision, although their agreement is pushed to $0.01\%$ with high precision (Lesgourgues, 2011b).
+SymBoltz computes $P(k)$ efficiently, but the extra steps needed for $C_{l}$ are not yet as optimized as they are in CLASS.
+Section 4 gives a more thorough performance comparison for $P(k)$.
 3.4 Differentiable Fisher forecasting
 Fisher forecasting is a technique to predict how strong parameter constraints that can be placed by data with given errors. It needs derivatives that can be computed with automatic differentiation.
-Near a peak $\boldsymbol{p}_{0}$, where derivatives vanish, a log-likelihood $\log L$ of parameters $\boldsymbol{p}$ is approximated by the Taylor series
-$\log L(\boldsymbol{p})\approx\log L(\boldsymbol{p}_{0})-\sum_{i,j}F_{ij}(\boldsymbol{p}_{0})(p_{i}-\bar{p}_{i})(p_{j}-\bar{p}_{j}),$ (8)
-where $\boldsymbol{F}$ is the Fisher (information) matrix with elements
+Near a peak $\bar{\boldsymbol{p}}$, where derivatives vanish, a log-likelihood $\log L$ of parameters $\boldsymbol{p}$ is approximated by the Taylor series
+$\log L(\boldsymbol{p})\approx\log L(\bar{\boldsymbol{p}})-\sum_{i,j}F_{ij}(\bar{\boldsymbol{p}})(p_{i}-\bar{p}_{i})(p_{j}-\bar{p}_{j}),$ (8)
+where $\boldsymbol{F}$ is the Fisher information matrix with elements
 $F_{ij}(\boldsymbol{p})=-\frac{1}{2}\frac{\partial^{2}\log L(\boldsymbol{p})}{\partial p_{i}\,\partial p_{j}}.$ (9)
-Intuitively, $\boldsymbol{F}$ measures how sharp the peak is, or how sensitive $L$ is in different directions in parameter space.
-Fisher forecasting is powered by the Cramér-Rao bound $\big\lvert\smash{C_{ij}}\big\rvert\geq\big\lvert\smash{F^{-1}_{ij}}\big\rvert$ for the covariance $C_{ij}$ between parameters $p_{i}$ and $p_{j}$.
-It is an equality for a Gaussian, for which the likelihood (8) is exact.
-Thus, inverting $\boldsymbol{F}(\boldsymbol{p}_{0})$ gives the tightest parameter constraints one can hope for.
-To demonstrate differentiable Fisher forecasting with SymBoltz, we make the best possible CMB (TT) measurement: one of $\bar{C}_{l}$ over the full sky with errors only due to cosmic variance
+Intuitively, $\boldsymbol{F}$ measures how sharp the peak is or how sensitive $L$ is in different directions in parameter space.
+Fisher forecasting is powered by the Cramér–Rao bound $\big\lvert\smash{C_{ij}}\big\rvert\geq\big\lvert\smash{F^{-1}_{ij}}\big\rvert$ for the covariance $C_{ij}$ between parameters $p_{i}$ and $p_{j}$.
+It is an equality for a Gaussian, for which the likelihood expansion (8) is exact.
+Thus, inverting $\boldsymbol{F}(\bar{\boldsymbol{p}})$ gives the tightest possible parameter constraints.
+To demonstrate differentiable Fisher forecasting with SymBoltz, we make the best possible CMB (TT) measurement of $\bar{C}_{l}$ over the full sky with errors only due to cosmic variance
 $\sigma_{l}=\sqrt{\frac{2}{2l+1}}\bar{C}_{l}.$ (10)
@@
 $F_{ij}=\sum_{l}\frac{\partial C_{l}}{\partial p_{i}}\frac{1}{\sigma_{l}^{2}}\frac{\partial C_{l}}{\partial p_{j}}.$ (12)
-The derivatives $\partial C_{l}/\partial p_{i}$ are usually found with error-prone finite differences by carefully tuning the step.
+The derivatives $\partial C_{l}/\partial p_{i}$ are usually found with error-prone finite differences and careful step size tuning.
 SymBoltz avoids this problem altogether with automatic differentiation: {codebox}
-vary = [ M.g.h, M.c.Ω₀, M.b.Ω₀, M.b.YHe, M.ν.Neff, M.h.m_eV, M.X.w₀, M.X.wₐ, M.I.ln_As1e10, M.I.ns,]genprob = parameter_updater(prob, vary)jl, ls = SphericalBesselCache(100:25:1000), 100:1000Cl(p) = spectrum_cmb(:TT, genprob(p), jl, ls)p₀ = map(par -> p[par], vary)dCl_dp = ForwardDiff.jacobian(Cl, p₀)Here vary orders the parameters to differentiate with respect to, and genprob generates a new problem with updated parameters p.
+vary = [ M.g.h, M.c.Ω₀, M.b.Ω₀, M.b.YHe, M.ν.Neff, M.h.m_eV, M.X.w₀, M.X.wₐ, M.I.ln_As1e10, M.I.ns,]genprob = parameter_updater(prob, vary)jl, ls = SphericalBesselCache(100:25:1000), 100:1000Cl(p) = spectrum_cmb(:TT, genprob(p), jl, ls)p₀ = map(par -> p[par], vary)dCl_dp = ForwardDiff.jacobian(Cl, p₀)Here, vary orders the parameters to differentiate with respect to, and genprob generates a new problem with updated parameters p.
 This prepares a vector-to-vector function Cl(p) that is differentiated by ForwardDiff.jacobian to compute $\partial C_{l}/\partial p_{i}$ with dual numbers, as shown in Fig.
@@
 5.
-They are in excellent agreement with finite difference results from CLASS, but these required significant tuning of precision parameters and step sizes.
-Differentiable Fisher forecasts have also been done with an Eisenstein-Hu transfer function fit by Campagne et al. (2023), and through an Einstein-Boltzmann solver by Hahn et al. (2024).
+They agree with finite difference results from CLASS, but these required significant tuning of precision parameters and step sizes.
+Differentiable Fisher forecasts have also been demonstrated with an Eisenstein–Hu transfer function fit by Campagne et al.
+(2023), and through the Einstein–Boltzmann solver DISCO-EB by Hahn et al.
+(2024).
 Figure 4: Normalized derivatives $(\partial C_{l}/\partial p_{i})/C_{l}$ of a CMB $T$ power spectrum with respect to cosmological parameters $p_{i}$ from SymBoltz and automatic differentiation (AD; colored lines) versus CLASS and central finite differences (FD; gray dashes).
@@
 3.5 Differentiable MCMC sampling with supernova data
-Finally, we show that SymBoltz can output differentiable results for gradient-based MCMC samplers, like the No-U-Turn Sampler (NUTS).
-It uses gradients to sample parameter space more efficiently than the ?blind?
-Metropolis-Hastings algorithm.
-We predict apparent magnitudes $m(z)=M+\mu(z)$ of Type Ia supernovae at redshift $z$ and fit to 1048 Pantheon observations $(z_{i},m_{i})$ by Scolnic et al.
+Finally, SymBoltz can output differentiable results for gradient-based MCMC samplers, such as the No-U-Turn Sampler (NUTS).
+It uses gradients to sample parameter space more efficiently than the random walking Metropolis–Hastings algorithm.
+We fit 1048 Pantheon observations $(z_{i},m_{i})$ of Type Ia supernovae by Scolnic et al.
+(2018) to predicted apparent magnitudes $m(z)=M+\mu(z)$ at redshift $z$.
+Their standard absolute magnitude is $M\approx-19.3$, and $\mu(z)=5\lg(d_{L}(z)/10\text{ pc})$ is the distance modulus of the background-derived luminosity distance [^17] This expression is valid for any $\Omega_{k0}$ with complex $\operatorname{sinc}(x)=\sin(x)/x$, but can be split into branches for positive, negative, and zero $\Omega_{k0}$.
+$d_{L}(z)=\frac{c}{H_{0}}\frac{\chi(z)}{a(z)}\operatorname{sinc}\Big(H_{0}\sqrt{-\Omega_{k0}}\,\chi(z)\Big).$ (13)
+We define the likelihood $L$ by a multivariate normal distribution of the 1048 supernovae with covariance matrix $C$ given by Scolnic et al.
 (2018).
-Their standard absolute magnitude is $M\approx-19.3$, and $\mu(z)=5\lg(d_{L}(z)/10\text{ pc})$ is the distance modulus from the background-derived luminosity distance [^14] This expression is valid for any $\Omega_{k0}$ with complex $\operatorname{sinc}(x)=\sin(x)/x$, but can be split into branches for positive, negative and zero $\Omega_{k0}$.
-$d_{L}(z)=\frac{c}{H_{0}}\frac{\chi(z)}{a(z)}\operatorname{sinc}\Big(H_{0}\sqrt{-\Omega_{k0}}\,\chi(z)\Big).$ (13)
-We define the likelihood $L$ by a multivariate normal distribution of the 1048 supernovae with (constant) covariance matrix $C$ given in detail by Scolnic et al.
-(2018).
-To compute $L$, we interface SymBoltz with the probabilistic programming framework Turing.jl (Fjelde et al., 2025), which allows a high-level description of the probabilistic model equivalent to the log-likelihood
+To compute $L$, we use the probabilistic programming framework Turing.jl [^18] https://github.com/TuringLang/Turing.jl by Fjelde et al.
+(2025) and specify a probabilistic model equivalent to the log-likelihood
 $\log L=-\frac{1}{2}\sum_{i,j}C^{-1}_{ij}\big(m(z_{i})-m_{i}\big)\big(m(z_{j})-m_{j}\big).$ (14)
-We use the NUTS sampler in Turing.
+We used the NUTS sampler in Turing.
 The code for this example can be found in the paper’s notebook.
-For the flat $w_{0}\text{CDM}$ model ($\Omega_{k0}=0$, $w_{a}=0$, fixed $\Omega_{r0}$) it gives the constraints in fig.
+For the flat $w_{0}\text{CDM}$ model ($\Omega_{k0}=0$, $w_{a}=0$, fixed $\Omega_{r0}$), it gives the constraints in Fig.
 6.
-The differentiable computation is not yet fast enough for MCMCs with perturbation-derived spectra.
-The results here are not groundbreaking, but show that the differentiable pipeline is functional.
-This is a first step to differentiating the solver quickly, and we aim to accelerate this computation with future work.
-Differentiable Boltzmann codes have yet to demonstrate this ability, but more advanced MCMCs with a differentiable background have been done by Campagne et al. (2023), for example.
-Figure 6: Parameter inference on 1048 Type Ia supernovae from Pantheon data, using 5000 MCMC samples with the gradient-based NUTS sampler in Turing.jl and differentiable predictions from SymBoltz.
+Differentiable computations are not yet fast enough for MCMCs with perturbation-derived spectra.
+These preliminary results show that the differentiable pipeline works, and we hope to speed it up with future work.
+Differentiable Boltzmann codes have yet to show use with gradient-based MCMC samplers, but Campagne et al.
+(2023) has done more advanced MCMCs with a differentiable background, for example.
+Figure 6: Parameter inference on 1048 Type Ia supernovae from Pantheon data, using 5000 MCMC samples with the gradient-based NUTS sampler in Turing and differentiable predictions from SymBoltz.
 4 Performance comparison to CLASS
-Figure 7: Work vs.
-precision for computing the matter power spectrum $P(k)$ with SymBoltz and CLASS.
+Figure 7: Work versus precision for computing the matter power spectrum $P(k)$ with SymBoltz and CLASS.
 SymBoltz is approximation-free and run with several implicit ODE solvers.
-CLASS is run with approximations on/off, its implicit/explicit evolvers ndf15/rk and precision settings in appendix B, but tight coupling is always approximated.
+CLASS is run with approximations on/off, its implicit/explicit evolvers ndf15/rk, and precision settings in Appendix B, but tight coupling is always approximated.
 For each setup, $P(k)$ is computed by integrating perturbation $k$-modes with ODE solver tolerances $10^{-2}$-$10^{-9}$ that controls the adaptive stepping.
-For each tolerance, we record the best of 3 runtimes and the $L_{2}$-compressed error $\smash{\big(\sum_{i=1}^{N}\!\big(P(k_{i})/\bar{P}(k_{i})-1\big)^{2}/N\big)^{1/2}}$ relative to high-precision spectra $\bar{P}(k)$ from both codes with approximations off, tolerance $10^{-10}$ and FBDF/ndf15.
-Both codes solve a $w_{0}w_{a}\text{CDM}$ model with equal parameters, $N=114$ equal wavenumbers $k_{i}$, multipole cutoff $l_{\text{max}}=16$ for all species, default sampling of massive neutrino momenta and parallelize over $k$.
-Times are from a Linux laptop with Intel i7-12800H CPU.
+For each tolerance, we record the best of 3 runtimes and the $L_{2}$-compressed error $\smash{\big(\sum_{i=1}^{N}\!\big(P(k_{i})/\bar{P}(k_{i})-1\big)^{2}/N\big)^{1/2}}$ relative to high-precision spectra $\bar{P}(k)$ from both codes with approximations off, tolerance $10^{-10}$, and FBDF/ndf15.
+Both codes solve a $w_{0}w_{a}\text{CDM}$ model with equal parameters, $N=114$ equal wavenumbers $k_{i}$, multipole cutoff $l_{\text{max}}=16$ for all species, default sampling of massive neutrino momenta, and parallelization over $k$.
+Times are from a Linux laptop with an Intel i7-12800H CPU.
 Figure 7 compares performance versus precision of the matter power spectrum $P(k)$ computed with SymBoltz and CLASS.
-This computation solves the background, thermodynamics and perturbations for several $k$ and reads off $P(k)$ at the final time.
-The work is dominated by the perturbations, so it is a relevant test of SymBoltz’ approximation-free treatment of perturbations.
-The comparison shows how approximations and the ODE solver affect integration of the Einstein-Boltzmann equations.
+This computation solves the background, thermodynamics, and perturbations for several $k$ and reads off $P(k)$ at the final time.
+It is a relevant test of SymBoltz’ approximation-free treatment of the perturbations, as they dominate the computational work.
+The comparison shows how approximations and the ODE solver affect the integration of the equations.
 Without approximations, SymBoltz with Rodas5P is more than $10\times$ faster than CLASS with its implicit ndf15 evolver at comparable precision.
-When CLASS uses approximations and explicit evolver, SymBoltz is as fast while remaining approximation-free.
-To make this comparison as fair as possible, we configured both codes to integrate identical $k$-modes (decided by CLASS) without $k$-interpolation and use the same multipole cutoff $l_{\text{max}}$ for all species.
+When CLASS uses approximations and its explicit rk evolver, SymBoltz is as fast while remaining approximation-free.
+To make this comparison as fair as possible, we configured both codes to integrate identical $k$-modes (decided by CLASS) without $k$-interpolation and use equal multipole cutoffs $l_{\text{max}}$ for all species.
 However, the codes sample massive neutrino momenta $q$ differently.
-This greatly impacts the number of perturbation equations (see section A.7).
-SymBoltz’ 4 default momenta gives 127 perturbation equations in total.
+This greatly impacts the number of perturbation equations (see Appendix A.7).
+SymBoltz’ 4 default momenta give 127 perturbation equations in total.
 CLASS’ default $\texttt{tol\_ncdm\_newtonian}=10^{-5}$ gives 11 momenta and 245 equations (before approximations).
-Increasing this tolerance to $10^{-3}$ gives 5 momentum bins, 143 equations and is around $245/143\approx 1.7$ times faster, but introduces errors beyond $1\%$ in $P(k)$.
-We therefore leave both codes with default $q$-sampling, but note that CLASS could be up to $2\times$ faster at unchanged precision with sparser sampling.
-This would shift its curves down by $\lg 2\approx 0.3$.
+Increasing this tolerance to $10^{-3}$ gives 5 momenta and 143 equations.
+This runs around $245/143\approx 1.7$ times faster, but introduces errors beyond $1\%$ in $P(k)$.
+We therefore leave both codes with default $q$-sampling, but note that sparser sampling could make CLASS up to $2\times$ faster at unchanged precision and shift its curves down by $\lg 2\approx 0.3$.
 Howlett et al.
-(2012) (in appendix A) also find a precise 3-4 point scheme in CAMB, and Lee et al.
-(2025) speed up CLASS using integral equations for massive neutrinos in the fork CLASSIER.
-Despite lack of approximation schemes, SymBoltz is fast due to high-order implicit solvers like Rodas5P that take accurate and long steps, fast functions for $\boldsymbol{f}$ and the analytically generated $\boldsymbol{J}$, sparse matrix methods with a constant predetermined sparsity structure and efficient sampling of massive neutrino momenta.
-SymBoltz is currently slower than CAMB and CLASS for computing $C_{l}$.
-This is subject to more optimizations than $P(k)$, such as sampling of the source function $S(\tau,k)$, interpolation in $\tau$, $k$, and $l$, evaluation of spherical Bessel functions $j_{l}(x)$ and optimized line-of-sight quadrature.
-These are very well polished by CAMB and CLASS, and unmatched by SymBoltz at this time.
-A detailed performance analysis for $C_{l}$ is therefore not meaningful now (but see table 1).
-With more work, the $C_{l}$ optimizations can be added to SymBoltz.
-The purpose of the comparison made here is to show the viability of approximation-free perturbations, which was a major milestone in designing SymBoltz.
+(2012, in Appendix A) also found a precise three- to four-point scheme in CAMB, and Lee et al.
+(2025) sped up CLASS using integral equations for massive neutrinos.
+Despite the lack of approximation schemes, SymBoltz is fast thanks to high-order implicit solvers, fast generated functions for $\boldsymbol{f}$ and the analytical $\boldsymbol{J}$, sparse matrix methods with a constant precomputed sparsity pattern, and efficient sampling of massive neutrino momenta.
+However, SymBoltz currently computes $C_{l}$ slower than CAMB and CLASS.
+This computation is subject to more optimizations than $P(k)$, such as sampling of source functions $S(\tau,k)$, interpolation in $\tau$, $k$, and $l$, evaluation of spherical Bessel functions $j_{l}(x)$, and fast line-of-sight quadrature.
+These aspects are not yet as polished in SymBoltz as in CAMB and CLASS, so a detailed performance analysis for $C_{l}$ (beyond Table 1) is not meaningful now.
+Further work can add these optimizations to SymBoltz.
+The purpose of the comparison made here is to show the viability of approximation-free perturbations, which marks a milestone in the design of SymBoltz.
 5 Future work
 We hope SymBoltz continues to grow and that others can build on it. Its symbolic capabilities can be extended with utilities for:
-  • separation into finer computational stages (see section 2.1.3);
-  • gauge transformation of equations (e.g. Newtonian $\leftrightarrow$ synchronous gauge), as already possible in CAMB [^15] https://camb.readthedocs.io/en/latest/symbolic.html;
-  • perturbation theory utilities for deriving initial conditions and approximation schemes for new models;
-  • transformation of differential equations with respect to conformal time $\tau$ to another (one-to-one) independent variable (e.g. to get observables as direct functions of redshift $z$);
+  • separation into finer computational stages (see Sect. 2.1.3);
+  • gauge transformation of equations (e.g., Newtonian $\leftrightarrow$ synchronous gauge), as is already possible in CAMB; [^19] https://camb.readthedocs.io/en/latest/symbolic.html
+  • perturbation theory utilities to derive initial conditions and approximation schemes for new models;
+  • transformation of differential equations with respect to conformal time $\tau$ to another (one-to-one) independent variable (e.g., to get observables as direct functions of redshift $z$);
   • validation of dimensions in symbolic equations to catch modeling mistakes by tagging variables with physical units;
-  • symbolic tensor manipulation to aid work with modified gravity theories (as a standalone package).
-These tasks traditionally involve manual error-prone calculations, so such utilities could ease exploration of extended models.
-We particularly aim to improve performance for differentiable runs to the level of standard runs.
+  • symbolic tensor algebra for work with modified gravity.
+These tasks traditionally involve manual error-prone calculations, so such utilities could aid in the exploration of extended models.
+In particular, we also aim to improve performance for differentiable runs to the level of standard runs.
 An attractive milestone is fast reverse-mode gradients of scalar loss functions for gradient-based MCMCs.
 We anticipate the symbolic structure and Jacobian generation to be useful for this, as writing $\boldsymbol{u}(\tau,\boldsymbol{p})$ in the ODE (3) and differentiating it for $\boldsymbol{v}_{i}=\partial\boldsymbol{u}/\partial p_{i}$ gives $\mathrm{d}\boldsymbol{v}_{i}/\mathrm{d}\tau=\boldsymbol{J}\boldsymbol{v}_{i}+\partial\boldsymbol{f}/\partial p_{i}$, which could be constructed analytically.
-It is left for future work to generalize the code from scalar to vector and tensor perturbations, from flat to curved spacetimes, and from adiabatic to other initial conditions.
-Some additional models like quintessence dark energy and Brans-Dicke gravity are in the works, and we invite others to add extended models.
+Future work could also generalize the code from scalar to vector and tensor perturbations, from flat to curved spacetimes, and from adiabatic to other initial conditions.
+Some additional models for quintessence dark energy and Brans–Dicke gravity are in the works, and we invite others to add extended models.
 We hope SymBoltz can foster growth of modular interoperating packages for cosmology in Julia.
-For example, applications like higher-order perturbation theory (e.g.
-CLASS-PT by Chudaykin et al.
-(2020) and PyBird by D’Amico et al.
-(2021)), nonlinear boosting (e.g.
-HALOFIT by Takahashi et al.
-(2012) and HMCODE by Mead et al.
-(2021)), CMB lensing (e.g.
-CMBLensing.jl by Millea et al.
-(2020)) and nonlinear $N$-body simulations all rely on output from Boltzmann solvers.
-Supporting automatic differentiation across language barriers is hard, and an ecosystem of such packages in one language with support for automatic differentiation, like Julia or Python with JAX, could create a modular and differentiable cosmological modeling toolbox.
-A recent example of this is DISCO-DJ’s coupling of differentiable Boltzmann and $N$-body codes for end-to-end modeling from linear perturbations to non-linear structure formation (Hahn et al., 2024; List et al., 2025).
+For example, applications such as higher order perturbation theory (e.g., CLASS-PT by Chudaykin et al.
+2020 and PyBird by D’Amico et al.
+2021), nonlinear boosting (e.g., HALOFIT by Takahashi et al.
+2012 and HMCODE by Mead et al.
+2021), CMB lensing (e.g., CMBLensing.jl by Millea et al.
+2020), and nonlinear $N$-body simulations all rely on output from Boltzmann solvers.
+Supporting automatic differentiation across programming languages is hard, so an ecosystem of such packages in one language with support for automatic differentiation, such as Julia or Python with JAX, could create a modular and differentiable cosmological modeling toolbox.
+A recent example of this is DISCO-DJ’s coupling of differentiable Boltzmann and $N$-body codes for end-to-end modeling from linear perturbations to nonlinear structure formation (Hahn et al., 2024; List et al., 2025).
 6 Conclusion
-SymBoltz is a fresh Julia package for solving the linear Einstein-Boltzmann equations.
+SymBoltz is a fresh Julia package for solving the linear Einstein–Boltzmann equations.
 It features a symbolic-numeric interface where models are built from symbolic equations and compiled to efficient numerical code.
-This lets modelers prototype new models with few lines code at a high level.
+This lets users prototype new models with few lines of high-level code.
 It relaxes all approximation schemes found in older codes by solving a single set of stiff equations at all times with implicit integrators.
-This simplifies the equations and remains performant due to high-order implicit solvers with fast generated code for ODEs and analytical and sparse Jacobians.
-It is differentiable, so one can get accurate derivatives of any output quantity with respect to any input parameter.
-Output power spectra agree with CLASS to $0.1\%$ with standard precision and can be used fit to current data.
+This simplifies the equations and remains fast due to high-order implicit solvers that integrate efficient generated code for ODEs with analytical and sparse Jacobians.
+It is differentiable, so we can get accurate derivatives of any output quantity with respect to any input parameter.
+Output power spectra agree with CLASS to $0.1\%$ with standard precision, so SymBoltz can be used to fit current data.
 Symbolic-numericinterfaceApproximation-freenessDifferentiabilityDetect sparsity of $\boldsymbol{J}$Calculate $\boldsymbol{J}$ robustlyEasy to specify model
-Figure 8: A symbolic-numeric interface, approximation-freeness and differentiability are three main features of SymBoltz that form a synergy.
-Differentiability computes the ODE Jacobian $\boldsymbol{J}$ accurately and efficiently; which is used by implicit solvers to integrate stiff ODEs without approximations; which makes it easier to write down equations in simple symbolic form; which is used to find the sparsity pattern of $\boldsymbol{J}$; which speeds up $\boldsymbol{J}$ and the implicit solver.
-This is a self-reinforcing design.
-Other recent codes like PyCosmo, Bolt.jl and DISCO-EB also incorporate some of these features.
+Figure 8: A symbolic-numeric interface, approximation-freeness, and differentiability are three main features of SymBoltz that form a synergy.
+Differentiability computes the ODE Jacobian $\boldsymbol{J}$ accurately and efficiently; which implicit solvers use to integrate stiff ODEs without approximations; which makes it easy to write equations in simple symbolic form; which provide the exact sparsity pattern of $\boldsymbol{J}$; in turn speeding up $\boldsymbol{J}$ and the solver.
+This loop creates a self-reinforcing design.
+Other recent codes, such as PyCosmo, Bolt, and DISCO-EB, also incorporate some of these features.
 SymBoltz combines them and reinforces its design with a synergy between them, as explained in Fig.
 8.
-Together, these new codes offer a simpler alternative in high-level languages that complements traditional codes like CAMB and CLASS that are historically highly tuned around approximations in low-level languages.
+Together, these new codes offer a simpler alternative in high-level languages that complements traditional codes such as CAMB and CLASS, which are historically highly tuned around approximations in low-level languages.
 However, new codes need more work to reach the level of features and maturity of CAMB and CLASS that have grown over many years.
-SymBoltz version 0.14.1 includes the flat Newtonian gauge metric, General Relativity, cold dark matter, photons, baryons and RECFAST recombination, massless and massive neutrinos, the cosmological constant and $w_{0}w_{a}$ dark energy with scalar perturbations, distances and matter and CMB spectra.
-Forward-mode automatic differentiation works for single-run purposes like Fisher forecasting.
+SymBoltz version 1.0.0 includes the flat Newtonian gauge metric, general relativity, cold dark matter, photons, baryons and RECFAST recombination, massless and massive neutrinos, the cosmological constant, and $w_{0}w_{a}$ dark energy with scalar perturbations, distances, and matter and CMB spectra.
+Forward-mode automatic differentiation works for single-run purposes such as Fisher forecasting.
 Fast reverse-mode scalar loss gradients, particularly through the perturbations, is a future milestone that could challenge emulators in gradient-based MCMC sampling.
-No differentiable Boltzmann code is yet capable of this.
-SymBoltz is easy to install from https://github.com/hersle/SymBoltz.jl.
-Documentation [^16] https://hersle.github.io/SymBoltz.jl/ is available, and the code is tested with continuous integration to remain correct and up-to-date (see appendix C).
-All questions and contributions are welcome in the repository!
+No differentiable Boltzmann code is capable of this yet.
+SymBoltz is publicly available [^20] https://github.com/hersle/SymBoltz.jl and easy to install.
+Documentation is also available, [^21] https://hersle.github.io/SymBoltz.jl/ and the code is tested with continuous integration to remain up-to-date (see Sect.
+C).
+All questions and contributions are welcome in the repository.
 We hope SymBoltz proves useful and that it offers valuable ideas on the Boltzmann solver market.
-Acknowledgments
-I thank the Julia community for creating an open ecosystem of scientific packages underlying SymBoltz.
-I thank Aayush Sabharwal and Christopher Rackauckas for developing ModelingToolkit.jl and OrdinaryDiffEq.jl and answering questions.
+Acknowledgements.
+I thank the Julia community for creating an extensive open ecosystem of scientific packages underlying SymBoltz.
+I thank Aayush Sabharwal and Christopher Rackauckas for developing ModelingToolkit and OrdinaryDiffEq and answering questions.
 I thank Hans A.
 Winther for helpful suggestions and feedback on the code and drafts of this paper.
-I thank Julien Lesgourgues, Thomas Tram and others for developing and documenting CLASS.
+I thank Julien Lesgourgues, Thomas Tram, and others for developing and thoroughly documenting CLASS.
+I thank the anonymous referee for detailed and constructive comments that helped improve the quality of this paper.
+This research was supported by the Research Council of Norway under project number 325113.
 Appendix A List of equations and practical implementation details
@@
 We hope this can be a useful reference for others.
-The list is structured like SymBoltz with one component per subsection.
-Variables are in units where $G=c=H_{0}=1$?
-and temperatures are in $\mathrm{K}$, unless otherwise is stated.
-These units are chosen because $G$, $c$ and $H_{0}$ can be divided out from the Einstein equations as natural units, but this requires some conversion in the recombination equations that depend explicitly on $H_{0}$.
-In other words, times are in units of $1/H_{0}$, distances in $c/H_{0}$ and masses in $c^{3}/H_{0}G$.
-When $G$, $c$ or and $H_{0}$ appear explicitly in the equations, they are in SI units and used only to convert from SI units into the dimensionless units.
-The equations closely follow the conventions in the seminal paper by Ma and Bertschinger (1995) and very closely matches the source code of SymBoltz with Unicode characters.
+The list mirrors the internal structure of SymBoltz with one component per subsection.
+Unless otherwise stated, temperatures are in $\mathrm{K}$ and other variables are in units where $G=c=H_{0}=1$.?
+These units are chosen because $G$, $c$, and $H_{0}$ can be divided out from the Einstein equations as natural units (but this requires some conversion in the recombination equations that depend explicitly on $H_{0}$).
+In other words, times are in units of $1/H_{0}$, distances in $c/H_{0}$, and masses in $c^{3}/H_{0}G$.
+When $G$, $c$, and $H_{0}$ appear explicitly in the equations, they are in SI units and used only to convert from SI units into the dimensionless units.
+The equations closely follow the conventions in the seminal paper by Ma and Bertschinger (1995) and very closely match the source code of SymBoltz with Unicode characters.
 It is far outside the scope of this paper to derive the equations and explain the meaning of every variable.
 The independent variable is conformal time $\tau$, and all derivatives $^{\prime}=\mathrm{d}/\mathrm{d}\tau$ are with respect to it (in units of $1/H_{0}$).
-This is perhaps the most common parametrization in the literature, and is natural because most equations are autonomous with respect to $\tau$ (i.e.
-$f(\tau,\boldsymbol{u})\rightarrow f(\boldsymbol{u})$, except some multipole truncation schemes that use $1/(k\tau)$, but these are unphysical).
-By default, integration starts from the early time $\tau=\tau_{i}=10^{-6}$, and is terminated when the scale factor crosses $a=1$, and the corresponding time $\tau=\tau_{0}$ is labeled today.
+This is perhaps the most common parameterization in the literature, and it is natural because most equations are autonomous with respect to $\tau$ (i.e., $\boldsymbol{f}(\boldsymbol{u},\tau)\rightarrow\boldsymbol{f}(\boldsymbol{u})$, except some multipole truncation schemes that have factors of $1/(k\tau)$, but these are not of physical origin).
+By default, integration starts from the early time $\tau=\tau_{i}=10^{-6}$ and terminates when the scale factor crosses $a=1$ at the time $\tau=\tau_{0}$ today.
 A.1 Metric and spacetime $(g)$
@@
 $g_{0i}=g_{i0}=-a^{2}(1+2\Psi)\delta_{0i},\qquad g_{ij}=a^{2}(1-2\Phi)\delta_{ij},\qquad z=\frac{1}{a}-1,\qquad\mathscr{H}=\frac{a^{\prime}}{a},\qquad H=\frac{\mathscr{H}}{a},\qquad\chi=\tau_{0}-\tau.$
-Here $\mathscr{H}$ and $H$ are the conformal and cosmic Hubble factors (in units where $\mathscr{H}_{0}=H_{0}=1$).
+Here, $\mathscr{H}$ and $H$ are the conformal and cosmic Hubble factors (in units where $\mathscr{H}_{0}=H_{0}=1$).
 The scale factor $a$ is related to redshift $z$, and $\chi$ is the lookback time from today that appears in some integral solutions.
@@
 The gravitational theory of the $\mathrm{\Lambda CDM}$ model is general relativity governed by the Einstein field equations $G_{\mu\nu}=8\pi T_{\mu\nu}$.
-By default, SymBoltz solves for the metric variables $a$, $\Phi$ and $\Psi$ with $(\mu,\nu)=(0,0)$ in the background (1st Friedmann equation) and $(\mu,\nu)=\{(0,0),(i,j)\}$ in the perturbations:
+By default, SymBoltz solves for the metric variables $a$, $\Phi$, and $\Psi$ with $(\mu,\nu)=(0,0)$ in the background (first Friedmann equation), and $(\mu,\nu)=\{(0,0),(i,j)\}$ in the perturbations:
 $a^{\prime}=\sqrt{\frac{8\pi}{3}\rho}\,a^{2},\qquad\Phi^{\prime}=-\mathscr{H}\Psi-\frac{k^{2}}{3\mathscr{H}}\Phi-\frac{4\pi}{3}\frac{a^{2}}{\mathscr{H}}{\delta\rho},\qquad\Psi=-\Phi-12\pi\bigg(\frac{a}{k}\bigg)^{2}\Pi.$
-Note that it is also possible to evolve other redundant combinations of the Einstein equations (such as the acceleration equation).
-The equations are coupled to total densities $\rho$, $\delta\rho$, pressures $P$ and anisotropic stresses $\Pi$ for an arbitrary set of species $s$ that are present in the cosmological model:
+We note that it is also possible to evolve other redundant combinations of the Einstein equations, such as the acceleration equation.
+The equations are coupled to total densities $\rho$ and $\delta\rho$, pressures $P$ and $\delta P$, and anisotropic stress $\Pi$ for whichever set of species $s$ that are present in the cosmological model:
 $\rho=\sum_{s}\rho_{s},\qquad P=\sum_{s}P_{s},\qquad\delta\rho=\sum_{s}\delta\rho_{s}=\sum_{s}\delta_{s}\rho_{s},\qquad\delta P=\sum_{s}\delta P_{s}=\sum_{s}\delta_{s}\rho_{s}c_{s,s}^{2},\qquad\Pi=\sum_{s}\Pi_{s}=\sum_{s}(\rho_{s}+P_{s})\sigma_{s}.$
@@
 It only reacts to total stress-energy components.
-All species must therefore define $\rho$, $P$, $\delta\rho$, $\delta P$ and $\Pi$ explicitly, even if zero.
+All species must therefore define $\rho$, $P$, $\delta\rho$, $\delta P$, and $\Pi$ explicitly, even if they are zero.
 This requirement is somewhat pedantic, but helps isolate components from each other for greater reuse when composing models.
 The scale factor $a$ is initialized as the nonlinear solution of the Friedmann equation constrained to $\mathscr{H}=1/\tau$ (motivated by its radiation-dominated solution $a=\sqrt{\Omega_{r0}}\,\tau$).
-The constraint potential is initialized to $\Psi=20C/(15+4f_{\nu})$ with the (arbitrary) integration constant $C=1/2$ and initial energy density fraction $f_{\nu}=(\rho_{\nu}+\rho_{h})/(\rho_{\nu}+\rho_{h}+\rho_{\gamma})$ of all (massless and massive) neutrino species relative to all species that are radiative at early times.
-The evolved potential $\Phi$ is initialized accordingly from the constraint equation (the found solution is close to $\Phi=(1+2f_{\nu}/5)\Psi$, but providing both $\Phi$ and $\Psi$ explicitly leads to overdetermined initialization equations due to the constraint equation).
-The next sections present the $\mathrm{\Lambda CDM}$ section of the ?library of species? that are available in SymBoltz.
+The constraint potential is initialized to $\Psi=20C/(15+4f_{\nu})$ with the (arbitrary) integration constant $C=1/2$ and initial energy density fraction $f_{\nu}=(\rho_{\nu}+\rho_{h})/(\rho_{\nu}+\rho_{h}+\rho_{\gamma})$ of all (massless and massive) neutrino species relative to all species that are radiation-like at early times.
+The evolved potential $\Phi$ is initialized accordingly from the constraint equation (the result is close to $\Phi=(1+2f_{\nu}/5)\Psi$, but providing both $\Phi$ and $\Psi$ explicitly leads to overdetermined initialization equations that violate the constraint equation for $\Psi$).
+The next sections present the $\mathrm{\Lambda CDM}$ part of the ?library of species?
+that are available in SymBoltz.
 A.3 Cold dark matter $(c)$
-Cold dark matter is a non-relativistic and non-interacting species that follows very simple equations:
+Cold dark matter is a nonrelativistic and noninteracting species that follows very simple equations:
 $w=0,\qquad{c_{s}^{2}}=w,\qquad P=0,\qquad\rho=\frac{\rho_{0}}{a^{3}},\qquad\delta^{\prime}=-\theta+3\Phi^{\prime},\qquad\theta^{\prime}=-\mathscr{H}\theta+k^{2}\Psi,\qquad u=\frac{\theta}{k},\qquad\sigma=0.$
-Initial conditions are adiabatic with $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$. The species is parametrized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today.
+Initial conditions are adiabatic with $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$. The species is parameterized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today.
 A.4 Baryons $(b)$
-Baryons are also non-relativistic, but interact with photons through Compton scattering and are subject to recombination physics.
+Baryons are also nonrelativistic, but interact with photons through Compton scattering and are subject to recombination physics.
 This significantly complicates their behavior.
-SymBoltz currently implements equations from RECFAST [^17] https://www.astro.ubc.ca/people/scott/recfast.html version 1.5.2 (Seager et al., 1999; Wong et al., 2008; Scott and Moss, 2009):
+SymBoltz currently implements equations from RECFAST [^22] https://www.astro.ubc.ca/people/scott/recfast.html version 1.5.2 (Seager et al.
+1999; Wong et al.
+2008; Scott and Moss 2009):
 $w=0,\qquad P=0,\qquad\rho=\frac{\rho_{0}}{a^{3}},\qquad{f_{\mathrm{He}}}=\frac{{Y_{\mathrm{He}}}}{\frac{m_{\mathrm{He}}}{m_{\mathrm{H}}}\big(1-{Y_{\mathrm{He}}}\big)},\qquad n_{\mathrm{H}}=\frac{(1-Y_{\mathrm{He}})\rho}{m_{\mathrm{H}}},\qquad n_{\mathrm{He}}=f_{\mathrm{He}}n_{\mathrm{H}},\qquad{c_{s}^{2}}=\frac{k_{B}}{\mu c^{2}}\bigg({T_{b}}-\frac{T_{b}^{\prime}}{3\mathscr{H}}\bigg),$
@@
 Helium rates and coefficients are even more complicated. First, Helium includes contributions from singlet states ($\mathrm{He}_{1}$}):
-$\alpha_{\mathrm{He}_{1}}=\frac{q_{1}}{\sqrt{\frac{T_{b}}{T_{2}}}\Big(1+\sqrt{\frac{T_{b}}{T_{2}}}\Big)^{1-p_{1}}\Big(1+\sqrt{\frac{T_{b}}{T_{1}}}\Big)^{1+p_{1}}},\qquad\beta_{\mathrm{He}_{1}}=4\frac{\alpha_{\mathrm{He}_{1}}}{\lambda_{\mathrm{e}}^{3}}\exp\big({-βE_{\mathrm{He}_{1}}^{\infty,2s}}\big),\qquad K_{\mathrm{He}_{1}}=\frac{1}{K_{\mathrm{He}_{1}^{0}}^{-1}+K_{\mathrm{He}_{1}^{1}}^{-1}+K_{\mathrm{He}_{1}^{2}}^{-1}},$
+$\alpha_{\mathrm{He}_{1}}=\frac{q_{1}}{\sqrt{\frac{T_{b}}{T_{2}}}\Big(1+\sqrt{\frac{T_{b}}{T_{2}}}\Big)^{1-p_{1}}\Big(1+\sqrt{\frac{T_{b}}{T_{1}}}\Big)^{1+p_{1}}},\qquad\beta_{\mathrm{He}_{1}}=4\frac{\alpha_{\mathrm{He}_{1}}}{\lambda_{\mathrm{e}}^{3}}\exp\big({-\beta E_{\mathrm{He}_{1}}^{\infty,2s}}\big),\qquad K_{\mathrm{He}_{1}}=\frac{1}{K_{\mathrm{He}_{1}^{0}}^{-1}+K_{\mathrm{He}_{1}^{1}}^{-1}+K_{\mathrm{He}_{1}^{2}}^{-1}},$
 $K_{\mathrm{He}_{1}^{0}}^{-1}=\frac{8\pi H}{\big(\lambda_{\mathrm{He}_{1}}^{2p,1s}\big)^{3}},\qquad K_{\mathrm{He}_{1}^{1}}^{-1}=-\exp(-\tau_{\mathrm{He}_{1}})K_{\mathrm{He}_{1}^{0}}^{-1},\qquad K_{\mathrm{He}_{1}^{2}}^{-1}=\frac{A_{2p_{1}}}{3\big(1+0.36\,\gamma_{2p_{1}}^{0.86}\big)n_{\mathrm{He}}(1-X_{\mathrm{He}}^{+})},\qquad\tau_{\mathrm{He}_{1}}=\frac{3A_{2p_{1}}n_{\mathrm{He}}(1-X_{\mathrm{He}}^{+})}{K_{\mathrm{He}_{1}^{0}}^{-1}},$
@@
 $\gamma_{2p_{3}}=\frac{3A_{2p_{3}}f_{\mathrm{He}}c^{2}(1-X_{\mathrm{He}}^{+})}{8\pi\sigma_{{\mathrm{He}_{3}}}\sqrt{\frac{2\pi}{\beta m_{\mathrm{He}}c^{2}}}\big(f_{\mathrm{He}_{3}}^{2p,1s}\big)^{3}(1-X_{\mathrm{H}}^{+})},\qquad C_{\mathrm{He}_{3}}=\frac{A_{2p_{3}}\bigg(\frac{1-\exp({-\tau_{\mathrm{He}_{3}}})}{\tau_{\mathrm{He}_{3}}}+\frac{1}{3\big(1+0.66\,\gamma_{2p_{3}}^{0.9}\big)}\bigg)\exp\big({-\beta E_{\mathrm{He}_{3}}^{2p,2s}}\big)}{A_{2p_{3}}\bigg(\frac{1-\exp({-\tau_{\mathrm{He}_{3}}})}{\tau_{\mathrm{He}_{3}}}+\frac{1}{3\big(1+0.66\,\gamma_{2p_{3}}^{0.9}\big)}\bigg)\exp\big({-\beta E_{\mathrm{He}_{3}}^{2p,2s}}\big)+\beta_{\mathrm{He}_{3}}}.$
-Every variable that does not occur on the left side of an equation is either a constant or a parameter.
-This includes $Y_{\mathrm{He}}$, fudge factors and wavenumbers, frequencies and energies for atomic transitions.
-Some important variables are the baryon temperature $T_{b}$, photon temperature $T_{\gamma}$, mean molecular weight $\mu$, baryon sound speed $c_{s}^{2}$, optical depth $\kappa$, visibility function $v$ and the free electron fraction $X_{\mathrm{e}}$ (conventionally relative to Hydrogen, so $X_{\mathrm{e}}>1$ in presence of Helium).
-Please consult the code and RECFAST references cited above for more details.
+Every variable that does not occur on the left side of an equation is either a constant or a parameter given in the code.
+This includes $Y_{\mathrm{He}}$, fudge factors, and wavenumbers, frequencies, and energies for atomic transitions.
+Some important variables defined above are the baryon temperature $T_{b}$, photon temperature $T_{\gamma}$, mean molecular weight $\mu$, baryon sound speed $c_{s}^{2}$, optical depth $\kappa$, visibility function $v$, and the free electron fraction $X_{\mathrm{e}}$ (conventionally relative to Hydrogen, so $X_{\mathrm{e}}>1$ in presence of Helium).
+We refer to the code and RECFAST references cited above for more details.
 Unlike other RECFAST implementations, SymBoltz does not approximate the stiff Peebles equations at early times by Saha approximations (although $X_{\mathrm{He}}^{++}$ is given by a Saha equation at all times).
@@
 Atomic calculations are done in SI units and converted to SymBoltz’ dimensionless units by factors of $H_{0}$ in SI units.
-The differential equation for $T_{b}^{\prime}$ is very stiff and sensitive to $T_{b}-T_{\gamma}$, but $T_{b}\approx T_{\gamma}$ in the early universe, so we rewrite it to a more stable differential equation for $\Delta T^{\prime}=T_{b}^{\prime}-T_{\gamma}^{\prime}$ instead, initialize $\Delta T=0$ and observe $T_{b}=\Delta T+T_{\gamma}$.
-The optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\tau^{\prime})\mathrm{d}\tau^{\prime}$ is really a line-of-sight integral into the past, but is integrated together with the background ODEs by initializing $\kappa(\tau_{i})=0$ to an arbitrary value, integrating the differential equation for $\kappa^{\prime}$ and subtracting the final value of $\kappa(\tau_{0})$ (i.e.
-$\int_{\tau_{0}}^{\tau}=\int_{\tau_{0}}^{\tau_{i}}+\int_{\tau_{i}}^{\tau}=\int_{\tau_{i}}^{\tau}-\int_{\tau_{i}}^{\tau_{0}}$).
+The differential equation for $T_{b}^{\prime}$ is very stiff and sensitive to $T_{b}-T_{\gamma}$, but $T_{b}\approx T_{\gamma}$ in the early universe, so we rewrite it to a more stable differential equation for $\Delta T^{\prime}=T_{b}^{\prime}-T_{\gamma}^{\prime}$ instead, initialize $\Delta T=0$, and observe $T_{b}=\Delta T+T_{\gamma}$.
+The optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\tau^{\prime})\mathrm{d}\tau^{\prime}$ is really a line-of-sight integral into the past, but we solve it together with the background ODEs by initializing $\kappa(\tau_{i})=0$ to an arbitrary value, integrating the differential equation for $\kappa^{\prime}$, and subtracting the final value of $\kappa(\tau_{0})$ (i.e., $\int_{\tau_{0}}^{\tau}=\int_{\tau_{0}}^{\tau_{i}}+\int_{\tau_{i}}^{\tau}=\int_{\tau_{i}}^{\tau}-\int_{\tau_{i}}^{\tau_{0}}$).
 There is no tight-coupling approximation.
-Initial conditions are full ionization $X_{\mathrm{H}}^{+}=X_{\mathrm{He}}^{+}=1$, thermal equilibrium $T_{b}=T_{\gamma}$ ($\Delta T=0$), the arbitrary $\kappa=0$, and adiabatic perturbations $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$.
-The baryon species is parametrized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today and the primordial Helium mass fraction $Y_{\mathrm{He}}$.
+Initial conditions are fully ionized fractions $X_{\mathrm{H}}^{+}=X_{\mathrm{He}}^{+}=1$, temperatures $T_{b}=T_{\gamma}$ ($\Delta T=0$) in equilibrium, the arbitrary $\kappa=0$, and adiabatic perturbations $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$.
+The baryon species is parameterized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today, and the primordial Helium mass fraction $Y_{\mathrm{He}}$.
 SymBoltz solves thermodynamics equations together with the background equations, while some other codes treat these as separate stages.
 There is no meaningful performance improvement from doing this, as the size of the background (and thermodynamics) ODEs is so small.
-This makes a clear distinction between the background with all 0th order equations of motion, and the perturbations with all 1st order equations.
-It also makes it easy to create exotic models where the thermodynamics couple to the background, for example.
-Note that RECFAST uses fitting functions to emulate the results of more physically accurate and expensive simulations.
+Solving these stages together creates a clear distinction between the background with all zeroth-order equations of motion, and the perturbations with all first-order equations.
+It also makes it possible to create exotic models where the thermodynamics affect the background, for example.
+We note that RECFAST uses fitting functions to emulate the results of more physically accurate and expensive simulations.
 These are tuned to work for the $\mathrm{\Lambda CDM}$ model.
-SymBoltz would therefore benefit from including more physically accurate recombination models for safer use with modified models.
+SymBoltz would therefore benefit from including more realistic recombination models for safer use with modified models.
 A.5 Photons $(\gamma)$
-Photons are massless and therefore ultra-relativistic.
-Unlike non-relativistic particles, one must account for the direction $\cos\theta=\boldsymbol{p}\cdot\boldsymbol{k}/\lvert\boldsymbol{p}\rvert\lvert\boldsymbol{k}\rvert$ of their momenta $\boldsymbol{p}$ relative to the Fourier wavenumber $\boldsymbol{k}$.
+Photons are massless and therefore ultrarelativistic.
+Unlike nonrelativistic particles, we must account for the direction $\cos\theta=\boldsymbol{p}\cdot\boldsymbol{k}\,/\,\lvert\boldsymbol{p}\rvert\cdot\lvert\boldsymbol{k}\rvert$ of their momenta $\boldsymbol{p}$ relative to the Fourier wavenumber $\boldsymbol{k}$.
 This results in a theoretically infinite hierarchy of equations for Legendre multipoles $l$, which in practice must be truncated at some maximum multipole $l_{\text{max}}$:
@@
 $G_{l_{\text{max}}}^{\prime}=kG_{l_{\text{max}}-1}-\frac{l_{\text{max}}+1}{\tau}G_{l_{\text{max}}}+{\kappa^{\prime}}G_{l_{\text{max}}},\qquad\Pi=F_{2}+G_{0}+G_{2},\qquad{{\Theta}_{l}}=\frac{F_{l}}{4}.$
-The equations for $F_{l}^{\prime}$ and $G_{l}^{\prime}$ apply for $2\leq l<l_{\text{max}}$.
-There are no tight-coupling, radiation-streaming or ultra-relativistic fluid approximations.
-Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e.
-$\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e.
-$\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=-\frac{8}{15}\frac{k}{\kappa^{\prime}}F_{1}$, $G_{0}=\frac{5}{16}F_{2}$, $G_{1}=-\frac{1}{16}\frac{k}{\kappa^{\prime}}F_{2}$, $G_{2}=\frac{1}{16}F_{2}$, and $F_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}F_{l-1}$ and $G_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}G_{l-1}$ for $3\leq l\leq l_{\text{max}}$.
-The species is parametrized by its temperature $T_{0}$ today, which in turn sets the density parameters $\Omega_{0}=\frac{\pi^{2}}{15}\frac{(k_{B}T_{0})^{4}}{(\hbar c)^{3}}\frac{8\pi G}{3H_{0}^{2}}$ and $\rho_{0}=\frac{8\pi}{3}\Omega_{0}$ today.
+The equations for $F_{l}^{\prime}$ and $G_{l}^{\prime}$ hold for $2\leq l<l_{\text{max}}$.
+There are no tight-coupling, radiation-streaming or ultrarelativistic fluid approximations.
+Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e., $\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e., $\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=-\frac{8}{15}\frac{k}{\kappa^{\prime}}F_{1}$, $G_{0}=\frac{5}{16}F_{2}$, $G_{1}=-\frac{1}{16}\frac{k}{\kappa^{\prime}}F_{2}$, $G_{2}=\frac{1}{16}F_{2}$, $F_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}F_{l-1}$, and $G_{l}=-\frac{l}{2l+1}\frac{k}{\kappa^{\prime}}G_{l-1}$ for $3\leq l\leq l_{\text{max}}$.
+The species is parameterized by its temperature $T_{0}$ today, which in turn sets the density parameters $\Omega_{0}=\frac{\pi^{2}}{15}\frac{(k_{B}T_{0})^{4}}{(\hbar c)^{3}}\frac{8\pi G}{3H_{0}^{2}}$ and $\rho_{0}=\frac{8\pi}{3}\Omega_{0}$ today.
 A.6 Massless neutrinos $(\nu)$
 Massless neutrinos behave similarly to photons, but decouple from interactions with other species in the very early universe.
-One must only account for this interaction in initial conditions, while their evolution equations are a simpler case of the photons’:
+We must only account for this interaction in initial conditions, while their evolution equations are a simpler case of the photon equations:
 $T=\frac{T_{0}}{a},\qquad w=\frac{1}{3},\qquad{c_{s}^{2}}=\frac{1}{3},\qquad P=\frac{\rho}{3},\qquad\rho=\frac{\rho_{0}}{a^{4}},\qquad\delta={F_{0}},\qquad\theta=\frac{3}{4}kF_{1},\qquad\sigma=\frac{F_{2}}{2},$
 $F_{0}^{\prime}=-kF_{1}+4\Phi^{\prime},\qquad F_{1}^{\prime}=\frac{k}{3}\big({F_{0}}-2F_{2}+4\Psi\big),\qquad F_{l}^{\prime}=\frac{k}{2l+1}\big(lF_{l-1}-(l+1)F_{l+1}\big),\qquad F_{l_{\text{max}}}^{\prime}=kF_{l_{\text{max}}-1}-\frac{l_{\text{max}}+1}{\tau}F_{l_{\text{max}}}.$
-The equations for $F_{l}^{\prime}$ apply for $2\leq l<l_{\text{max}}$.
-There is no ultra-relativistic fluid approximation.
-Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e.
-$\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e.
-$\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=\frac{2}{15}(k\tau)^{2}\Psi$ and $F_{l}=\frac{l}{2l+1}k\tau F_{l-1}$.
-The species is parametrized by the effective number $N_{\text{eff}}$, the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ and temperature $T_{0}$ today.
+The equations for $F_{l}^{\prime}$ hold for $2\leq l<l_{\text{max}}$.
+There is no ultrarelativistic fluid approximation.
+Initial conditions are adiabatic with $F_{0}=-2\Psi$ (i.e., $\delta/(1+w)=-\frac{3}{2}\Psi$), $F_{1}=\frac{2}{3}k\tau\Psi$ (i.e., $\theta=\frac{1}{2}k^{2}\tau\Psi$), $F_{2}=\frac{2}{15}(k\tau)^{2}\Psi$, and $F_{l}=\frac{l}{2l+1}k\tau F_{l-1}$.
+The species is parameterized by the effective number $N_{\text{eff}}$, the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today, and temperature $T_{0}$ today.
 If photons are present, they default to $T_{\nu 0}=\big(\frac{4}{11}\big)^{1/3}T_{\gamma 0}$ and $\Omega_{\nu 0}=N_{\text{eff}}\frac{7}{8}\big(\frac{4}{11}\big)^{4/3}\Omega_{\gamma 0}$.
@@
 Massive neutrinos are the most complicated species in the $\mathrm{\Lambda CDM}$ model (alongside baryon recombination).
-In essence, the species we have looked at so far have Boltzmann equations where the momenta of their distribution function can be integrated out analytically in non-relativistic and ultra-relativistic limits.
-This means that their stress-energy components are linked by trivial equations of state and sound speeds, for example, and their effect can be parametrized by a simple density parameter $\Omega_{0}$.
-On the other hand, massive neutrinos have intermediate masses that fall between the non-relativistic and ultra-relativistic limits.
+In essence, the species we have looked at so far have Boltzmann equations where the momenta of their distribution function can be integrated out analytically in nonrelativistic or ultrarelativistic limits.
+This means that their stress-energy components are linked by trivial equations of state and sound speeds, for example, and their effect can be parameterized by a simple density parameter $\Omega_{0}$.
+On the other hand, massive neutrinos have intermediate masses that fall between the nonrelativistic and ultrarelativistic limits.
 Integrals over their distribution function must be computed numerically.
-This is very expensive if done naively, and it is extremely important to choose a quadrature scheme that minimizes the number of sampled points.
-Fortunately, the momentum integrals have a structure that can be exploited: they are all in the form weighted form $I[g(x)]=\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)$, where $f(x)=1/(e^{x}+1)$ is the equilibrium distribution function and $g(x)$ is an arbitrary function of the dimensionless momentum $x=pc/k_{B}T$ (see Ma and Bertschinger (1995) for more details).
-One can generally approximate $I[g(x)]\approx\sum_{i}W_{i}\,g(x_{i})$ with a weighted quadrature scheme with points $x_{i}$ and weights $W_{i}$ (more on this after the equations).
+This is very expensive if done naively, and it is extremely important to choose a quadrature scheme that samples as few momenta as possible.
+Fortunately, the momentum integrals have a structure that can be exploited: they are all in the weighted form $I[g(x)]=\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)g(x)$, where $f(x)=1/(e^{x}+1)$ is the equilibrium distribution function and $g(x)$ is an arbitrary function of the dimensionless momentum $x=pc/k_{B}T$ (see Ma and Bertschinger 1995 for more details).
+We can generally approximate $I[g(x)]\approx\sum_{i}W_{i}\,g(x_{i})$ with a weighted quadrature scheme with points $x_{i}$ and weights $W_{i}$ (more on this after the equations).
 In other words, the integral operator $\int_{0}^{\infty}\mathrm{d}x\,x^{2}f(x)$ is effectively replaced by the discrete summation operator $\sum_{i}\!W_{i}$ for some weights $W_{i}$.
@@
 $I_{0}=\sum_{i}\!W_{i}E_{i}\psi_{i,0},\qquad I_{1}=\sum_{i}\!W_{i}x_{i}\psi_{i,1},\qquad I_{2}=\sum_{i}\!W_{i}\frac{x_{i}^{2}}{E_{i}}\psi_{i,2},\qquad\delta=\frac{I_{0}}{I_{\rho}},\qquad\sigma=\frac{2I_{2}}{3I_{\rho}+I_{P}},\qquad u=\frac{3I_{1}}{3I_{\rho}+I_{P}},\qquad\theta=ku.$
-Initial conditions are $\psi_{i,0}=-\frac{1}{4}(-2\Psi)\big(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\big)_{i}$, $\psi_{i,1}=-\frac{1}{3}\frac{E_{i}}{x_{i}}\frac{1}{2}k\tau\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\big)_{i}$, $\psi_{i,2}=-\frac{1}{2}\frac{1}{15}(k\tau)^{2}\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d\ln x}}\big)_{i}$ and $\psi_{i,l}=0$.
-This integrates to adiabatic $\delta/(1+w)$, $\theta$ and $\sigma$ similarly to massless neutrinos.
-Free parameters are the temperature today $T_{0}$, mass $m$ of a single neutrino and degeneracy factor $N=\sum_{i=1}^{N}m_{i}/m$ for describing multiple neutrinos with equal mass.
-The degeneracy factor defaults to $N=3$, and the temperature to $T_{h0}=\big(\frac{4}{11}\big)^{1/3}T_{\gamma 0}$ if photons are present, as for massless neutrinos.
-Here the equation for $\psi_{i,l}^{\prime}$ applies for $2\leq l\leq l_{\text{max}}$, and expressions $g_{i}=g(x_{i})$ indexed by $i$ are evaluated with the momentum quadrature point $x=x_{i}$.
+Initial conditions are $\psi_{i,0}=-\frac{1}{4}(-2\Psi)\big(\frac{\mathrm{d}\ln f}{\mathrm{d}\ln x}\big)_{i}$, $\psi_{i,1}=-\frac{1}{3}\frac{E_{i}}{x_{i}}\frac{1}{2}k\tau\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d}\ln x}\big)_{i}$, $\psi_{i,2}=-\frac{1}{2}\frac{1}{15}(k\tau)^{2}\Psi\big(\frac{\mathrm{d}\ln f}{\mathrm{d}\ln x}\big)_{i}$, and $\psi_{i,l}=0$.
+When integrated over momenta, they are equivalent to adiabatic $\delta/(1+w)$, $\theta$, and $\sigma$, similarly to massless neutrinos.
+Free parameters are the temperature $T_{0}$ today, single neutrino mass $m$, and degeneracy factor $N=\sum_{i=1}^{N}m_{i}/m$ for describing multiple neutrinos with the same mass.
+They default to $N=3$, and $T_{h0}=\big(\frac{4}{11}\big)^{1/3}T_{\gamma 0}$ if photons are present, as for massless neutrinos.
+Here, the equations for $\psi_{i,l}^{\prime}$ hold for $2\leq l\leq l_{\text{max}}$, and the $i$-indexed expressions $g_{i}=g(x_{i})$ are evaluated with the momentum quadrature point $x=x_{i}$.
 The reduction to dimensionless momenta $x=pc/k_{B}T$ (the argument of $\exp$ in $f$) is deliberate because it makes numerics more well-defined and the quadrature scheme independent of $m$ and all other cosmological parameters.
@@
 The substituted integrand is then passed to an adaptive algorithm in QuadGK.jl [^23] https://github.com/JuliaMath/QuadGK.jl that computes quadrature points $u_{i}$ and weights $W_{i}$ by performing weighted integrals against several test functions $g(x)$.
-Finally, the corresponding momenta $x_{i}=x(u_{i})$ are returned along with the weights $W_{i}$, from which one can approximate the integral $I[g(x)]\approx\sum_{i}W_{i}g(x_{i})$ against any $g(x)$.
-SymBoltz tests this numerical quadrature scheme against the analytical result $I[x^{n-2}]=\int_{0}^{\infty}\mathrm{d}x\,x^{n}/(e^{x}+1)=(1-2^{-n})\zeta(n+1)\Gamma(n+1)$ for $2\leq n\leq 8$.
-We assume this to be a reasonable test for the integrals encountered in the equations above.
-Agreement is excellent with $L=100$, which yields relative errors below $10^{-6+n-N}$ for all $2\leq n\leq 8$ and $1\leq N\leq 5$.
+Finally, the corresponding momenta $x_{i}=x(u_{i})$ are returned along with the weights $W_{i}$, from which we can approximate the integral $I[g(x)]\approx\sum_{i}W_{i}g(x_{i})$ against any $g(x)$.
+We test this numerical quadrature scheme against the analytical result $I[x^{n-2}]=\int_{0}^{\infty}\mathrm{d}x\,x^{n}/(e^{x}+1)=(1-2^{-n})\zeta(n+1)\Gamma(n+1)$ for $2\leq n\leq 8$.
+We assume this is a reasonable test for the integrals encountered in the equations above.
+The agreement is excellent with $L=100$, which yields relative errors below $10^{-6+n-N}$ for all $2\leq n\leq 8$ and $1\leq N\leq 5$.
 SymBoltz defaults to $N=4$ momenta, for which this relative error is less than $10^{-6}$ for $n\leq 4$, for example.
 It also agrees well with CLASS using default settings.
-Note that this momentum quadrature strategy is generic with respect to the distribution function $f(x)$ and substitution $u(x)$, so it can easily be modified for other particle species whose distribution function cannot be integrated out.
-CAMB (Lewis, 2025) and CLASS (Lesgourgues and Tram, 2011) apply similar weighted quadrature strategies.
+We note that this momentum quadrature strategy is generic with respect to the distribution function $f(x)$ and substitution $u(x)$, so it can easily be modified for other particle species whose distribution function cannot be integrated out.
+CAMB [^24] https://cosmologist.info/notes/CAMB.pdf and CLASS (Lesgourgues and Tram 2011) apply similar weighted quadrature strategies.
 They also get away with only a handful of sampled momenta, but the precise details of the computation differ slightly.
@@
 $w=-1,\qquad\rho=\rho_{0},\qquad P=-\rho,\qquad\delta=0,\qquad\theta=0,\qquad\sigma=0,\qquad u=0.$
-It is parametrized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today.
-This is set to $\Omega_{\Lambda 0}=1-\sum_{s\neq\Lambda}\Omega_{s0}$ if all species $s$ have a $\Omega_{s0}$ parameter and General Relativity is the theory of gravity.
-This constraint comes from the 1st Friedmann equation today.
+It is parameterized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today.
+If all species $s$ have a $\Omega_{s0}$ parameter and general relativity is the theory of gravity, it is set to $\Omega_{\Lambda 0}=1-\sum_{s\neq\Lambda}\Omega_{s0}$.
+This constraint comes from the first Friedmann equation today.
 A.9 Primordial power spectrum ($I$)
-SymBoltz computes the inflationary primordial power spectrum parametrized by the amplitude $A_{s}$ and tilt $n_{s}$:
+SymBoltz computes the inflationary primordial power spectrum parameterized by the amplitude $A_{s}$ and tilt $n_{s}$:
 $P_{0}(k)=\frac{2\pi^{2}}{k^{3}}A_{s}\bigg(\frac{k}{k_{\text{p}}}\bigg)^{n_{s}-1}.$
 A.10 Matter power spectrum
-SymBoltz computes the matter power spectrum for some desired set of species $s$, which are presumably matter-like at late times (e.g. $s\in\{c,b,h\}$):
+SymBoltz computes the matter power spectrum for some desired set of species $s$, which are presumably matter-like at late times (e.g., $s\in\{c,b,h\}$):
 $P(k,\tau)=P_{0}(k)\big|\Delta(\tau,k)\big|^{2}\quad\text{with}\quad\Delta=\delta+\frac{3\mathscr{H}}{k^{2}}\theta=\frac{\sum_{s}\delta\rho_{s}}{\sum_{s}\rho_{s}}+\frac{3\mathscr{H}}{k^{2}}\frac{\sum_{s}(\rho_{s}+P_{s})\theta_{s}}{\sum_{s}(\rho_{s}+P_{s})}.$
-Here $\Delta$ is the total gauge-independent overdensity with total $\delta$ and $\theta$ computed by summing the components of the energy-momentum tensor that are additive.
+Here, $\Delta$ is the total gauge-independent overdensity with total $\delta$ and $\theta$ computed by summing the components of the energy-momentum tensor that are additive.
 A.11 CMB power spectrum and line-of-sight integration
-SymBoltz finds photon temperature and polarization multipoles today for any $l$ by computing the line-of-sight integrals
+SymBoltz finds photon temperature and polarization multipoles today for any $l$ by computing line-of-sight integrals:
 $\Theta_{l}^{\mathrm{T}}\big(\tau_{0},k\big)=\int_{\tau_{i}}^{\tau_{0}}S_{T}(\tau,k)j_{l}\big(k(\tau_{0}-\tau)\big)\mathrm{d}\tau\quad\text{with}\quad S_{T}=v\bigg(\frac{\delta_{\gamma}}{4}+\Psi+\frac{\Pi_{\gamma}}{16}\bigg)+e^{-\kappa}\big(\Psi+\Phi\big)^{\prime}+\frac{\big(vu_{b}\big)^{\prime}}{k}+\frac{3}{16k^{2}}\big(v\Pi_{\gamma}\big)^{\prime\prime},$
@@
 This drastically speeds up the computation over including all $l$ in an enormous set of coupled perturbation ODEs.
-SymBoltz performs the integrals with the trapezoid method using the substitution $u(\tau)=\tanh(\tau)$, which adds more points in the early universe when sampled uniformly, using $768$ points by default.
-Here $j_{l}$ are the spherical Bessel functions of the first kind.
-SymBoltz is not yet generalized to non-flat geometries, where they are replaced by hyperspherical functions.
+SymBoltz performs the integrals with the trapezoid method using around $1250$ uniformly sampled $\tau$, by default.
+Here, $j_{l}$ are the spherical Bessel functions of the first kind.
+SymBoltz is not yet generalized to curved spacetimes, where they are replaced by hyperspherical functions.
 The cross-correlated angular spectrum between $\mathrm{A},\mathrm{B}\in\{\mathrm{T},\mathrm{E}\}$ is then computed from
@@
 First, the source function $S_{i}(\tau)=S(\tau,k_{i})$ is found on an initial grid of wavenumbers $k_{i}$ by integrating their perturbations.
-The method then iterates over each interval $(k_{i},k_{i+1})$, integrates the perturbations for the center $k_{i+1/2}=(k_{i}+k_{i+1})/2$ and finds $S_{i+1/2}(\tau)=S(\tau,k_{i+1/2})$.
-The idea is then to compare this to a linearly interpolated source $S_{\text{int}}(\tau)=(S_{i}(\tau)+S_{i+1}(\tau))/2$ with some error measure err, for example from $\texttt{err}^{2}=\int\mathrm{d}\tau\big(S_{\text{int}}(\tau)-S_{i+1/2}(\tau)\big)^{2}$.
+The method then iterates over each interval $(k_{i},k_{i+1})$, integrates the perturbations for the center $k_{i+1/2}=(k_{i}+k_{i+1})/2$, and finds $S_{i+1/2}(\tau)=S(\tau,k_{i+1/2})$.
+The idea is then to compare this to a linearly interpolated source $S_{\text{int}}(\tau)=(S_{i}(\tau)+S_{i+1}(\tau))/2$ with some error measure err, for example from the $L_{2}$-like $\texttt{err}^{2}=\int\mathrm{d}\tau\big(S_{\text{int}}(\tau)-S_{i+1/2}(\tau)\big)^{2}$.
 If the interpolation is poor and err is greater than some absolute and relative tolerances, the left and right half-intervals $(k_{i},k_{i+1/2})$ and $(k_{i+1/2},k_{i+1})$ are recursively refined in the same way.
-This gives $S(\tau,k)$ by integrating as few kk–modes as possible.
-Instead of manually creating a non-uniform $k$-grid with many hyperparameters, the process is controlled by only one tolerance.
+This gives $S(\tau,k)$ while integrating as few kk–modes as possible.
+Instead of manually creating a nonuniform $k$-grid with many hyperparameters, the process is controlled by only one tolerance.
 It parallelizes over each interval and can also be done in $\log k$.
 Appendix B Precision parameters for CLASS
-When comparing results to CLASS in section 3, we use CLASS version 3.3.4 with the following precision parameters: {codebox}
+When comparing results to CLASS in Sect. 3, we use CLASS version 3.3.4 with the following precision parameters: {codebox}
 tight_coupling_approximation = 5 # compromise_CLASS; cannot disable TCAtight_coupling_trigger_tau_c_over_tau_h = 0.01 # turn off TCA earliertight_coupling_trigger_tau_c_over_tau_k = 0.001 # turn off TCA earlierradiation_streaming_approximation = 3 # turn off RSAncdm_fluid_approximation = 3 # turn off NCDMFAur_fluid_approximation = 3 # turn off UFAevolver = 1 # for implicit ndf15; 0 for explicit rktol_perturbations_integration = 1.0e-6 # less noise; varied in work-precision diagramtol_ncdm_newtonian = 1.0e-5 # default sampling of massive neutrino momentabackground_Nloga = 6000 # for stable derivativesWe run CLASS with output = mPk to compute only $P(k)$, and output = tCl, pCl to compute only $C_{l}^{\text{TT}}$, $C_{l}^{\text{EE}}$ and $C_{l}^{\text{TE}}$ with the code’s most specialized and optimized paths.
-CLASS and SymBoltz are configured to use the same Newtonian gauge, $l_{\text{max}}$ for all species, RECFAST recombination model, $\tanh$-like reionization parametrization, $w_{0}w_{a}$ dark energy equations and cosmological parameter values.
+CLASS and SymBoltz are configured to use the same Newtonian gauge, $l_{\text{max}}$ for all species, RECFAST recombination model, $\tanh$-like reionization parameterization, $w_{0}w_{a}$ dark energy equations, and cosmological parameter values.
 The settings above disable as many approximations as possible and reduce the impact of the tight-coupling approximation, which cannot be disabled.
@@
 Likewise, the tolerance for the perturbations ODEs is set one notch smaller than the default $10^{-5}$ to make the output stable.
-We use CLASS’ default sampling of massive neutrino momenta, as discussed in section 4.
-When running CLASS with all approximations on in tables 1 and 7, the first block of parameters above are not used.
+We use CLASS’ default sampling of massive neutrino momenta, as discussed in Sect.
+4.
+When running CLASS with all approximations on in Table 1 and Fig.
+7, the first block of parameters above are not used.
 In the work-precision comparison in Fig.
@@
 SymBoltz’ code repository is set up with continuous integration that runs several tests and builds updated documentation pages every time changes to the code are committed.
-In particular, this compares the solution for $\mathrm{\Lambda CDM}$ with CLASS for many variables solved by the background, thermodynamics and perturbations (using the options write_background, write_thermodynamics and k_output_values).
-These are the basis for all derived quantities like luminosity distances, matter and CMB power spectra, which are also compared.
+In particular, this compares the solution for $\mathrm{\Lambda CDM}$ with CLASS for many variables solved by the background, thermodynamics, and perturbations (using the options write_background, write_thermodynamics, and k_output_values).
+These are the basis for all derived quantities, such as luminosity distances, matter, and CMB power spectra, which are also compared.
 The checks pass when the quantities agree within a small tolerance.
-We do not compare directly against more codes like CAMB, but CLASS has already been compared extensively with CAMB with excellent agreement (Lesgourgues, 2011b).
-The comparison takes a lot of space and is not included here, but is found in the documentation linked from SymBoltz’ repository.
-Another test checks that integration of the background and perturbations equations are stable throughout parameter space.
-As the equations are very stiff and SymBoltz does not rely on approximations for relieving it, one could imagine that the integration would be stable for some parameter values and unstable for others.
-The test creates a box in parameter space $\pm 50\%$ around a fiducial set of realistic parameter values, draws several sets of parameter values from that space with Latin hypercube sampling (to efficiently cover parameter space) and integrates the background and perturbations for each such set.
-All parameter samples are found to integrate successfully without warnings and errors.
+We do not compare directly against additional codes such as CAMB, but CLASS has already been compared extensively with CAMB with excellent agreement (Lesgourgues 2011b).
+This comparison is found in SymBoltz’ documentation.
+[^25] https://hersle.github.io/SymBoltz.jl/stable/comparison/
+Another test checks that integration of the background and perturbations equations is stable throughout parameter space.
+As SymBoltz integrates stiff equations without approximations, we could imagine that the integration would be stable for some parameter values and unstable for others.
+The test creates a box in parameter space $\pm 50\%$ around a fiducial set of realistic parameter values, draws several sets of parameter values from that space with Latin hypercube sampling (to efficiently cover the box), and integrates the background and perturbations for each such set.
+All samples are found to integrate successfully.